### v2

# On degenerate metric resolvent: local monotonicity, restricted maximality and beyond

Abstract
We study the basic properties of resolvent equipped with degenerate metric, under local monotonicity and restricted maximality, and further propose several conditions for the well-defineness and convergence of the fixed point iterations. The results help to understand the behaviours of many operator splitting algorithms in the kernel space of degenerate metric.
keywords:
Resolvent, degenerate metric, proximal point algorithm (PPA), operator splitting algorithms, local monotonicity, restricted maximality.
{AMS}
47H05, 49M29, 49M27, 90C25
1 Introduction
1.1 Motivation
In nonlinear analysis and optimization, a generic and central assignment is to obtain numerical solutions to the inclusion problem
$\textrm{find }x\in\mathcal{H},\quad\textrm{s.t. }0\in\mathcal{A}x$ (1)
where $\mathcal{H}$ is a real Hilbert space and $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is a set-valued operator. Built on the classical proximal point algorithm (PPA) in 1970s [25, 28], many splitting methods have been developed in the literature during the past five decades. Very recently, a degenerate preconditioned version of PPA was studied in [11], which provided a powerful framework for analyzing various existing splitting methods and also for devising new and flexible schemes. The degenerate preconditioned PPA can be described by the following fixed-point iterate
$x^{k+1}=\mathcal{T}x^{k},\quad\textrm{where\ }\mathcal{T}:=(\mathcal{Q}+\mathcal{A})^{-1}\mathcal{Q},$ (2)
where $\mathcal{Q}:\mathcal{H}\mapsto\mathcal{H}$ is a linear, bounded, self-ajoint and positive semi-definite operator. To ensure the iterate to be well-defined, the operate $\mathcal{T}$ has to be single-valued and have full domain. However, this key condition was only stated as a basic assumption in [11, Definition 2.1] and hence motivates us to find sufficient guarantees for the operator $\mathcal{T}$.
Among many operator splitting algorithms prevails the degeneracy phenomenon, which makes the behaviours of (2) more subtle. For instance, the celebrated Douglas–Rachford splitting (DRS), for solving $\min_{u\in\mathcal{U}}(u)+g(u)$, reads as
$\left\lfloor\begin{array}[]{lll}u^{k+1}&=&\arg\min_{u}g(u)+\frac{1}{2\tau}\big{\|}u-z^{k}\big{\|}^{2},\\ w^{k+1}&=&\arg\min_{w}f(w)+\frac{1}{2\tau}\big{\|}w-(2u^{k+1}-z^{k})\big{\|}^{2},\\ z^{k+1}&=&z^{k}+(w^{k+1}-u^{k+1}),\end{array}\right.$ (3)
which exactly fits the form of (2) according to
$\mathcal{H}=\mathcal{U}^{3},\ x^{k}=\begin{bmatrix}u^{k}\\ w^{k}\\ z^{k}\end{bmatrix},\ \mathcal{A}=\begin{bmatrix}\tau\partial g&I&-I\\ -I&\tau\partial f&I\\ I&-I&0\end{bmatrix},\ \mathcal{Q}=\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&0&I\end{bmatrix}.$ (4)
Here, we adopt a different (but essentially equivalent) scheme (3) from [11, Eq.(1.6)] for two reasons: (i) (3) explicitly preserves both solution trajectory $\{u^{k}\}_{k\in\mathbb{N}}$ and shadow sequence $\{z^{k}\}_{k\in\mathbb{N}}$; (ii) the degenerate metric $\mathcal{Q}$ in (4) is easier to analyze than that one in [11, Eq.(1.5)]. Then, one may ask: Can we conclude from (4) only (without ad hoc analysis) the well-definedness of the iterations and the convergence of $\{u^{k}\}_{k\in\mathbb{N}}$? More broadly, this further motivates us to study a general setting of (2) and find the sufficient conditions on $\mathcal{A}$ and $\mathcal{Q}$, to guarantee the well-definedness of $\mathcal{T}$ and the convergence of $\{x^{k}\}_{k\in\mathbb{N}}$.
1.2 Related works
The metric resolvent $\mathcal{T}$ was also coined the term warped resolvent in [14, Definition 1.1] or $F$-resolvent [8, Definition 4.1] in various contexts. [8, Sect. 3] assumed $\mathcal{Q}$ to be maximally monotone, strictly monotone, 3 monotone and surjective; [14, Definition 1.1] assumed $\mathsf{ran}\mathcal{Q}\subseteq\mathsf{ran}(\mathcal{A}+\mathcal{Q})$ and disjoint injectivity of $\mathcal{A}+\mathcal{Q}$, and [14, Proposition 3.9] further listed specific requirements on $\mathcal{A}$ and $\mathcal{Q}$, such as surjectivity, uniform/strict/strong monotonicity. However, they are much stronger than, and fundamentally different from our degenerate setting.
To our knowledge, [17, Theorem 3.3] first studied the convergence for the limiting case of Chambolle–Pock algorithm [15], which actually corresponds to a degenerate metric $\mathcal{Q}$ (see [35, Eq.(32)] for details). This technique was later adopted in [22, Theorem 3.4], which extended the discussion to more general problems in a finite-dimensional setting. At almost the same time of [22], [12, Sect. 3] also discussed the degenerate metric for a class of preconditioned ADMM in an infinite dimensional space. [11, Sect. 2], where the authors performed a systematic study for this issue. The works of [35, 37] further exemplified the degeneracy with many practical examples and developed the reduced forms of resolvent to remove the redundancy. more theoretical results of [11] on some special cases have been reported in [7]. The degeneracy has been further extended to derive various new algorithms [36, 1, 31, 24].
1.3 Contributions
This work aims to provide a fundamental analysis and understanding of the degenerate metric resolvent. The contributions are in order.
(1) Sect. 2 studies the single-valued, $\mathcal{Q}$-firmly nonexpansive and $\mathcal{Q}$-demiclosed properties of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ under local monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$, which greatly weakens the usual monotone property of $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ in literature. Moreover, Lemma 2.14 develops several equivalent forms of $\mathcal{T}$ based on a conjugate view of $\mathcal{A}^{-1}$ via generalized Moreau’s identity.
(2) We in Sect. 3 give the full domain condition based on a conjugate view and develop a restricted Minty’s theorem based on a notion of restricted maximaly monotonicity and propose specific conditions on $\mathcal{A}$ and $\mathcal{Q}$ for the full domain. Sect. 4 further presents the sufficient and necessary conditions for the single-valuedness of $\mathcal{T}$ from both primal and conjugate perspectives. Our results extends the existing ones, from injectivity to restricted injectivity.
(3) We in Theorem 5.4 first show the weak convergence in the range space of $\mathcal{Q}$ under restricted maximal monotonicity only, and further show in Theorem 5.11 the whole convergence of $\{x^{k}\}_{k\in\mathbb{N}}$ under additional conditions. The comparisons with existing works are also discussed in details. Finally, Sect. 6 exemplifies the results with several splitting algorithms.
1.4 Notations, assumptions and preliminaries
We use standard notations and concepts from convex analysis and variational analysis, which can all be found in the classical and recent monographs [29, 30, 5, 9].
[…]
As mentioned in Sect. 1.1, we focus on linear metric. More specifically, we put the following assumption on $\mathcal{Q}$ throughout the paper:
Assumption 1
(cf. [35, Assumption 2.1])
(i) [Basic] $\mathcal{Q}:\mathcal{H}\mapsto\mathcal{H}$ is linear, bounded, self-adjoint and positive semi-definite, and $\mathsf{dom}\mathcal{Q}=\mathcal{H}$;
(ii) [Degeneracy] $\mathcal{Q}$ is degenerate, i.e., $\mathsf{ker}\mathcal{Q}\backslash\{0\}\neq\varnothing$;
(iii) [Closedness] $\mathcal{Q}$ is closed, i.e., $\overline{\mathsf{ran}\mathcal{Q}}=\mathsf{ran}\mathcal{Q}$.
Remark 1
(i) In particular, we assume the closedness of range of $\mathcal{Q}$, since (1) it is often satisfied by many splitting algorithms; (2) it considerably simplifies the analysis. A simple criterion for the closedness is: $\exists\alpha>0$, such that $\|x\|_{\mathcal{Q}}^{2}\geq\alpha\|x\|^{2},\ \forall x\in\mathsf{ran}\mathcal{Q}$ [5, Fact 2.26].
(ii) Assumption 1 is much weaker than the rigorous conditions of $\mathcal{Q}$ listed in [14, Proposition 3.9], e.g., strict monotonicity, surjectivity. We do not put assumption on $\mathcal{A}$ here, since the monotonicity of $\mathcal{A}$ should be closely linked to the degeneracy of $\mathcal{Q}$, and deserves particular investigation in the main part.
Before our exposition, some basic but important properties of $\mathcal{Q}$ are collected below.
Fact 1
Under Assumption 1, the following hold.
(i) [5, Fact 2.25] An orthogonal decomposition of $\mathcal{H}$ is: $\mathcal{H}=\mathsf{ran}\mathcal{Q}+\mathsf{ker}\mathcal{Q}$.
(ii) [5, Proposition 3.30] Any point $x\in\mathcal{H}$ can be decomposed as $x=x_{\mathsf{r}}+x_{\mathsf{k}}$, where $x_{\mathsf{r}}:=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x$ and $x_{\mathsf{k}}:=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}x$ are the $\mathsf{ran}\mathcal{Q}$- and $\mathsf{ker}\mathcal{Q}$-parts of $x$.
(iii) [10, Theorem 4] $\mathcal{Q}$ admits an (actually unique) self-adjoint square-root, denoted as $\sqrt{\mathcal{Q}}$, such that $\mathcal{Q}=\sqrt{\mathcal{Q}}\sqrt{\mathcal{Q}}$.
[…]
Fact 2
Given $\mathcal{Q}$ satisfying Assumption 1, then $\langle\cdot|\cdot\rangle_{\mathcal{Q}}$ is weak-strong continuous, i.e., if $x^{k}\rightharpoonup x^{\star}$ and $u^{k}\rightarrow u^{\star}$, then $\langle x^{k}|u^{k}\rangle_{\mathcal{Q}}\rightarrow\langle x^{\star}|u^{\star}\rangle_{\mathcal{Q}}$.
Proof 1.1.
We deduce that
$\langle x^{k}|u^{k}\rangle_{\mathcal{Q}}-\langle x^{\star}|u^{\star}\rangle_{\mathcal{Q}}$ $=$ $\langle x^{k}-x^{\star}|u^{\star}\rangle_{\mathcal{Q}}+\langle x^{k}|u^{k}-u^{\star}\rangle_{\mathcal{Q}}$
$\leq$ $\langle x^{k}-x^{\star}|\mathcal{Q}u^{\star}\rangle+\big{\|}\mathcal{Q}x^{k}\big{\|}\cdot\big{\|}u^{k}-u^{\star}\big{\|}$
$\rightarrow$ 0,since xkx0 and $0,\qquad\text{since $x^{k}-x^{\star}\rightharpoonup 0$ and $u^{k}-u^{\star}\rightarrow 0$}$
which shows that $\langle x^{k}|u^{k}\rangle_{\mathcal{Q}}-\langle x^{\star}|u^{\star}\rangle_{\mathcal{Q}}\rightarrow 0$.
2 Local monotonicity related to degenerate metric
Almost all of the existing results of metric resolvent were developed based on the usual assumption that $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is maximally monotone. This part will show that this assumption can be weakened to local monotonicity under degenerate metric.
Sect. 24 focus on the single step of (2), which is rewritten for simplicity as:
$0\in\mathcal{A}y+\mathcal{Q}(y-x)\Longleftrightarrow y\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x:=\mathcal{T}x.$ (5)
Here, we use $y\in\mathcal{T}x$, since $\mathcal{T}$ is not guaranteed to be single-valued for the moment.
2.1 Motivations and definitions
Observing (5), the monotonicity of $\mathcal{A}$ should be associated with the degeneracy of $\mathcal{Q}$, which gives rise to the following important definition, which also appeared in [11].
Definition 2.1.
Under the condition of $\mathsf{ran}\mathcal{A}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing$, $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ is monotone, if
$\langle x-y|u-v\rangle\geq 0,\ \forall(x,u)\in\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q}),\ \forall(y,v)\in\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q}).$
This is similar to the notion of so-called local monotonicity [5, Definition 25.7, Proposition 25.8], which defines the monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times C)$, where $C$ is an open convex subset of $\mathcal{H}$, such that $C\cap\mathsf{ran}\mathcal{A}\neq\varnothing$. Our Definition 2.1 uses a subspace $\mathsf{ran}\mathcal{Q}$ instead of $C$. The essential difference here is that $\mathsf{int}\ \mathsf{ran}\mathcal{Q}=\varnothing$ but $\mathsf{int}\ C\neq\varnothing$. We will see in Sect. 3 that due to this difference, it is difficult to incorporate maximality into this local monotonicity.
Definition 2.1 requires $\mathcal{A}$ to be monotone in $\mathsf{ran}\mathcal{A}\cap\mathsf{ran}\mathcal{Q}$ only, regardless of the behaviours of $\mathcal{A}$ in $\mathsf{ran}\mathcal{A}\cap\mathsf{ker}\mathcal{Q}$. Clearly, this definition is much weaker than the usual (globally) monotone assumption of $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$, since nothing is enforced for $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ker}\mathcal{Q})$ in Definition 2.1.
Definition 2.1 was used in [11, Definition 2.4, Proposition 2.5] to describe $\mathcal{Q}$-monotonicity based on an inverse image of degenerate metric $\mathcal{Q}$. As a comparison, we now take a conjugate [^1] Let us explain why it is called ‘conjugate’. If $\mathcal{A}=\partial f$ for some $f\in\Gamma(\mathcal{H})$, then $\mathcal{A}^{-1}=\partial f^{*}$, where $f^{*}$ is a conjugate function of $f$. Hence, we call the analysis based on $\mathcal{A}^{-1}$ as a conjugate view. route based on $\mathcal{A}^{-1}$ and obtain the following basic observation.
Fact 1.
$\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ is monotone, if and only if $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone.
Fact 1 transforms the local monotonicity to that of $\mathcal{A}^{-1}$. This will lead us to a conjugate perspective in later developments. This fact is an extension of [5, Proposition 20.10] to the local monotonicity within a subspace. The basic condition of $\mathsf{ran}\mathcal{A}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing$ in Definition 2.1 is also closely related to the monotonicity of $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$. Indeed, if $\mathsf{ran}\mathcal{A}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing$, then $\exists x\in\mathcal{H}$, such that $\mathcal{Q}x\in\mathsf{ran}\mathcal{A}=\mathsf{dom}\mathcal{A}^{-1}$. It implies $\mathsf{dom}(\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}})\supseteq\{\mathcal{Q}x\}\neq\varnothing$. This is a very basic prerequisite for any viable operator.
2.2 Basic properties of $\mathcal{T}$ under local monotonicity
If we drop the usual assumption of $\mathcal{A}$ as being maximal monotone from $\mathcal{H}$ to $2^{\mathcal{H}}$, it should be careful to enforce new conditions on $\mathcal{A}$ as a new start, such that the requirements on $\mathcal{A}$ are kept minimal. In addition, since we has not dealt with the full domain issue of $\mathcal{T}$ (which will be discussed in Sect. 3), we use $\mathsf{dom}\mathcal{T}$ in this section. We for the moment put Definition 2.1 as
Assumption 2
(i) [Basic viability] $\mathsf{ran}\mathcal{A}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing$;
(ii) [Local monotonicity] $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ is monotone.
We are going to show that many nice properties of $\mathcal{T}$ presented in literature still hold under local monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ only.
Many results rely on the following basic but important observation.
Fact 2.
Given $\mathcal{T}$ as (5) under Assumption 1, the following hold:
(i) $\mathcal{Q}=\mathcal{Q}\circ\mathcal{P}_{\mathsf{ran}\mathcal{Q}}=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{Q}$;
(ii) $\mathcal{T}\circ\mathcal{P}_{\mathsf{ran}\mathcal{Q}}=\mathcal{T}$, i.e., $\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}=\mathcal{T}$.
2.2.1 Single-valuedness of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$
The following simple but crucial fact will be used repeatedly in our later developments.
Lemma 2.2.
Under Assumptions 1 and 2, $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ (or equivalently saying, $\mathcal{Q}\circ\mathcal{T}$) is always single-valued.
Proof 2.3.
Suppose $y_{1}$ and $y_{2}$ are two distinct outcomes of the operator $\mathcal{T}$ under the same input $x\in\mathsf{dom}\mathcal{T}$, i.e., $y_{1}\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$ and $y_{2}\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$. Then, we have
$\mathcal{Q}(x-y_{1})\in\mathcal{A}y_{1},\quad\mathcal{Q}(x-y_{2})\in\mathcal{A}y_{2}.$ (6)
Then,
[…]
$=$ $\big{\langle}y_{1}-y_{2}|\mathcal{Q}(x-y_{1})-\mathcal{Q}(x-y_{2})\big{\rangle}$
[…]
which yields $\|y_{1}-y_{2}\|_{\mathcal{Q}}=0$, i.e., $\mathcal{Q}y_{1}=\mathcal{Q}y_{2}$, or equivalently, $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}y_{1}=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}y_{2}$.
2.2.2 $\mathcal{Q}$-firm nonexpansiveness of $\mathcal{T}$
[8, Proposition 4.2] and [12, Lemma 3.2] showed that $\mathcal{T}$ is $\mathcal{Q}$-firmly nonexpansive under maximally monotone $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ and non-degenerate $\mathcal{Q}$. This was also discussed in [34, Sect. 2]. As emphasized in [11, Lemma 2.6], this key property is preserved under local monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ for degenerate case.
Fact 3.
[11, Lemma 2.6] Under Assumptions 1 and 2, $\mathcal{T}$ given as (2) is $\mathcal{Q}$-firmly nonexpansive, i.e.,
$\big{\|}\mathcal{T}x-\mathcal{T}y\big{\|}_{\mathcal{Q}}^{2}\leq\big{\langle}\mathcal{T}x-\mathcal{T}y\big{|}\mathcal{Q}(x-y)\big{\rangle},\quad\forall x,y\in\mathsf{dom}\mathcal{T},$
or equivalently,
$\big{\|}\mathcal{T}x-\mathcal{T}y\big{\|}_{\mathcal{Q}}^{2}+\big{\|}(\mathcal{I}-\mathcal{T})x-(\mathcal{I}-\mathcal{T})y\big{\|}_{\mathcal{Q}}^{2}\leq\big{\|}x-y\big{\|}_{\mathcal{Q}}^{2},\quad\forall x,y\in\mathsf{dom}\mathcal{T}.$ (7)
Remark 2.4.
There is still a slight difference between our result and [11, Lemma 2.6]. The latter required $\mathcal{T}$ to be well-defined, which, however, is not assumed in our result. Nevertheless, we clarify that all the expressions in above inequalities are well-defined, even if $\mathcal{T}$ is multi-valued. Indeed, $\|\mathcal{T}x\|_{\mathcal{Q}}=\|\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{T}x\|_{\mathcal{Q}}$ and $\langle\mathcal{T}x|\mathcal{Q}y\rangle=\langle\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{T}x|\mathcal{Q}y\rangle$, $\forall(x,y)\in\mathsf{dom}\mathcal{T}\times\mathcal{H}$. They are well-defined, since $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ is single-valued (by Lemma 2.2). To put it more explicitly, combining with Fact 2, one can equivalently rewrite (7) as $\forall x_{\mathsf{r}},y_{\mathsf{r}}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{dom}\mathcal{T})$:
$\big{\|}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}})-(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(y_{\mathsf{r}})\big{\|}_{\mathcal{Q}}^{2}+\big{\|}(\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}})-(\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(y_{\mathsf{r}})\big{\|}_{\mathcal{Q}}^{2}\leq\big{\|}x_{\mathsf{r}}-y_{\mathsf{r}}\big{\|}_{\mathcal{Q}}^{2}.$ (8)
2.2.3 $\mathcal{Q}$-demiclosedness of $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}$
We first extend the notion of demiclosedness [5, Definition 4.26] and Browder’s demiclosedness principle [5, Theorem 4.27] to any $\mathcal{Q}$-firmly nonexpansive operator $\mathcal{S}$.
Definition 2.5 ($\mathcal{Q}$-demiclosedness).
Under Assumption 1, let $D$ be a nonempty weakly sequentially closed subset of $\mathsf{ran}\mathcal{Q}$, let $\mathcal{S}:D\mapsto\mathsf{ran}\mathcal{Q}$, and let $u\in\mathsf{ran}\mathcal{Q}$. Then $\mathcal{S}$ is $\mathcal{Q}$-demiclosed at $u$ if, for every sequence $\big{\{}x^{k}\big{\}}_{k\in\mathbb{N}}$ in $D$ and every $x\in D$ such that $x^{k}\rightharpoonup x$ and $\mathcal{S}x^{k}\rightarrow u$, we have $\mathcal{S}x=u$. In addition, $\mathcal{S}$ is $\mathcal{Q}$-demiclosed if it is demiclosed at every point in $D$.
Lemma 2.6 ($\mathcal{Q}$-demiclosedness principle).
Under Assumption 1, let $D$ be a nonempty weakly sequentially closed subset of $\mathsf{ran}\mathcal{Q}$ and let $\mathcal{S}:D\mapsto\mathsf{ran}\mathcal{Q}$ be a $\mathcal{Q}$-nonexpansive operator, i.e.,
$\big{\|}\mathcal{S}x_{1}-\mathcal{S}x_{2}\big{\|}_{\mathcal{Q}}\leq\big{\|}x_{1}-x_{2}\big{\|}_{\mathcal{Q}},\quad\forall x_{1},x_{2}\in D\subseteq\mathsf{ran}\mathcal{Q}\cap\mathsf{dom}\mathcal{S}.$
Then $\mathcal{I}-\mathcal{S}$ is $\mathcal{Q}$-demiclosed.
Proof 2.7.
Let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence in $D\subseteq\mathsf{ran}\mathcal{Q}\cap\mathsf{dom}\mathcal{S}$, such that $x^{k}\rightharpoonup x$ and $x^{k}-\mathcal{S}x^{k}\rightarrow u$. Since $D$ is weakly sequentially closed, and $\mathsf{ran}\mathcal{Q}$ is closed (by Assumption 1), we have $x\in D$ and $u\in\mathsf{ran}\mathcal{Q}$. For every $k\in\mathbb{N}$, it follows from the $\mathcal{Q}$-nonexpansiveness of $\mathcal{S}$ that
$\big{\|}x-\mathcal{S}x-u\big{\|}_{\mathcal{Q}}^{2}$ (9)
$=$ $\big{\|}x^{k}-\mathcal{S}x-u\big{\|}_{\mathcal{Q}}^{2}-\big{\|}x^{k}-x\big{\|}_{\mathcal{Q}}^{2}-2\big{\langle}x^{k}-x\big{|}x-\mathcal{S}x-u\big{\rangle}_{\mathcal{Q}}$
$=$ $\big{\|}x^{k}-\mathcal{S}x^{k}-u\big{\|}_{\mathcal{Q}}^{2}+2\big{\langle}x^{k}-\mathcal{S}x^{k}-u\big{|}\mathcal{S}x^{k}-\mathcal{S}x\big{\rangle}_{\mathcal{Q}}$
$+$ $\big{\|}\mathcal{S}x^{k}-\mathcal{S}x\big{\|}_{\mathcal{Q}}^{2}-\big{\|}x^{k}-x\big{\|}_{\mathcal{Q}}^{2}-2\big{\langle}x^{k}-x\big{|}x-\mathcal{S}x-u\big{\rangle}_{\mathcal{Q}}$
$\leq$ $\big{\|}x^{k}-\mathcal{S}x^{k}-u\big{\|}_{\mathcal{Q}}^{2}+2\big{\langle}x^{k}-\mathcal{S}x^{k}-u\big{|}\mathcal{S}x^{k}-\mathcal{S}x\big{\rangle}_{\mathcal{Q}}$
-- $2\big{\langle}x^{k}-x\big{|}x-\mathcal{S}x-u\big{\rangle}_{\mathcal{Q}}.$
By assumption, $x^{k}-\mathcal{S}x^{k}-u\rightarrow 0$ and $x^{k}-x\rightharpoonup 0$ as $k\rightarrow\infty$, and hence
$\mathcal{S}x^{k}-\mathcal{S}x=(\mathcal{S}x^{k}-x^{k})+(x^{k}-x)+x-\mathcal{S}x\rightharpoonup x-\mathcal{S}x-u.$
Taking the limit as $k\rightarrow\infty$ in (9) and appealing to Fact 2, we obtain $\|x-\mathcal{S}x-u\|_{\mathcal{Q}}=0$, which leads to $x-\mathcal{S}x=u$ since $x,\mathcal{S}x,u\in\mathsf{ran}\mathcal{Q}$.
Remark 2.8.
The $\mathcal{Q}$-demiclosedness essentially confines the demiclosed property within the subspace $\mathsf{ran}\mathcal{Q}$, which indicates that the weak and strong convergences take place in $\mathsf{ran}\mathcal{Q}$ only. Actually, if $\mathsf{ran}\mathcal{Q}$ is regarded as the ambient space instead of $\mathcal{H}$, there is no difference between Definition 2.5, Lemma 2.6 and the standard definitions of demiclosedness (cf. [5, Definition 4.26, Theorem 4.27]). However, this restriction makes sense: Lamma 2.9 will conclude $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}$ is $\mathcal{Q}$-demiclosed, but $\mathcal{I}-\mathcal{T}$ is not necessarily so.
Lemma 2.9 ($\mathcal{Q}$-demiclosedness of $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}$).
Given $\mathcal{T}$ as (5) under Assumptions 1 and 2, let $D$ be a nonempty closed subset of $\mathsf{ran}\mathcal{Q}\cap\mathsf{dom}\mathcal{T}$, then $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}:D\mapsto\mathsf{ran}\mathcal{Q}$ is $\mathcal{Q}$-demiclosed.
Proof 2.10.
In view of Lemma 2.6, it suffices to show that $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}:D\mapsto\mathsf{ran}\mathcal{Q}$ is $\mathcal{Q}$-nonexpansive. Indeed, this immediately follows from Fact 3.
2.3 Conjugate view based on generalized Moreau’s decomposition identity
We now, based on the local monotonicity, develop some equivalent forms of $\mathcal{T}$, which also generalize the classic Moreau’s decomposition identity [5, Propositioin 23.20]. The following fact will be used in Lemma 2.12.
Fact 4.
Under Assumptions 1 and 2, the following hold.
(i) $\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}$ is single-valued;
(ii) $(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}$ is single-valued;
(iii) $\sqrt{\mathcal{Q}}\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}$ is single-valued.
Proof 2.11.
(i) Given $x\in\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$, let $\{y_{1},y_{2}\}\subseteq(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x$, then, $y_{1}+\mathcal{A}^{-1}\mathcal{Q}y_{1}\owns x$ and $y_{2}+\mathcal{A}^{-1}\mathcal{Q}y_{2}\owns x$, leading to $y_{2}-y_{1}\in\mathcal{A}^{-1}\mathcal{Q}y_{1}-\mathcal{A}^{-1}\mathcal{Q}y_{2}$. Then,
$0\geq-\|y_{1}-y_{2}\|_{\mathcal{Q}}^{2}=\big{\langle}\mathcal{Q}y_{1}-\mathcal{Q}y_{2}\big{|}y_{2}-y_{1}\big{\rangle}\geq 0,$
where the last inequality is due to $y_{2}-y_{1}\in\mathcal{A}^{-1}\mathcal{Q}y_{1}-\mathcal{A}^{-1}\mathcal{Q}y_{2}$ and the local monotonicity of $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$. Finally, we have $\mathcal{Q}y_{1}=\mathcal{Q}y_{2}$, which shows the single-valuedness of $\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}$.
(ii) Take $x\in\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}\big{)}$. Assume, by contradiction, that $\{y_{1},y_{2}\}\subseteq(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}x$, then, $y_{1}+\mathcal{Q}\mathcal{A}^{-1}y_{1}\owns\mathcal{Q}x$ and $y_{2}+\mathcal{Q}\mathcal{A}^{-1}y_{2}\owns\mathcal{Q}x$. This implies that $\{y_{1},y_{2}\}\subseteq\mathsf{ran}\mathcal{Q}$. Let $y_{1}=\mathcal{Q}v_{1}$ and $y_{2}=\mathcal{Q}v_{2}$, we have $\mathcal{Q}v_{1}+\mathcal{Q}\mathcal{A}^{-1}\mathcal{Q}v_{1}\owns\mathcal{Q}x$ and $\mathcal{Q}v_{2}+\mathcal{Q}\mathcal{A}^{-1}\mathcal{Q}v_{2}\owns\mathcal{Q}x$. Thus, $v_{1}+\mathcal{A}^{-1}\mathcal{Q}v_{1}\owns x+s_{1}$ and $v_{2}+\mathcal{A}^{-1}\mathcal{Q}v_{2}\owns x+s_{2}$, $\forall s_{1},s_{2}\in\mathsf{ker}\mathcal{Q}$. Thus, due to the monotonicity of $\mathcal{A}^{-1}$, we have:
00 0 $\leq$ $\langle\mathcal{Q}v_{1}-\mathcal{Q}v_{2}|t_{1}-t_{2}\rangle\quad\forall t_{i}\in\mathcal{A}^{-1}\mathcal{Q}v_{i}\text{\ for\ }i=1,2$
$=$ $\langle\mathcal{Q}v_{1}-\mathcal{Q}v_{2}|(x+s_{1}-v_{1})-(x+s_{2}-v_{2})\rangle\quad\text{choose\ }t_{i}=x+s_{i}-v_{i}\text{\ for\ }i=1,2$
$=$ $\langle v_{1}-v_{2}|\mathcal{Q}(s_{1}-v_{1}-s_{2}+v_{2})\rangle$
$=$ $-\|v_{1}-v_{2}\|_{\mathcal{Q}}^{2},\quad\text{since $s_{1},s_{2}\in\mathsf{ker}\mathcal{Q}$}$
which leads to $\mathcal{Q}v_{1}=\mathcal{Q}v_{2}$, i.e., $y_{1}=y_{2}$.
(iii) It suffices to show that $\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}$ is single-valued. Take $x\in\mathsf{dom}\big{(}\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}\big{)}$. Assume, by contradiction, that $\{y_{1},y_{2}\}\subseteq\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}x$. Then, $y_{1}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}y_{1}\owns\sqrt{\mathcal{Q}}x$ and $y_{2}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}y_{2}\owns\sqrt{\mathcal{Q}}x$. This implies that $\{y_{1},y_{2}\}\subseteq\mathsf{ran}\mathcal{Q}$. Let $y_{1}=\sqrt{\mathcal{Q}}v_{1}$ and $y_{2}=\sqrt{\mathcal{Q}}v_{2}$, we have $\sqrt{\mathcal{Q}}v_{1}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}v_{1}\owns\sqrt{\mathcal{Q}}x$ and $\sqrt{\mathcal{Q}}v_{2}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}v_{2}\owns\sqrt{\mathcal{Q}}x$. Thus, $v_{1}+\mathcal{A}^{-1}\mathcal{Q}v_{1}\owns x+s_{1}$ and $v_{2}+\mathcal{A}^{-1}\mathcal{Q}v_{2}\owns x+s_{2}$, $\forall s_{1},s_{2}\in\mathsf{ker}\mathcal{Q}$. Thus, due to the monotonicity of $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$, we have:
00 0 $\leq$ $\langle\mathcal{Q}v_{1}-\mathcal{Q}v_{2}|t_{1}-t_{2}\rangle\quad\forall t_{i}\in\mathcal{A}^{-1}\mathcal{Q}v_{i}\text{\ for\ }i=1,2$
$=$ $\langle\mathcal{Q}v_{1}-\mathcal{Q}v_{2}|(x+s_{1}-v_{1})-(x+s_{2}-v_{2})\rangle\quad\text{choose\ }t_{i}=x+s_{i}-v_{i}\text{\ for\ }i=1,2$
$=$ $\langle v_{1}-v_{2}|\mathcal{Q}(s_{1}-v_{1}-s_{2}+v_{2})\rangle$
$=$ $-\|v_{1}-v_{2}\|_{\mathcal{Q}}^{2},\quad\text{since $s_{1},s_{2}\in\mathsf{ker}\mathcal{Q}$}$
which leads to $\mathcal{Q}v_{1}=\mathcal{Q}v_{2}$, i.e., $y_{1}=y_{2}$.
We now extend the classical Moreau’s identity to the degenerate case. We stress that this extension holds under local monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ only. This greatly weakens the previous result of [35, Lemma 2.3].
Lemma 2.12 (Generalized Moreau’s decomposition identity).
Under Assumption 1, we have:
(i) $\mathcal{I}\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}+(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}$ for any input $x\in\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}=\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$;
(ii) Under Assumption 2, it holds that $\mathcal{Q}=\mathcal{Q}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}+\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}$ for any input $x\in\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}=\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$.
Proof 2.13.
(i) For any input $x\in\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}$, pick $y$, such that $y\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$. Then, we develop $\mathcal{Q}x\in\mathcal{A}y+\mathcal{Q}y\Longrightarrow\mathcal{Q}(x-y)\in\mathcal{A}y\Longrightarrow y\in\mathcal{A}^{-1}\mathcal{Q}(x-y)\Longrightarrow x-y\in x-\mathcal{A}^{-1}\mathcal{Q}(x-y)\Longrightarrow x-y\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x\Longrightarrow x\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x+(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$, from which follows that $\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}\subseteq\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$.
Conversely, For any input $x\in\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$, pick $y$, such that $y\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x$. Then, we develop $y+\mathcal{A}^{-1}\mathcal{Q}y\owns x\Longrightarrow x-y\in\mathcal{A}^{-1}\mathcal{Q}y\Longrightarrow\mathcal{Q}y\in\mathcal{A}(x-y)\Longrightarrow\mathcal{Q}(y-x+x)\in\mathcal{A}(x-y)\Longrightarrow\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})(x-y)\Longrightarrow x-y\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x\Longrightarrow x\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x+(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$, from which follows that $\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}\supseteq\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$.
Both directions show (i), with $\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}=\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$.
(ii) Applying $\mathcal{Q}$ to both sides of the first inclusion (i), we obtain
$\mathcal{Q}\in\mathcal{Q}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}+\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}.$ (10)
The single-valuedness of $\mathcal{Q}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}$ and $\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}$ has been shown in Lemma 2.2 and Fact 4-(i) [^2] Notice that here Assumption 2 is required for single-valuedness.. Finally, $\in$’ in (10) can be properly replaced by ‘$=$’.
Several useful equalities are further developed below, which provide more alternatives to Lemma 2.12.
Lemma 2.14.
Under Assumptions 1 and 2, the following equalities hold:
$\mathcal{Q}-\mathcal{Q}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}$
[…]
where $\sqrt{\mathcal{Q}}$ is defined as Fact 1-(iii).
Proof 2.15.
First, note that the single-valuedness of all the expressions here has been shown in Fact 4. We show the equalities following the direction of $(1)\Longrightarrow(2)\Longrightarrow(3)\Longrightarrow(4)\Longrightarrow(1)$.
(1): Lemma 2.12-(ii).
(2): We begin with $\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}$. Take $x\in\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$. Let $v\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x$ and $y=\mathcal{Q}v$. Then, we deduce that $v+\mathcal{A}^{-1}\mathcal{Q}v\owns x\Longrightarrow\mathcal{Q}v+\mathcal{Q}\mathcal{A}^{-1}\mathcal{Q}v\owns\mathcal{Q}x\Longrightarrow y+\mathcal{Q}\mathcal{A}^{-1}y\owns\mathcal{Q}x\Longrightarrow y=(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}x$. This also shows that $\mathsf{dom}\big{(}\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}\subseteq\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}\big{)}$.
(3): We begin with $(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}$. Take $x\in\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}\big{)}$. Let $y=(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}x$, then, $y+\mathcal{Q}\mathcal{A}^{-1}y\owns\mathcal{Q}x$. This implies that $y\in\mathsf{ran}\mathcal{Q}$. Let $y=\sqrt{\mathcal{Q}}v$ with $v\in\mathsf{ran}\mathcal{Q}$ [^3] Strictly speaking, we should write $y=\sqrt{\mathcal{Q}}(v+s)$ with $s\in\mathsf{ker}\mathcal{Q}$. This is essentially equivalent to $y=\sqrt{\mathcal{Q}}v$.. Then, the inclusion becomes $\sqrt{\mathcal{Q}}v+\mathcal{Q}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}v\owns\mathcal{Q}x\Longrightarrow v+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}v\owns\sqrt{\mathcal{Q}}x\Longrightarrow v=\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}x\Longrightarrow y=\sqrt{\mathcal{Q}}v=\sqrt{\mathcal{Q}}\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}x$. This also shows that $\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}\big{)}\subseteq\mathsf{dom}\big{(}\sqrt{\mathcal{Q}}\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}\big{)}$.
(4): It suffices to show $\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}=\Big{(}\mathcal{I}+\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}^{-1}\sqrt{\mathcal{Q}}$. Take $x\in\mathsf{dom}\big{(}\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}\big{)}$. Let $y=\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}x$, then, $y+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}y\owns\sqrt{\mathcal{Q}}x$. This implies that $y\in\mathsf{ran}\mathcal{Q}$, which already satisfies the domain restricted in $\mathsf{ran}\mathcal{Q}$, i.e., $y+\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}|_{\mathsf{ran}\mathcal{Q}}(y)\owns\sqrt{\mathcal{Q}}x$, leading to the right-hand side of (4).
(4$\Longrightarrow$(1): It suffices to show that $\Big{(}\mathcal{I}+\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}^{-1}\sqrt{\mathcal{Q}}=\sqrt{\mathcal{Q}}-\sqrt{\mathcal{Q}}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}$. Take $x\in\mathsf{dom}\big{(}\big{(}\mathcal{I}+(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}})|_{\mathsf{ran}\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}\big{)}$. Let $y=\big{(}\mathcal{I}+(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}})|_{\mathsf{ran}\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}x$, then, $y+(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}})|_{\mathsf{ran}\mathcal{Q}}(y)\owns\sqrt{\mathcal{Q}}x$. This implies that $y\in\mathsf{ran}\mathcal{Q}$, and thus, the above inclusion can be simplified as $y+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}y\owns\sqrt{\mathcal{Q}}x$, and let $y=\sqrt{\mathcal{Q}}v$ with $v\in\mathsf{ran}\mathcal{Q}$. Then, we have $\sqrt{\mathcal{Q}}v+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}v\owns\sqrt{\mathcal{Q}}x\Longrightarrow v+\mathcal{A}^{-1}\mathcal{Q}v\owns x\Longrightarrow\mathcal{A}^{-1}\mathcal{Q}v\owns x-v\Longrightarrow\mathcal{Q}v\in\mathcal{A}(x-v)\Longrightarrow\mathcal{Q}(v-x+x)\in\mathcal{A}(x-v)\Longrightarrow\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})(x-v)\Longrightarrow x-v\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x\Longrightarrow v\in x-(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x\Longrightarrow y=\sqrt{\mathcal{Q}}v=\sqrt{\mathcal{Q}}(\mathcal{I}-(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q})x$. This also shows that $\mathsf{dom}\big{(}\big{(}\mathcal{I}+(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}})|_{\mathsf{ran}\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}\big{)}\subseteq\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}$.
Finally, we show all the equalities with exactly same domains.
Remark 2.16.
For non-degenerate $\mathcal{Q}$, Lemma 2.14-(2) can be simplified as
$\mathcal{I}$ $=$ $\mathcal{Q}(\mathcal{A}+\mathcal{Q})^{-1}+(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}$ (11)
$=$ $(\mathcal{A}^{-1}+\mathcal{Q}^{-1})^{-1}\mathcal{Q}^{-1}+\mathcal{Q}(\mathcal{A}+\mathcal{Q})^{-1}.$
This is exactly [8, Corollary 7.2]. However, (11) is no longer valid for degenerate case, since it is not guaranteed that $\mathsf{ran}(\mathcal{A}+\mathcal{Q})=\mathcal{H}$. In fact, we will see in Sect. 3.3.1 that $\mathcal{A}+\mathcal{Q}$ is often non-surjective in many examples.
3 Full domain
It is now ready to discuss the full domain issue. Since it is not easy to verify $\mathsf{ran}(\mathcal{A}+\mathcal{Q})\supseteq\mathsf{ran}\mathcal{Q}$, we aim to propose several alternatives to this condition by exploring the interaction between the monotonicity of $\mathcal{A}$ with the degeneracy of $\mathcal{Q}$. More specifically, we wish to develop an analogous result to the well-known Minty’s Theorem [5, Theorem 21.1] for the degenerate case.
3.1 Restricted maximality: Motivations and definitions
Definition 2.1 defines the local monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$. The well-known Minty’s theorem inspires us that the full domain of $\mathcal{T}$ should be related to somewhat ‘maximality’ of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$. The general concept of local maximal monotonicity has been given in [5, Definition 25.7]. However, it is not easy to properly incorporate $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ with ‘maximality’ as [5, Definition 25.7]. Indeed, suppose we enforce the local maximality as follows:
[A tentative definition] $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ is called ‘local maximally monotone’, if for any given $(x,u)\in\mathcal{H}\times\mathsf{ran}\mathcal{Q}$, we have:
$(x,u)\in\mathsf{gra}\mathcal{A}\Longleftrightarrow\forall(y,v)\in\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q}),\ \langle x-y|u-v\rangle\geq 0.$ (12)
In other words, for every $(x,u)\in(\mathcal{H}\times\mathsf{ran}\mathcal{Q})\backslash\mathsf{gra}\mathcal{A}$, there exists $(y,v)\in(\mathcal{H}\times\mathsf{ran}\mathcal{Q})\cap\mathsf{gra}\mathcal{A}$ such that $\langle x-y|u-v\rangle<0$.
Let us check if the definition of (12) is appropriate, under the full domain condition of $\mathcal{T}$. Following a very similar manner with the proof of Minty’s Theorem [5, Theorem 21.1], we fix $(x,u)\in\mathcal{H}\times\mathsf{ran}\mathcal{Q}$, such that
$\forall(y,v)\in\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q}),\quad\langle x-y|u-v\rangle\geq 0.$ (13)
What we need to prove for the maximality is $u\in\mathcal{A}x\cap\mathsf{ran}\mathcal{Q}$. First, it is noticed that the above inner product in (12) is actually only performed in $\mathsf{ran}\mathcal{Q}$, i.e., $\langle x-y|u-v\rangle=\langle x_{\mathsf{r}}-y_{\mathsf{r}}|u-v\rangle\geq 0$, since $u,v\in\mathsf{ran}\mathcal{Q}$, and the kernel parts $x_{\mathsf{k}}$ and $y_{\mathsf{k}}$ are neglected and not properly controlled by the inner product. Since $\mathsf{ran}(\mathcal{A}+\mathcal{Q})\supseteq\mathsf{ran}\mathcal{Q}$, then $\exists y\in\mathcal{H}$, such that
$v\in\mathcal{A}y\cap\mathsf{ran}\mathcal{Q}\textrm{\ and \ }\mathcal{Q}x+u=v+\mathcal{Q}y\in(\mathcal{A}+\mathcal{Q})y.$
Then, it follows that $u-v=\mathcal{Q}(y-x)$. By substituting into (13), we obtain $0\leq\langle x-y|u-v\rangle=-\|x-y\|_{\mathcal{Q}}^{2}$, resulting in $\mathcal{Q}x=\mathcal{Q}y$, $x_{\mathsf{r}}=y_{\mathsf{r}}$ and thus, $u=v$. Then, we obtain $u\in\mathcal{A}y\cap\mathsf{ran}\mathcal{Q}$ with $y_{\mathsf{r}}=x_{\mathsf{r}}$, but fail to reach the desired result of $u\in\mathcal{A}x\cap\mathsf{ran}\mathcal{Q}$. In other words, this shows that under the full domain condition of $\mathcal{T}$, $\exists(x,u)\in(\mathcal{H}\times\mathsf{ran}\mathcal{Q})\backslash\mathsf{gra}\mathcal{A}$, such that $\langle x-y|u-v\rangle\geq 0$, $\forall(y,v)\in(\mathcal{H}\times\mathsf{ran}\mathcal{Q})\cap\mathsf{gra}\mathcal{A}$.
Therefore, the above reasoning shows that the full domain condition fails to yield the local maximality of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$, as (12) suggests. The problem is that $\langle\cdot|\cdot\rangle_{\mathcal{Q}}$ is not a valid inner product in $\mathcal{H}$, and we can only obtain $x_{\mathsf{r}}=y_{\mathsf{r}}$, but not $x=y$ in the whole space $\mathcal{H}$. This inspires us to define the maximality restricted to $\mathsf{ran}\mathcal{Q}$ in both domain and range of $\mathcal{A}$, in order to remove the redundancy of the inner product of (12). More precisely, considering $v\in\mathcal{A}y\cap\mathsf{ran}\mathcal{Q}$, we also wish to restrict $y$—the domain of $\mathcal{A}$—in $\mathsf{ran}\mathcal{Q}$ as well, and discard the kernel component $y_{\mathsf{k}}$. It is much easier to finish this work based on $\mathcal{A}^{-1}$ rather than $\mathcal{A}$ according to the following derivation: $v\in\mathcal{A}y\cap\mathsf{ran}\mathcal{Q}\Longleftrightarrow v\in\mathcal{A}y{\rm\ and\ }v\in\mathsf{ran}\mathcal{Q}\Longleftrightarrow y\in\mathcal{A}^{-1}v{\rm\ and\ }v\in\mathsf{ran}\mathcal{Q}\Longleftrightarrow y_{\mathsf{r}}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}v{\rm\ and\ }v\in\mathsf{ran}\mathcal{Q}$. This gives us a new operator—$\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$. To proceed with Fact 1, we develop the following equivalence.
Fact 5.
$\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone, if and only if $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone.
Proof 3.1.
If $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone, then we have:
$\langle u-v|x-y\rangle\geq 0,\ \forall(u,x)\in\mathsf{gra}(\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}),\ \forall(v,y)\in\mathsf{gra}(\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}).$
Note that $\langle u-v|x-y\rangle=\langle u-v|x_{\mathsf{r}}-y_{\mathsf{r}}\rangle$, since $u,v\in\mathsf{ran}\mathcal{Q}$. This is equivalent to saying that
$\langle u-v|x_{\mathsf{r}}-y_{\mathsf{r}}\rangle\geq 0,\ \forall(u,x_{\mathsf{r}})\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}),\ \forall(v,y_{\mathsf{r}})\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}).$
This shows that $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone.
Conversely, if $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone, then we have:
$\langle u-v|\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}y\rangle\geq 0,\ \forall(u,x)\in\mathsf{gra}(\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}),\ \forall(v,y)\in\mathsf{gra}(\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}).$
Note that $\langle u-v|\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}y\rangle=\langle\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(u-v)|x-y\rangle=\langle u-v|x-y\rangle\geq 0$, since $u,v\in\mathsf{ran}\mathcal{Q}$. This shows that $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone.
A key principle behind Fact 5 is that the monotonicity is defined by an inner product, which inherently pass the restriction of one side of inner product to another.
As summary, Facts 1 and 5 establish
Lemma 3.2.
The following 3 statements are equivalent.
(i) $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ is monotone;
(ii) $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone;
(iii) $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone.
The main advantage of defining $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ over other two options is that $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ restricts both domain and range to $\mathsf{ran}\mathcal{Q}$, and thus (hopefully) can be properly assigned with the notion of ‘restricted maximality’. Here, we distinguish ‘restricted’ from ‘local’: local means the maximality is associated with some open neighborhood $C\subset\mathcal{H}$, whereas restricted means the maximality is with respect to a subspace, which is not an open subset.
Definition 3.3.
$\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, if for any given $(u,x)\in\mathsf{ran}\mathcal{Q}\times\mathsf{ran}\mathcal{Q}$, we have:
$(u,x)\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}})\Longleftrightarrow\forall(v,y)\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}),\ \langle u-v|x-y\rangle\geq 0.$
Observing Definition 3.3, if the subspace $\mathsf{ran}\mathcal{Q}$ is regarded as the ambient space, $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ obeys a standard definition of maximal monotone operator. Thus, it immediately follows from classic Minty’s theorem [5, Theorem 21.1] that
Theorem 3.4 (Restricted Minty’s theorem).
Let $\mathcal{B}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ be monotone. Then $\mathcal{B}$ is maximally monotone, if and only if $\mathsf{ran}(\mathcal{I}+\mathcal{B})=\mathsf{ran}\mathcal{Q}$.
Then, replacing $\mathcal{B}$ by our specific $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$, we obtain
Corollary 3.5.
Let $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ be monotone. Then $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, if and only if $\mathsf{ran}(\mathcal{I}+\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}})=\mathsf{ran}\mathcal{Q}$.
3.2 Various conditions for full domain
This part will show how Corollary 3.5 is connected to the full domain condition of $\mathcal{T}$. Lemma 3.6 is a bridge for this connection.
Lemma 3.6.
Under Assumptions 1 and 2, the following two statements are equivalent.
(i) $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone;
(ii) $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone.
Proof 3.7.
(i)$\Longrightarrow$(ii): If $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, then for any given $(u,x)\in\mathsf{ran}\mathcal{Q}\times\mathsf{ran}\mathcal{Q}$, we have:
$(u,x)\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}})\Longleftrightarrow\forall(v,y)\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}),\ \langle u-v|x-y\rangle\geq 0.$
Since $u,x,v,y\in\mathsf{ran}\mathcal{Q}$, let $u=\sqrt{\mathcal{Q}}u^{\prime}$ and $v=\sqrt{\mathcal{Q}}v^{\prime}$ for $u^{\prime},v^{\prime}\in\mathsf{ran}\mathcal{Q}$. Then, $x\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}(u)=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}u=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}u^{\prime}$, and similarly, $y\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}v^{\prime}$. Then, we obtain
$(u^{\prime},x)\in\mathsf{gra}\Big{(}\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}\Longleftrightarrow\forall(v^{\prime},y)\in\mathsf{gra}\Big{(}\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)},\ \langle\sqrt{\mathcal{Q}}(u^{\prime}-v^{\prime})|x-y\rangle\geq 0.$
Let $x^{\prime}=\sqrt{\mathcal{Q}}x$ and $y^{\prime}=\sqrt{\mathcal{Q}}y$, then it becomes
$(u^{\prime},x^{\prime})\in\mathsf{gra}\Big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}\Longleftrightarrow\forall(v^{\prime},y^{\prime})\in\mathsf{gra}\Big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)},\ \langle u^{\prime}-v^{\prime}|x^{\prime}-y^{\prime}\rangle\geq 0.$
This shows that $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone.
(ii)$\Longrightarrow$(i): If $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, then for any given $(u,x)\in\mathsf{ran}\mathcal{Q}\times\mathsf{ran}\mathcal{Q}$, we have:
$(u,x)\in\mathsf{gra}(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}})\Longleftrightarrow\forall(v,y)\in\mathsf{gra}(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}),\ \langle u-v|\sqrt{\mathcal{Q}}x-\sqrt{\mathcal{Q}}y\rangle\geq 0.$
Let $x^{\prime}=\sqrt{\mathcal{Q}}x$, $y^{\prime}=\sqrt{\mathcal{Q}}y$, $u^{\prime}=\sqrt{\mathcal{Q}}u$, $v^{\prime}=\sqrt{\mathcal{Q}}v$, we have:
$(u^{\prime},x^{\prime})\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}})\Longleftrightarrow\forall(v^{\prime},y^{\prime})\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}),\ \langle u^{\prime}-v^{\prime}|x^{\prime}-y^{\prime}\rangle\geq 0.$
This shows the maximal monotonicity of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$.
Then, we obtain a sufficient and necessary condition for the full domain of $\mathcal{T}$.
Theorem 3.8.
$\mathcal{T}$ given as (5) has full domain, if and only if $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ (or equivalently saying, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$) $:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ is maximally monotone.
Proof 3.9.
Combine Lemma 2.14-(4), Corollary 3.5 and Lemma 3.6.
The next question naturally arises: what is the condition on $\mathcal{A}$ and $\mathcal{Q}$, for which $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ is maximally monotone? The following important corollary gives a sufficient condition for maximal monotonicity of $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$.
Corollary 3.10.
Under Assumptions 1 and 2, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, if the following conditions are fulfilled:
(i) $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is maximally monotone;
(ii) $0\in\mathsf{sri}(\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A})$ [^4] The definition of strong relative interior $(\mathsf{sri})$ is given by [5, Definition 6.9]..
In particular, if $\mathcal{Q}$ is non-degenerate, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, if and only if $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is maximally monotone.
Proof 3.11.
In view of [5, Corollary 25.6 or Proposition 25.41-(iv)].
Finally, as summary, under Assumptions 1 and 2, the full domain condition of $\mathcal{T}$ can be equivalently (i.e., sufficient and necessary) expressed as
Assumption 3
$\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ or $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ is maximally monotone;
or further improved as the following sufficient condition in many practical situations:
Assumption 4
(i) $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is maximally monotone;
(ii) $0\in\mathsf{sri}(\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A})$.
Remark 3.12.
[…]
$\mathcal{A}=\begin{bmatrix}\partial f&0\\ -B&\partial g\end{bmatrix}:\mathcal{X}\times\mathcal{Y}\mapsto 2^{\mathcal{X}}\times 2^{\mathcal{Y}},\quad\mathcal{Q}=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}:\mathcal{X}\times\mathcal{Y}\mapsto\mathcal{X}\times\{0\},$
[…]
3.3 Further discussions
3.3.1 Several examples
We here list several examples to show
(1) $\mathcal{A}+\mathcal{Q}$ is often non-surjective;
(2) $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ may not be maximally monotone for degenerate $\mathcal{Q}$, even if $\mathcal{A}$ is maximally monotone;
(3) Assumption 4 is tight, but not necessary for the full domain of $\mathcal{T}$.
Example 3.13.
Consider a toy example
$\mathcal{H}=\mathbb{R}^{2},\quad\mathcal{A}=\partial f,\text{\ where\ }f(x,y)=|x|,\quad\mathcal{Q}=\begin{bmatrix}0&0\\ 0&1\end{bmatrix}.$
Then, it is easy to see that $\mathsf{ran}\mathcal{A}=[-1,1]\times\{0\}$, $\mathsf{ran}\mathcal{Q}=\{0\}\times\mathbb{R}$, $\mathsf{ran}(\mathcal{A}+\mathcal{Q})=[-1,1]\times\mathbb{R}\neq\mathcal{H}=\mathbb{R}^{2}$, but it already suffices to guarantee that $\mathsf{ran}(\mathcal{A}+\mathcal{Q})\supset\mathsf{ran}\mathcal{Q}$. One further have $\mathsf{ran}\mathcal{A}-\mathsf{ran}\mathcal{Q}=[-1,1]\times\mathbb{R}$, which contains 0 as a relative interior.
Is $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ maximally monotone? First, noting that $\mathsf{dom}\mathcal{A}^{-1}=\mathsf{ran}\mathcal{A}=[-1,1]\times\{0\}$, $\mathcal{A}^{-1}$ is given as:
$\mathcal{A}^{-1}:(x,y)\mapsto\left\{\begin{array}[]{ll}\mathbb{R}_{+}\times\mathbb{R},&\textrm{if\ }x=1,\ y=0;\\ \mathbb{R}_{-}\times\mathbb{R},&\textrm{if\ }x=-1,\ y=0;\\ \{0\}\times\mathbb{R},&\textrm{if\ }x\in(-1,1),\ y=0.\\ \end{array}\right.$
where +=$\mathbb{R}_{+}=\{x\in\mathbb{R}:x\geq 0\}$, =$\mathbb{R}_{-}=\{x\in\mathbb{R}:x\leq 0\}$.
We now check whether $\mathsf{ran}(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})=\{0\}\times\mathcal{R}=\mathsf{ran}\mathcal{Q}$. Indeed, we develop:
$\mathsf{ran}\Big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}$ $\supseteq$ $\Big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}\begin{bmatrix}0\\ 0\end{bmatrix}$
$=$ $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\begin{bmatrix}0\\ 0\end{bmatrix}=\sqrt{\mathcal{Q}}(\{0\}\times\mathbb{R})$
$=$ $\{0\}\times\mathbb{R}=\mathsf{ran}\mathcal{Q},$
which shows $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone by restricted Minty’s theorem (Theorem 3.4 or Corollary 3.5).
Example 3.14.
In [11, Remark 2.2], the authors proposed an interesting counter-example that deserves particular treatment. The problem is
$\mathcal{H}=\mathbb{R}^{2},\mathcal{A}=\partial f,\text{\ where\ }f(x,y)=\max\{e^{y}-x,0\}\text{\ and\ }\mathcal{Q}=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}.$
Based on basic subdifferential calculus, one obtains $\mathcal{A}$ given as
$\mathcal{A}:(x,y)\mapsto\left\{\begin{array}[]{ll}(-1,e^{y}),&\textrm{if\ }y>\ln x,x>0\textrm{\ or\ }x\leq 0,\forall y\in\mathbb{R};\\ (0,0),&\textrm{if\ }y<\ln x,x>0;\\ {}[-1,0]\times[0,e^{y}],&\textrm{if\ }y=\ln x,x>0.\end{array}\right.$
Thus, $\mathcal{A}+\mathcal{Q}$ is given as
$\mathcal{A}+\mathcal{Q}:(x,y)\mapsto\left\{\begin{array}[]{ll}(x-1,e^{y}),&\textrm{if\ }y>\ln x,x>0\textrm{\ or\ }x\leq 0,\forall y\in\mathbb{R};\\ (x,0),&\textrm{if\ }y<\ln x,x>0;\\ (u,v)\in[x-1,x]\times[0,e^{y}],&\textrm{if\ }y=\ln x,x>0.\end{array}\right.$
We then deduce that $\mathsf{ran}\mathcal{A}=[-1,0]\times\mathbb{R}_{+}$, $\mathsf{ran}(\mathcal{A}+\mathcal{Q})=(-\infty,-1]\times(0,+\infty)\cup(-1,+\infty]\times[0,+\infty)=\mathbb{R}\times[0,+\infty)\backslash(-\infty,-1]\times\{0\}$. Clearly, $\mathsf{ran}(\mathcal{A}+\mathcal{Q})\nsupseteq\mathsf{ran}\mathcal{Q}=\mathbb{R}\times\{0\}$. Note that $(\mathcal{A}+\mathcal{Q})^{-1}(x,0)=\varnothing$ for $x\leq-1$, though $(x,0)\in\mathsf{ran}\mathcal{Q}$.
Since $\mathcal{T}$ is not of full domian, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ must not be maximally monotone by Theorem 3.8. Let us check it. First, we have $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}:(x,0)\mapsto\mathbb{R}_{++}\times\{0\}$, $\forall x\in[-1,0]$, where ++=$\mathbb{R}_{++}=\{x\in\mathbb{R}:x>0\}$. Then, $\mathsf{ran}(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})=(-1,+\infty)\times\{0\}\neq\mathbb{R}\times\{0\}$, and $(-\infty,-1]\nsubseteq\mathsf{ran}(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})$. Thus, by restricted Minty’s theorem, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is not maximally monotone.
Finally, it is easy to check that Assumption 4-(ii) is not satisfied. Indeed, $\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A}=\mathbb{R}\times(-\infty,0]$, and thus, $0\notin\mathsf{sri}(\mathsf{ran}\mathcal{A}-\mathsf{ran}\mathcal{Q})$. This shows that Assumption 4 is tight.
Example 3.15.
We consider Example 3.14 but with the metric $\mathcal{Q}=\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$. Then $\mathcal{A}+\mathcal{Q}$ is given as
$\mathcal{A}+\mathcal{Q}:(x,y)\mapsto\left\{\begin{array}[]{ll}(1,e^{y}+y),&\textrm{if\ }y>\ln x,x>0\textrm{\ or\ }x\leq 0,\forall y\in\mathbb{R};\\ (0,y),&\textrm{if\ }y<\ln x,x>0;\\ {}[-1,0]\times[y,e^{y}+y],&\textrm{if\ }y=\ln x,x>0.\end{array}\right.$
We then deduce that $\mathsf{ran}(\mathcal{A}+\mathcal{Q})=[-1,0]\times\mathbb{R}$. Clearly, $\mathsf{ran}(\mathcal{A}+\mathcal{Q})\supseteq\mathsf{ran}\mathcal{Q}=\{0\}\times\mathbb{R}$. By Theorem 3.8, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ must be maximally monotone. Indeed, if the input $y=0$, we then have $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}:0\mapsto\{0\}\times\mathbb{R}$. Then, $\mathsf{ran}(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})\supseteq(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})(0)=(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})(0)=\{0\}\times\mathbb{R}=\mathsf{ran}\mathcal{Q}$, and thus by restricted Minty’s theorem, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone.
However, $\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A}=[0,1]\times\mathbb{R}$, $0\notin\mathsf{sri}(\mathsf{ran}\mathcal{A}-\mathsf{ran}\mathcal{Q})$. This shows that the condition of $0\in\mathsf{sri}(\mathsf{ran}\mathcal{A}-\mathsf{ran}\mathcal{Q})$ is sufficient but not necessary condition for the full domain of $\mathcal{T}$.
3.3.2 Other related results
We now have additional remarks based on some known results in [5, Chap. 25].
Fact 6.
Under Assumptions 1 and 4-(i), the following hold.
(i) $\mathcal{Q}$ is maximally monotone;
(ii) $\mathcal{A}+\mathcal{Q}$ is maximally monotone;
(iii) $\mathcal{Q}$ is $3^{*}$ monotone.
Proof 3.16.
(i) [5, Corollary 20.28];
(ii) [5, Corollary 25.5];
(iii) Fact 1-(iv) and [5, Example 25.17].
Then, it is easy to conclude
Proposition 3.17.
Under Assumptions 1 and 4-(i), $\overline{\mathsf{ran}}(\mathcal{A}+\mathcal{Q})\supseteq\mathsf{ran}\mathcal{Q}$.
Proof 3.18.
By Fact 6, and $\mathsf{dom}\mathcal{A}\subseteq\mathsf{dom}\mathcal{Q}=\mathcal{H}$, then [5, Theorem 25.24] yields
$\overline{\mathsf{ran}}(\mathcal{A}+\mathcal{Q})=\overline{\mathsf{ran}\mathcal{A}+\mathsf{ran}\mathcal{Q}}\supseteq\overline{0+\mathsf{ran}\mathcal{Q}}=\overline{\mathsf{ran}}\mathcal{Q}=\mathsf{ran}\mathcal{Q},$
where the last equality is due to the closedness of $\mathsf{ran}\mathcal{Q}$.
The above result always holds, without Assumption 4-(ii). One can check Example 3.14, where $\mathsf{ran}(\mathcal{A}+\mathcal{Q})\nsupseteq\mathsf{ran}\mathcal{Q}$ but $\overline{\mathsf{ran}}(\mathcal{A}+\mathcal{Q})=\mathbb{R}\times\mathbb{R}_{+}\supset\mathbb{R}\times\{0\}=\mathsf{ran}\mathcal{Q}$. Unfortunately, it is generally difficult to investigate the relation between $\overline{\mathsf{ran}}(\mathcal{A}+\mathcal{Q})$ and $\mathsf{ran}(\mathcal{A}+\mathcal{Q})$. We believe there are some connections between Assumption 4-(ii) and $\overline{\mathsf{ran}}(\mathcal{A}+\mathcal{Q})$. This is left to future work.
4 Single-valuedness
Assumption 3 (or 4) has ensured the full domain of $\mathcal{T}$. Let us now study the condition, under which $\mathcal{T}$ is single-valued, not only in its projection onto $\mathsf{ran}\mathcal{Q}$.
4.1 Injectivity: A primal view
As mentioned in Sect. 1.2, a standard condition for the single-valuedness of $\mathcal{T}$ is disjoint injectivity of $\mathcal{A}+\mathcal{Q}$ [6, Theorem 2.1-(ix)], [14, Proposition 3.8]. [13, Proposition 2.1-(2)] further proposed a sufficient and necessary condition is that $\mathcal{A}+\mathcal{Q}$ is injective on $\mathsf{ran}\mathcal{Q}$. Proposition 4.2 further relaxes this condition to its disjoint injectivity on $\mathsf{ran}\mathcal{Q}$ w.r.t. $\mathsf{ker}\mathcal{Q}$ only, which is defined as below.
Definition 4.1.
$\mathcal{B}$ is disjointly injective on $\mathsf{ran}\mathcal{Q}$ w.r.t. $\mathsf{ker}\mathcal{Q}$, if
$\mathcal{B}y_{1}\cap\mathcal{B}y_{2}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing\Longrightarrow y_{1,\mathsf{k}}=y_{2,\mathsf{k}},\quad\forall y_{1},y_{2}\in\mathsf{dom}\mathcal{A}.$ (14)
Proposition 4.2.
Under Assumptions 1, 2 and 3 (or 4), $\mathcal{T}$ is single-valued, if and only if $\mathcal{A}+\mathcal{Q}$ is disjointly injective on $\mathsf{ran}\mathcal{Q}$ w.r.t. $\mathsf{ker}\mathcal{Q}$.
Proof 4.3.
(1) Assuming (14) holds, take $x$, such that $\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}$, then, (14) yields $y_{1,\mathsf{k}}=y_{2,\mathsf{k}}$. On the other hand, from $\{y_{1},y_{2}\}\subseteq(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$ and by Lemma 2.2, we obtain that $y_{1,\mathsf{r}}=y_{2,\mathsf{r}}$. Combining both results leads to $y_{1}=y_{2}$.
(2) Assuming $\mathcal{T}$ is single-valued, then we need to show (14). Let us prove by contradiction. If (14) does not hold, it implies that $\exists y_{1},y_{2}$, such that
$(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing\Longrightarrow y_{1,\mathsf{k}}\neq y_{2,\mathsf{k}}.$
Take $x$, such that $\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}$. It has been shown in Lemma 2.2 that $y_{1,\mathsf{r}}=y_{2,\mathsf{r}}$. Combining with $y_{1,\mathsf{k}}\neq y_{2,\mathsf{k}}$, we obtain $y_{1}\neq y_{2}$. This means that $\mathcal{T}$ is not single-valued, which is contradiction.
4.2 Our result: A dual view
Let us now study the single-valued issue from a conjugate perspective.
Theorem 4.4.
Under Assumptions 1, 2 and 3 (or 4), $y$ is uniquely determined by (5) for any given $x\in\mathcal{H}$, if and only if one of the following equivalent conditions holds:
(i) $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ is single-valued;
(ii) For any input $x_{\mathsf{r}}\in\mathsf{ran}\mathcal{Q}$, given the corresponding $y_{\mathsf{r}}=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}})$, the constrained inclusion problem of $\mathcal{A}(y_{\mathsf{r}}+y_{\mathsf{k}})\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$, s.t. $\mathcal{Q}y_{\mathsf{k}}=0$, has a unique solution of $y_{\mathsf{k}}$.
(iii) Suppose there exists a linear, bounded and surjective operator $\mathcal{E}:\mathcal{H}\mapsto D^{\prime}$, for some real Hilbert space $D^{\prime}$, such that $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}=\mathcal{E}^{\top}\mathcal{E}$. Then, for any input $x_{\mathsf{r}}\in\mathsf{ran}\mathcal{Q}$ given the corresponding $y_{\mathsf{r}}=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}})$, the inclusion problem $\mathcal{A}(y_{\mathsf{r}}+\mathcal{E}^{\top}b)\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$ has a unique solution of $b\in D^{\prime}$.
Proof 4.5.
(i)$\Longleftrightarrow$(ii): From (5), we deduce
$(\mathcal{A}+\mathcal{Q})y\owns\mathcal{Q}x\Longrightarrow y\in\mathcal{A}^{-1}\mathcal{Q}(x-y)\Longrightarrow y\in\mathcal{A}^{-1}\mathcal{Q}(x-y)=\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}}),$
which results in
$y_{\mathsf{r}}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}}),\quad y_{\mathsf{k}}\in\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}}).$ (15)
Lemma 2.2 shows that $y_{\mathsf{r}}$ is uniquely determined by $x_{\mathsf{r}}$. Thus, the single-valuedness of $\mathcal{T}$ totally depends on the kernel part—$y_{\mathsf{k}}$, as given above. This completes the proof, noting that $x_{\mathsf{r}}$ and $y_{\mathsf{r}}$ are irrelevant to $y_{\mathsf{k}}$.
(ii)$\Longleftrightarrow$(iii): clear.
Remark 4.6.
(i) There is no need to enforce $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ to be single-valued, since the uniqueness of $y_{\mathsf{r}}$ has been guaranteed by Lemma 2.2 (i.e., Assumptions 1 and 2), and is irrelevant to single-valuedness of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$, which is, indeed, often multi-valued in practice.
(ii) The condition (i) is succinct; (ii) is more practical, since since $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ is sometimes difficult to evaluate (see Sect. 6). A big merit of condition (iii) is that the unconstrained inclusion is easier to solve.
Anyway, Theorem 4.4 gives rise to
Assumption 5
$\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ is single-valued.
We present the following result as summary for the well-definedness of $\mathcal{T}$.
Theorem 4.7.
Under Assumptions 1 and 2, $\mathcal{T}$ given as (5) is well-defined,
(i) if and only if Assumption 3 and 5;
(ii) if Assumption 4 and 5.
4.3 Consistency between primal and dual views
We now prove that Proposition 4.2 (a primal view) and Theorem 4.4 (a conjugate view) are essentially equivalent.
Proposition 4.8.
Under Assumptions 1, 2 and 3 (or 4), the following statements are equivalent.
(i) $\mathcal{A}+\mathcal{Q}$ is disjointly injective on $\mathsf{ran}\mathcal{Q}$ w.r.t. $\mathsf{ker}\mathcal{Q}$;
(ii) $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is single-valued.
Proof 4.9.
(i)$\Longrightarrow$(ii): Since $(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing$, pick a point $x$, such that
$\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}.$ (16)
Lemma 2.2 and (14) show that $y_{1,\mathsf{r}}=y_{2,\mathsf{r}}:=y_{\mathsf{r}}$ and $y_{1,\mathsf{k}}=y_{2,\mathsf{k}}:=y_{\mathsf{k}}$. On the other hand, (16) can be rewritten as (15), which shows the single-valuedness of $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$.
(ii)$\Longrightarrow$(i): We aim to prove (14). Let us prove it by contradiction. Assuming (14) does not hold, it implies that $\exists y_{1},y_{2}$, such that
$(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing\Longrightarrow y_{1,\mathsf{k}}\neq y_{2,\mathsf{k}}.$
Take $x$, such that $\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}$. This can be rewritten as $\{y_{1,\mathsf{k}},y_{2,\mathsf{k}}\}\subseteq\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}(x_{\mathsf{r}}-y_{\mathsf{r}})$, which means that $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is not single-valued. Contradiction is achieved.
4.4 More remarks and examples
Regarding $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$, be aware of the case where $\mathcal{A}$ is locally bounded everywhere [5, Definition 21.10], and thus, the input $x$ for $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ should be carefully selected to ensure $\mathcal{Q}x\in\mathsf{dom}\mathcal{A}^{-1}=\mathsf{ran}\mathcal{A}$. For example, consider the case of $\mathcal{A}=\partial f$, where $f=\|\cdot\|_{1}:\mathbb{R}^{n}\mapsto\mathbb{R}$. Then, $\mathsf{ran}\mathcal{A}=\mathsf{dom}\mathcal{A}^{-1}=[-1,1]^{n}$.
The following guiding example shows that the single-valuedness of $\mathcal{T}$ may depend on the choice of $\mathcal{Q}$. Given a function $f:\mathbb{R}^{2}\mapsto\mathbb{R}:(x_{1},x_{2})\mapsto|x_{1}|$, consider a minimization problem w.r.t. $x=(x_{1},x_{2})\in\mathbb{R}^{2}$:
$\min_{x\in\mathbb{R}^{2}}f(x)+\frac{1}{2}(x_{1}-b)^{2}\Longleftrightarrow 0\in\partial|\cdot|(x_{1}^{\star})+(x_{1}^{\star}-b),$ (17)
This is further written as the PPA form (2):
$\begin{bmatrix}0\\ 0\end{bmatrix}\in\begin{bmatrix}\partial|\cdot|&0\\ 0&0\end{bmatrix}\begin{bmatrix}x_{1}^{\star}\\ x_{2}^{\star}\end{bmatrix}+\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\begin{bmatrix}x_{1}^{\star}-b\\ x_{2}^{\star}-0\end{bmatrix}$
with the input variable $y=(b,0)$.
Obviously, the optimal solution set is $X^{\star}=\{x_{1}^{\star}\}\times\mathbb{R}$, where $x_{1}^{\star}$ is given by a well-known soft-thresholding function of $b$:
$x_{1}^{\star}=S(b):=\left\{\begin{array}[]{ll}b-1,&\textrm{if\ }b>1;\\ b+1,&\textrm{if\ }b<-1;\\ 0,&\textrm{if\ }|b|\leq 1.\end{array}\right.$
We also obtain
$\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(y_{\mathsf{r}}-x^{\star}_{\mathsf{r}})=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(y_{\mathsf{r}}-x^{\star}_{\mathsf{r}})=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\begin{bmatrix}b-S(b)\\ 0\end{bmatrix}=\{0\}\times\mathbb{R},$
which is multi-valued.
However, given the same $f$, (17) is changed as:
$\min_{x\in\mathbb{R}^{2}}f(x)+\frac{1}{2}(x_{2}-b)^{2}\Longleftrightarrow 0\in\partial\partial|\cdot|(x_{1}^{\star})+x_{2}^{\star}-b.$
This corresponds to
$\begin{bmatrix}0\\ 0\end{bmatrix}\in\begin{bmatrix}\partial|\cdot|&0\\ 0&0\end{bmatrix}\begin{bmatrix}x_{1}^{\star}\\ x_{2}^{\star}\end{bmatrix}+\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\begin{bmatrix}x_{1}^{\star}-0\\ x_{2}^{\star}-b\end{bmatrix}$
with input $y=(0,b)$. The unique solution is obviously $x^{\star}=(0,b)$. On the other hand, it is easy to check that $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(y_{\mathsf{r}}-x^{\star}_{\mathsf{r}})=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\begin{bmatrix}0\\ 0\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix}$. This shows that the single-valuedness of $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ complies with the uniqueness of solution, $\forall b\in\mathbb{R}$.
5 Convergence analysis
5.1 Convergence in the range space
All the results presented above are concerned with the single step of (5) only, and are not related to the existence of solution to (1). From now on, we focus on the iterative process (2) and aim at solving (1), and thus, begin to assume
Assumption 6
[Existence of solution to (1)] $\mathsf{zer}\mathcal{A}=\mathcal{A}^{-1}(0)\neq\varnothing$.
Remark 5.1.
Assumption 6 is quite a standard condition in literature, though it would be also interesting to investigate the behaviours of PPA (2) for the case of $\mathsf{zer}\mathcal{A}=\varnothing$, which is known as inconsistent feasibility setting in [3].
The following is an intermediate result in the the proof of [22, Theorem 3.4], which is crucial for our convergence analysis.
Lemma 5.2.
Under Assumptions 1, 2 and 6, given $\mathcal{T}$ as (5), then
$\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A}).$
Proof 5.3.
(1) $x\in\mathsf{zer}\mathcal{A}\Longrightarrow 0\in\mathcal{A}x\Longrightarrow\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})x\Longrightarrow x\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x\Longrightarrow\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x)=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x)\Longrightarrow\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x\in\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$. On the other hand, $x\in\mathsf{zer}\mathcal{A}\Longrightarrow\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$. Thus, we obtain $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})\subseteq\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$.
(2) Let $x\in\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})\Longrightarrow x=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})x\Longrightarrow\mathcal{Q}x=\mathcal{Q}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})x=\mathcal{Q}\mathcal{T}x$. On the other hand, $\mathcal{T}=(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\Longrightarrow 0\in\mathcal{A}(\mathcal{T}x)+\mathcal{Q}\mathcal{T}x-\mathcal{Q}x$, $\forall x\in\mathsf{dom}\mathcal{T}$. Combining both above, we then have $x\in\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})\Longrightarrow x\in\mathcal{A}(\mathcal{T}x)\Longrightarrow\mathcal{T}x\subseteq\mathsf{zer}\mathcal{A}\Longrightarrow(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x)\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$. Since $x=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x)$, we have $x\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$. Thus, $\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})\subseteq\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$.
Now, we discuss the convergence of (2). The convergence of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x^{k}\rightharpoonup\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x^{\star}$ has been shown in [22, Theorem 3.4] in a finite dimensional setting. Now, we extend this result to infinite dimensional case. For notational simplicity, we in the proof always denote $x_{\mathsf{r}}^{k}:=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x^{k}$, $\forall k\in\mathbb{N}$.
Theorem 5.4.
[Weak convergence in $\mathrm{ran}\mathcal{Q}$] Under Assumptions 1, 2, 3 and 6, let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by (2). Then $x_{\mathsf{r}}^{k}\rightharpoonup\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x^{\star}$ for some $x^{\star}\in\mathsf{zer}\mathcal{A}$, as $k\rightarrow\infty$.
Proof 5.5.
The proof is divided into 3 steps:
Step-1: For every $x_{\mathsf{r}}^{\star}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$, $\lim_{k\rightarrow\infty}$ $\|x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}}$ exists.
Step-2: $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ has at least one weak sequential cluster point lying in $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$;
Step-3: The weak sequential cluster point of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ is unique.
Step-1: First, we define $x_{\mathsf{r}}^{\star}\in\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$. We also have $x_{\mathsf{r}}^{\star}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$ by Lemma 5.2, which implies that $\exists x^{\star}\in\mathsf{zer}\mathcal{A}$, such that $x_{\mathsf{r}}^{\star}=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x^{\star}$. According to Fact 3, substituting $x=x^{k}$ and $y=x_{\mathsf{r}}^{\star}$ into (8), we have, by Fact 3, that
$\big{\|}x_{\mathsf{r}}^{k+1}-x_{\mathsf{r}}^{\star}\big{\|}_{\mathcal{Q}}^{2}\leq\big{\|}x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\big{\|}_{\mathcal{Q}}^{2}-\big{\|}x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{k+1}\|_{\mathcal{Q}}^{2}.$ (18)
(18) guarantees that $\{\|x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}}\}_{k\in\mathbb{N}}$ is non-increasing, and bounded from below (always being non-negative), and thus, convergent, i.e., $\lim_{k\rightarrow\infty}$ $\|x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}}$ exists.
Step-2: (18) implies the boundedness of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$. Indeed, from (18) and the closedness of $\mathsf{ran}\mathcal{Q}$, there exists $\alpha>0$, such that
$\alpha\big{\|}x_{\mathsf{r}}^{k}\big{\|}-\alpha\big{\|}x_{\mathsf{r}}^{\star}\big{\|}\leq\alpha\big{\|}x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\big{\|}\leq\big{\|}x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\big{\|}_{\mathcal{Q}}\leq\big{\|}x_{\mathsf{r}}^{0}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}},\quad\forall k\in\mathbb{N}.$
Thus, $\big{\|}x_{\mathsf{r}}^{k}\big{\|}\leq\big{\|}x_{\mathsf{r}}^{\star}\big{\|}+\frac{1}{\alpha}\big{\|}x_{\mathsf{r}}^{0}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}}:=\rho$, $\forall k\in\mathbb{N}$, and $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}\subseteq B(0,\rho)\cap\mathsf{ran}\mathcal{Q}$. $B(0,\rho)$ is weakly sequentially compact (by [5, Lemma 2.45]), and also is $B(0,\rho)\cap\mathsf{ran}\mathcal{Q}\subseteq B(0,\rho)$ (by [5, Lemma 1.34]). Then, $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ must possess at lease one weak sequential cluster point $v^{*}\in B(0,\rho)\cap\mathsf{ran}\mathcal{Q}$, i.e., there exists a subsequence $\{x_{\mathsf{r}}^{k_{i}}\}_{k\in\mathbb{N}}$ that weakly converges to $v^{*}$, as $k_{i}\rightarrow\infty$. We here need to show that $v^{*}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$, and more generally, every weak sequential cluster point of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ belongs to $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$. To this end, summing up (18) from $k=0$ to $K-1$, and taking $K\rightarrow\infty$, we have
$\sum_{k=0}^{\infty}\big{\|}x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{k+1}\big{\|}_{\mathcal{Q}}^{2}\leq\big{\|}x_{\mathsf{r}}^{0}-x_{\mathsf{r}}^{\star}\big{\|}_{\mathcal{Q}}^{2}<+\infty.$
This implies that $x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{k+1}=(\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}}^{k})\rightarrow 0$, and also $(\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}}^{k_{i}})\rightarrow 0$ as $k_{i}\rightarrow\infty$. Combining with $x_{\mathsf{r}}^{k_{i}}\rightharpoonup v^{*}$ and the $\mathcal{Q}$-demiclosedness of $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ (shown in Lemma 2.9), we obtain $(\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(v^{*})=0\Longrightarrow v^{*}\in\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$. Since $\{x_{\mathsf{r}}^{k_{i}}\}_{k_{i}\in\mathbb{N}}$ is an arbitrary weakly convergence subsequence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$, we conclude that every weak sequential cluster point of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$, and furthermore, in $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$ as well, again by Lemma 5.2.
Step-3: We need to show that $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ cannot have two distinct weak sequential cluster point in $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$. Indeed, let $v_{1}^{*},v_{2}^{*}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$ be two cluster points of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$. Since $\lim_{k\rightarrow\infty}\|x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}}$ exists as proved in Step-1, set $l_{1}=\lim_{k\rightarrow\infty}\|x_{\mathsf{r}}^{k}-v_{1}^{*}\|_{\mathcal{Q}}$, and $l_{2}=\lim_{k\rightarrow\infty}\|x_{\mathsf{r}}^{k}-v_{2}^{*}\|_{\mathcal{Q}}$. Take a subsequence $\{x_{\mathsf{r}}^{k_{i}}\}_{k_{i}\in\mathbb{N}}$ weakly converging to $v_{1}^{*}$, as $k_{i}\rightarrow\infty$. From the identity:
$\big{\|}x_{\mathsf{r}}^{k_{i}}-v_{1}^{*}\big{\|}_{\mathcal{Q}}^{2}-\big{\|}x_{\mathsf{r}}^{k_{i}}-v_{2}^{*}\big{\|}_{\mathcal{Q}}^{2}=\big{\|}v_{1}^{*}-v_{2}^{*}\big{\|}_{\mathcal{Q}}^{2}+2\big{\langle}v_{1}^{*}-v_{2}^{*}\big{|}v_{2}^{*}-x_{\mathsf{r}}^{k_{i}}\big{\rangle}_{\mathcal{Q}},$
taking $k_{i}\rightarrow\infty$ on both sides, the last term becomes:
$\lim_{k_{i}\rightarrow\infty}\big{\langle}v_{1}^{*}-v_{2}^{*}\big{|}v_{2}^{*}-x_{\mathsf{r}}^{k_{i}}\big{\rangle}_{\mathcal{Q}}=\lim_{k_{i}\rightarrow\infty}\big{\langle}v_{1}^{*}-v_{2}^{*}\big{|}v_{2}^{*}-x_{\mathsf{r}}^{k_{i}}\big{\rangle}_{\mathcal{Q}}=-\|v_{1}^{*}-v_{2}^{*}\|_{\mathcal{Q}}^{2},$
then, we deduce that $l_{1}^{2}-l_{2}^{2}=-\big{\|}v_{1}^{*}-v_{2}^{*}\big{\|}_{\mathcal{Q}}^{2}$. Similarly, take a subsequence $\{x^{k_{j}}\}_{k_{j}\in\mathbb{N}}$ weakly converging to $v_{2}^{*}$, as $k_{j}\rightarrow\infty$, which yields that $l_{1}^{2}-l_{2}^{2}=\big{\|}v_{1}^{*}-v_{2}^{*}\big{\|}_{\mathcal{Q}}^{2}$. Consequently, $\big{\|}v_{1}^{*}-v_{2}^{*}\big{\|}_{\mathcal{Q}}=0\Longrightarrow v_{1}^{*}=v_{2}^{*}$, since $v_{1}^{*},v_{2}^{*}\in\mathsf{ran}\mathcal{Q}$. This shows the uniqueness of the weak sequential cluster point.
Finally, combining the above 3 step and [5, Lemma 2.46], we conclude the weak convergence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$, and denote the weak limit as $x_{\mathsf{r}}^{\star}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$.
5.2 Discussions and related works
There are lots of technical details in the proof of Theorem 5.4 that need to be clarified, discussed and further compared to other related works.
First, Lemma 5.2 makes the connection between the degenerate PPA (2) and the original inclusion problem (1). One can always show the weak convergence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ in $\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$ without Lemma 5.2. This is Lemma 5.2 that tells us what the degenerate PPA (2) solves for is actually a point of $\mathsf{zer}\mathcal{A}$ projected onto the range space of degenerate metric.
Note that Assumption 5 is not needed in Theorem 5.4, $\mathcal{T}$ is allowed to be multi-valued in the proof, since we are only concerned with $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$, which is already well-defined by Assumptions 1 and 2. This is fundamentally different from [22, Theorem 3.4], where $\mathcal{T}$ is assumed to be well-defined. In addition, $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is also not necessarily maximally monotone here.
[22, Theorem 3.4] discussed the convergence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ in the finite-dimensional setting, where the proof heavily relied on the assumption of continuity of $\mathcal{T}$. We here argue that the continuity of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ would suffice for [22, Theorem 3.4], which is obviously true in finite-dimensional case by Fact 3. However, $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}$ is not weakly continuous generally in infinite-dimensional case. Instead, we exploited the weak-to-strong continuity of $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ in Lemma 2.9.
[…]
Our proof uses the $\mathcal{Q}$-demiclosedness of $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}$. Another alternative way is to use the maximal monotonicity of $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ (i.e., Assumption 3) to finish the proof as well.
Proof 5.6.
Applying both sides of Lemma 2.14-(4) to $x_{\mathsf{r}}^{k}$, we have:
$\sqrt{\mathcal{Q}}x_{\mathsf{r}}^{k}-\sqrt{\mathcal{Q}}x_{\mathsf{r}}^{k+1}=\Big{(}\mathcal{I}+\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}^{-1}\sqrt{\mathcal{Q}}x_{\mathsf{r}}^{k},$ (19)
where the single-valuedness of the right-hand side has been shown in the proof of Lemma 2.14. Let $q_{\mathsf{r}}^{k}:=\sqrt{\mathcal{Q}}x_{\mathsf{r}}^{k}$, $\mathcal{B}:=\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}\big{|}_{\mathsf{ran}\mathcal{Q}}$, then, (19) is rewritten as (by maximal monotonicity of $\mathcal{B}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ and Moreau’s identity):
$q_{\mathsf{r}}^{k+1}=q_{\mathsf{r}}^{k}-(\mathcal{I}+\mathcal{B})^{-1}q_{\mathsf{r}}^{k}=\big{(}\mathcal{I}+\mathcal{B}^{-1}\big{)}^{-1}q_{\mathsf{r}}^{k}\Longrightarrow\mathcal{B}^{-1}q_{\mathsf{r}}^{k+1}\owns q_{\mathsf{r}}^{k}-q_{\mathsf{r}}^{k+1}.$
It is known from the proof of Theorem 5.4 that $q_{\mathsf{r}}^{k}-q_{\mathsf{r}}^{k+1}\rightarrow 0$ and $q_{\mathsf{r}}^{k}\rightharpoonup q^{\star}$, as $k\rightarrow\infty$. Since $\mathcal{B}^{-1}$ is also maximally monotone (by [5, Proposition 20.22]), then $\mathsf{gra}\mathcal{B}^{-1}$ is sequentially closed in $(\mathsf{ran}\mathcal{Q})^{\text{weak}}\times(\mathsf{ran}\mathcal{Q})^{\text{strong}}$, i.e., $(\mathsf{ran}\mathcal{Q},\|\cdot\|_{\mathcal{Q}})\times(\mathsf{ran}\mathcal{Q},\langle\cdot|\cdot\rangle_{\mathcal{Q}})$, by [5, Proposition 20.38]. Thus, we obtain
$0\in\mathcal{B}^{-1}q^{\star}\Longrightarrow q^{\star}\in\mathcal{B}(0)=\sqrt{\mathcal{Q}}\mathcal{A}^{-1}(0)\Longrightarrow x_{\mathsf{r}}^{\star}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}(0)=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A}).$
Finally, we stress that $\{x^{k}\}_{k\in\mathbb{N}}$ may be unbounded due to the uncontrolled component of $x_{\mathsf{k}}^{k}:=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}x^{k}$. Thus, $\{x^{k}\}_{k\in\mathbb{N}}$ may not even possess any weak sequential cluster point. The only exception is the non-degenerate case, where the weak convergence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ is obviously equivalent to that of $\{x^{k}\}_{k\in\mathbb{N}}$. Assumption 3 in this case boils down to the usual maximal monotonicity of $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$.
5.3 Convergence in the whole space
To proceed further, if $\mathcal{T}$ is single-valued (i.e., Assumption 5), it is then natural to ask when the iterations (2) converge in the whole space. To obtain the convergence of $\{x^{k}\}_{k\in\mathbb{N}}$, the boundedness should be a basic requirement. To this end, [11, Theorem 2.9] assumed the Lipschitz continuity of $(\mathcal{A}+\mathcal{Q})^{-1}$. Here, we propose
Assumption 7
For any two inputs $x_{i,\mathsf{r}}\in\mathsf{ran}\mathcal{Q}$ for $i=1$ and 2, and the corresponding $y_{i,\mathsf{r}}=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{i,\mathsf{r}})$,
(i) the solution to the constrained inclusion of $\mathcal{A}(y_{i,\mathsf{k}}+y_{i,\mathsf{r}})\owns\mathcal{Q}(x_{i,\mathsf{r}}-y_{i,\mathsf{r}})$, s.t. $\mathcal{Q}y_{i,\mathsf{k}}=0$, is unique, and satisfies $\|y_{1,\mathsf{k}}-y_{2,\mathsf{k}}\|\leq\xi\|x_{1,\mathsf{r}}-x_{2,\mathsf{r}}\|$ for some constant $\xi$;
(ii) the solution to the inclusion $\mathcal{A}(\mathcal{E}^{\top}b_{i}+y_{i,\mathsf{r}})\owns\mathcal{Q}(x_{i,\mathsf{r}}-y_{i,\mathsf{r}})$ is unique, and satisfies $\|b_{1}-b_{2}\|\leq\xi\|x_{1,\mathsf{r}}-x_{2,\mathsf{r}}\|$ for some constant $\xi$, where $\mathcal{E}$ is defined in Theorem 4.4-(iii),
and show that both Assumption 7-(i) and (ii) are equivalent to the Lipschitz continuity of $(\mathcal{A}+\mathcal{Q})^{-1}$, presented below.
Proposition 5.7.
Under Assumptions 1, 2, 3 and 5, then the following 3 conditions are equivalent:
(i) Assumption 7-(i);
(ii) Assumption 7-(ii);
(iii) Lipschitz continuity of $(\mathcal{A}+\mathcal{Q})^{-1}$ [11, Theorem 2.9].
Before proving this, we present several simple facts.
Fact 7.
Under Assumptions 1 and 2, given $\mathcal{T}$ in (2), then $\forall x\in\mathsf{ran}\mathcal{Q}$, $\exists\lambda_{\inf},\lambda_{\sup}>0$, such that
(i) $\lambda_{\inf}\|x\|^{2}\leq\|x\|_{\mathcal{Q}}^{2}\leq\lambda_{\sup}\|x\|^{2}$;
(ii) $\lambda_{\inf}\|x\|\leq\|\mathcal{Q}x\|\leq\lambda_{\sup}\|x\|$;
(iii) $\|\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{T}x_{1}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{T}x_{2}\|\leq\sqrt{\frac{\lambda_{\sup}}{\lambda_{\inf}}}\|x_{1,\mathsf{r}}-x_{2,\mathsf{r}}\|$, $\forall x_{1},x_{2}\in\mathsf{dom}\mathcal{T}$.
Proof 5.8.
(i)-(ii): Due to the boundedness and closedness of $\mathcal{Q}$ as Assumption 1.
(iii) Let $y_{i}\in\mathcal{T}x_{i}$ for $i=1$ and 2. Since $y_{i,\mathsf{r}}:=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}y_{i}$ for $i=1$ and 2 is single-valued by Lemma 2.2, we then have:
[…]
Then, we prove Proposition 5.7.
Proof 5.9.
(i)$\Longleftrightarrow$(ii): $\|y_{1,\mathsf{k}}-y_{2,\mathsf{k}}\|^{2}=\|\mathcal{E}^{\top}b_{1}-\mathcal{E}^{\top}b_{2}\|^{2}=\|b_{1}-b_{2}\|^{2}$, since $\mathcal{E}\mathcal{E}^{\top}=I$.
(i)$\Longrightarrow$(iii):
$\|(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x_{1}-(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x_{2}\|$
$=$ $\|y_{1}-y_{2}\|=\|y_{1,\mathsf{r}}+y_{1,\mathsf{k}}-y_{2,\mathsf{r}}-y_{2,\mathsf{k}}\|$
$\leq$ $\|y_{1,\mathsf{r}}-y_{2,\mathsf{r}}\|+\|y_{1,\mathsf{k}}-y_{2,\mathsf{k}}\|$
[…]
(iii)$\Longrightarrow$(i):
$\|y_{1,\mathsf{k}}-y_{2,\mathsf{k}}\|$$\leq$ $\|y_{1}-y_{2}\|+\|y_{1,\mathsf{r}}-y_{2,\mathsf{r}}\|$
[…]
Remark 5.10.
All the conditions in Proposition 5.7 are equivalent to $\|\mathcal{T}x_{1}-\mathcal{T}x_{2}\|\leq\xi\|x_{1}-x_{2}\|_{\mathcal{Q}}$, which can be found in [34, Theorem 3.3] and [12, Lemma 3.2].
Then, the convergence in the whole space immediately follows.
Theorem 5.11.
Let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by (2), which satisfies Assumptions 1, 2, 3, 5, 6 and any condition of Proposition 5.7. Furthermore, if any of the following conditions hold:
(i) every weak sequential cluster point of $\{x^{k}\}_{k\in\mathbb{N}}$ lies in $\mathsf{Fix}\mathcal{T}$;
(ii) Assumption 4;
(iii) $\mathcal{I}-\mathcal{T}$ is demiclosed,
then $\{x^{k}\}_{k\in\mathbb{N}}$ weakly converges to some $x^{\star}\in\mathsf{zer}\mathcal{A}$, as $k\rightarrow\infty$.
Proof 5.12.
Proposition 5.7 guarantees the boundedness of $\{x^{k}\}_{k\in\mathbb{N}}$. Then there exist two subsequences $\{x^{k_{i}}\}_{k\in\mathbb{N}}$ and $\{x^{k_{j}}\}_{k\in\mathbb{N}}$, such that $x^{k_{i}}\rightharpoonup v_{1}^{*}$ and $x^{k_{j}}\rightharpoonup v_{2}^{*}$. By Step-3 of the proof of Theorem 5.4, we have $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}v_{1}^{*}=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}v_{2}^{*}$ and further $\mathcal{T}v_{1}^{*}=\mathcal{T}v_{2}^{*}$, i.e., all the weak sequential cluster points of $\{x^{k}\}_{k\in\mathbb{N}}$ share the same projection onto $\mathsf{ran}\mathcal{Q}$.
Condition (i): it implies that $v_{1}^{*}=\mathcal{T}v_{1}^{*}=\mathcal{T}v_{2}^{*}=v_{2}^{*}$. This shows that the weak sequential cluster point of $\{x^{k}\}_{k\in\mathbb{N}}$ is unique, and lies in $\mathsf{Fix}\mathcal{T}=\mathsf{zer}\mathcal{A}$. The conclusion is obtained by [5, Lemma 2.46].
Condition (ii): Assumption 4-(i) implies that $\mathsf{gra}\mathcal{A}$ is closed in $\mathcal{H}^{\text{weak}}\times\mathcal{H}^{\text{strong}}$. Take a subsequence $\{x^{k_{i}}\}_{k\in\mathbb{N}}$, such that $x^{k_{i}}\rightharpoonup v^{*}$. Then $\mathcal{T}x^{k_{i}}\rightharpoonup v^{*}$, since $x^{k_{i}}-\mathcal{T}x^{k_{i}}\rightarrow 0$. Considering $\mathcal{A}\mathcal{T}x^{k_{i}}\owns\mathcal{Q}(x^{k_{i}}-\mathcal{T}x^{k_{i}})$, combining with $\mathcal{T}x^{k_{i}}\rightharpoonup v^{*}$ and $x^{k_{i}}-\mathcal{T}x^{k_{i}}\rightarrow 0$, it then follows from the closedness of $\mathsf{gra}\mathcal{A}$ that $v^{\star}\in\mathsf{zer}\mathcal{A}=\mathsf{Fix}\mathcal{T}$, which leads to Condition (i).
Condition (iii): Considering the subsequence $\{x^{k_{i}}\}_{k\in\mathbb{N}}$ in (ii), the demiclosedness of $\mathcal{I}-\mathcal{T}$ yields that $(\mathcal{I}-\mathcal{T})v^{*}=0$, i.e., $v^{*}\in\mathsf{Fix}\mathcal{T}$, which also leads to Condition (i).
Conditions (i) and (ii) correspond to [11, Theorem 2.9, Corollary 2.10], respectively. Condition (iii), as mentioned in [11, Remark 2.11], is hard to control. The key problem here is that $\mathcal{T}$ is not necessarily nonexpansive, and thus the general demiclosedness of $\mathcal{I}-\mathcal{T}$ cannot be obtained by Browder’s principle [5, Theorem 4.27]. However, it can be tackled case-by-case, by exploiting particular structures of $\mathcal{A}$ and $\mathcal{Q}$, and additional assumptions. For instance, [12, Lemma 3.4], for a specific class of preconditioned ADMM, proposed several additional assumptions (e.g., compactness, weak closedness) to make $\mathcal{I}-\mathcal{T}$ demiclosed. [17, Theorem 3.3] seems the first attempt to establish the whole convergence under degenerate metric in a finite-dimensional setting, based on the continuity of proximity operator. Similar to [22, Theorem 3.4], this technique is not applicable for infinite-dimensional setting, since $\mathcal{T}$ is generally not weakly continuous.
6 Applications to operator splitting algorithms
6.1 Douglas–Rachford splitting (DRS)
The scheme of DRS has been given as (3), with the degenerate PPA interpretation of (4). The well-definedness of the DRS is obvious from (3). It can also be verified from (4) only via Assumption 4. First, it is clear that $\mathcal{A}:\mathcal{U}^{3}\mapsto 2^{\mathcal{U}}\times 2^{\mathcal{U}}\times\mathcal{U}$ is maximally monotone. Second, observe that $\mathsf{ran}\mathcal{A}=\mathcal{U}^{3}=\mathcal{H}$ and $0\in\mathsf{ran}\mathcal{Q}$, and thus, $0\in\mathsf{sri}(\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A})$, which fulfills Assumption 4-(ii). Then it follows that the scheme (3) admits any $(u,w,z)\in\mathcal{U}^{3}$ in any iterative step $k$.
A standard result in the seminal work [23] claims $z^{k}\rightharpoonup z^{\star}$ as $k\rightarrow\infty$, if $\tau>0$. This can also easily be obtained by our Theorem 5.4. The weak convergence of the solution trajectory $\{u^{k}\}_{k\in\mathbb{N}}$ had been a longstanding problem until the advent of [32]. To obtain this result, [32] used some sophisticated tools, e.g., quasi-Fejér convergence, strong/weak topologies. In [35, Theorem 4.2], the author used many variants of PPA forms of DRS (especially the complicated prediction+correction framework initialized by [21, 20, 19]). Now, we are going to show that the convergence of $\{u^{k}\}_{k\in\mathbb{N}}$ and $\{w^{k}\}_{k\in\mathbb{N}}$ can be directly obtained from (4) alone and via our Theorem 5.11. This is an improvement, compared to [35, Remark 3.1-(iii)], which once claimed that ‘it remains unclear from (4) that whether $\{u^{k}\}_{k\in\mathbb{N}}$ and $\{w^{k}\}_{k\in\mathbb{N}}$ are convergent’. This shows the great potential of using the PPA interpretation.
Proposition 6.1.
Let $\{u^{k},w^{k},z^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by (3). If $\tau>0$, then $(u^{k},w^{k},z^{k})\rightharpoonup(u^{\star},w^{\star},z^{\star})$ as $k\rightarrow\infty$, where $u^{\star}$ is an optimal solution to $\min_{u}f(u)+g(u)$.
Proof 6.2.
Since it is difficult to compute $\mathcal{A}^{-1}$, let us consider the inclusion of $\mathcal{A}(\mathcal{E}^{\top}b+y_{\mathsf{r}})\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$, and check if it has a unique solution of $b$ and if $b$ satisfies Assumption 7.
In the setting of (4), $\mathcal{E}=\begin{bmatrix}I&0&0\\ 0&I&0\end{bmatrix}$. Let $b=\begin{bmatrix}b_{1}\\ b_{2}\end{bmatrix}$, $x_{\mathsf{r}}=(0,0,a)$, and the corresponding $y_{\mathsf{r}}=(0,0,c)$ [^6] There is no need to find the specific expression of $c$ in terms of $a$. Of great importance here is the fact that $c$ is a nonexpansive mapping of $a$ by Fact 3., then the concrete form of $\mathcal{A}(\mathcal{E}^{\top}b+y_{\mathsf{r}})\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$ becomes:
$\begin{bmatrix}\tau\partial g&I&-I\\ -I&\tau\partial f&I\\ I&-I&0\end{bmatrix}\begin{bmatrix}b_{1}\\ b_{2}\\ c\end{bmatrix}\owns\begin{bmatrix}0\\ 0\\ a-c\end{bmatrix}.$
Putting $b_{2}=b_{1}-a+c$ (indicated by the last line) into the first line, we obtain $\tau\partial g(b_{1})+b_{1}-a+c-c=\tau\partial g(b_{1})+b_{1}-a\owns 0$, leading to $b_{1}=\operatorname{prox}_{\tau g}(a)$, and thus, $b_{2}=\operatorname{prox}_{\tau g}(a)-a+c=c-\operatorname{prox}_{g^{*}/\tau}(a)$. $b$ is obviously single-valued, which meets Assumption 5. Then, from the well-known nonexpansiveness of proximity operator (see [16, Lemma 2.4] for example), we have
$\|b-b^{\prime}\|$$\leq$ $\|b_{1}-b^{\prime}_{1}\|+\|b_{2}-b^{\prime}_{2}\|$
$=$ $\big{\|}\operatorname{prox}_{\tau g}(a)-\operatorname{prox}_{\tau g}(a^{\prime})\big{\|}+\big{\|}c-\operatorname{prox}_{g^{*}/\tau}(a)-c^{\prime}+\operatorname{prox}_{g^{*}/\tau}(a^{\prime})\big{\|}$
$\leq$ $\big{\|}a-a^{\prime}\big{\|}+\big{\|}a-a^{\prime}\big{\|}+\big{\|}c-c^{\prime}\big{\|}$
$\leq$ $3\big{\|}a-a^{\prime}\big{\|},$
where the last inequality is due to the $\|c-c^{\prime}\|\leq\|a-a^{\prime}\|$ by Fact 3. Finally, the weak convergence of $\{u^{k},w^{k},z^{k}\}_{k\in\mathbb{N}}$ follows by Theorem 5.11.
It is worth noting that the weak convergence of $\{u^{k}\}_{k\in\mathbb{N}}$ cannot be directly obtained from $u^{k}=\operatorname{prox}_{\tau g}(z^{k})$ and the weak convergence of $\{z^{k}\}_{k\in\mathbb{N}}$, since the proximity operator is not weak-to-weak continuous. The Lipschitz continuity of $u^{k}$ (w.r.t. and bounded by $z^{k}$) has to be incorporated into the degenerate PPA framework. In addition, similar to [32], Proposition 6.1 also removes the assumption of maximal monotonicity of $\partial f+\partial g$, but without using the instrumental result of [4, Corollary 3] as [32] did.
6.2 Alternating direction method of multipliers (ADMM)
Considering an equality constrained problem:
$\min_{s,t}f(s)+g(t),\quad\text{s.t.\ }As+Bt=0,$
the ADMM is given as [35, Eq.(22)]:
$\left\lfloor\begin{array}[]{lll}s^{k+1}&\in&\operatorname{Arg}\min_{s}f(s)+\frac{\tau}{2}\big{\|}As+Bt^{k}-\frac{1}{\tau}u^{k}\big{\|}^{2},\\ t^{k+1}&\in&\operatorname{Arg}\min_{t}g(t)+\frac{\tau}{2}\big{\|}As^{k+1}+Bt-\frac{1}{\tau}u^{k}\big{\|}^{2},\\ u^{k+1}&=&u^{k}-\tau(As^{k+1}+Bt^{k+1}).\\ \end{array}\right.$ (20)
Due to the equivalence between ADMM and DRS, we now transform the ADMM into a DRS form. Let $z^{k}=u^{k}+\tau B^{\top}t^{k}$, then, (20) becomes:
$\left\lfloor\begin{array}[]{lll}t^{k}&\in&\operatorname{Arg}\min_{t}g(t)+\frac{\tau}{2}\big{\|}Bt-\frac{1}{\tau}z^{k}\big{\|}^{2},\\ s^{k+1}&\in&\operatorname{Arg}\min_{s}f(s)+\frac{\tau}{2}\big{\|}As+2Bt^{k}-\frac{1}{\tau}z^{k}\big{\|}^{2},\\ z^{k+1}&=&z^{k}-\tau(As^{k+1}+Bt^{k}).\\ \end{array}\right.$ (21)
Similar to [35, Remark 3.1-(ii)], (21) can be rewritten in terms of infinal postcomposition [5, Definition 12.34]:
$\left\lfloor\begin{array}[]{lll}q^{k}&=&\arg\min_{q}G(q)+\frac{\tau}{2}\big{\|}q-\frac{1}{\tau}z^{k}\big{\|}^{2},\\ r^{k+1}&=&\arg\min_{r}F(r)+\frac{\tau}{2}\big{\|}r+2q^{k}-\frac{1}{\tau}z^{k}\big{\|}^{2},\\ z^{k+1}&=&z^{k}-\tau(r^{k+1}+q^{k}),\\ \end{array}\right.$ (22)
where $F=A\triangleright f$ and $G=B\triangleright g$ are infimal postcompositions of $f$ and $g$ by $A$ and $B$, respectively. Then, it is easy to see that (22) exactly fits into the form of DRS (3). Finally, by Proposition 6.1, it is easy to show the weak convergence of $\big{\{}q^{k},r^{k},z^{k}\big{\}}_{n\in\mathbb{N}}$.
The limiting case of the well-known Chambolle–Pock (CP) algorithm [15, Algorithm 1] can also be analyzed in a similar manner, based an augmentation technique [26], which equivalently expresses the CP algorithm as the DRS form (also see [35, Sect. 5.3] for details).
6.3 Augmented Lagrangian method (ALM)
Let us now consider a constrained optimization problem:
$\min_{u\in\mathcal{U}}f(u),\qquad\text{s.t.\ }Bu=b,$
where $B:\mathcal{U}\mapsto\mathcal{Y}$. The ALM reads as (see [33, Eq.(7.2)] for example):
$\left\lfloor\begin{array}[]{lll}u^{k+1}&\in&\operatorname{Arg}\min_{u}f(u)+\frac{\tau}{2}\big{\|}Bu-b-\frac{1}{\tau}p^{k}\big{\|}^{2},\\ p^{k+1}&=&p^{k}-\tau(Bu^{k+1}-b).\\ \end{array}\right.$ (23)
Using the notion of infimal postcomposition mentioned above, we obtain the equivalent form to (23):
$\left\lfloor\begin{array}[]{lll}q^{k+1}&=&\arg\min_{q}F(q)+\frac{\tau}{2}\big{\|}q-b-\frac{1}{\tau}p^{k}\big{\|}^{2},\\ p^{k+1}&=&p^{k}-\tau(q^{k+1}-b),\\ \end{array}\right.$ (24)
where $F^{*}=B\triangleright f$. The corresponding PPA interpretation is
$\begin{bmatrix}0\\ 0\end{bmatrix}\in\underbrace{\begin{bmatrix}\partial F&-I\\ I&\partial l\end{bmatrix}}_{\mathcal{A}}\begin{bmatrix}q^{k+1}\\ p^{k+1}\end{bmatrix}+\underbrace{\begin{bmatrix}0&0\\ 0&\frac{1}{\tau}I\end{bmatrix}}_{\mathcal{Q}}\begin{bmatrix}q^{k+1}-q^{k}\\ p^{k+1}-p^{k}\end{bmatrix},$ (25)
where the linear function $l$ is defined as $l=-\langle\cdot|b\rangle$.
Theorem 6.3.
Let $\{q^{k},p^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by (24). If $\tau>0$, then $(q^{k},p^{k})\rightharpoonup(b,p^{\star})$ as $k\rightarrow\infty$.
Proof 6.4.
First, observing (25) obtains that $\mathcal{A}$ is maximally monotone, $\mathsf{ran}\mathcal{A}=\mathcal{Y}^{2}=\mathcal{H}$, and thus, $0\in\mathsf{sri}(\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A})$, which fulfills Assumption 4. Then it follows that the scheme (24) admits any input $x=(q,p)\in\mathcal{Y}^{2}$.
Then, we check Assumptions 5 and 7. Here, $\mathcal{E}=\begin{bmatrix}I&0\end{bmatrix}$. Given an input $x_{\mathsf{r}}=(0,a)$ and the corresponding output $y_{\mathsf{r}}=(0,c)$, the inclusion of $\mathcal{A}(\mathcal{E}^{\top}v+y_{\mathsf{r}})\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$ becomes:
$\begin{bmatrix}\partial F&-I\\ I&\partial l\end{bmatrix}\begin{bmatrix}v\\ c\end{bmatrix}\owns\begin{bmatrix}0\\ \frac{1}{\tau}(a-c)\end{bmatrix}.$
The second line yields that $v=\frac{1}{\tau}(a-c)+b$, which is obviously single-valued (satisfying Assumption 5). And furthermore, $\|v-v^{\prime}\|=\frac{1}{\tau}\|(a-a^{\prime})-(c-c^{\prime})\|\leq\frac{2}{\tau}\|a-a^{\prime}\|$, i.e., $\frac{2}{\tau}$-Lipschitz continuous. Finally, the weak convergence of $\{q^{k},p^{k}\}_{k\in\mathbb{N}}$ follows by Theorem 5.11.
The basic Bregman iteration [27] can be analyzed in a similar manner, due to its essential equivalence to the ALM.
7 Concluding remarks
This work provided a fundamental understanding of degenerate metric resolvent and its associated convergence issue. Our presented results can be further exemplified with more applications:
  • Analyze the existence and uniqueness of solution to the regularized least-squares problem: $\min_{x\in\mathbb{R}^{n}}f(x)+\frac{1}{2}\|Ax-b\|^{2}$, particularly including (generalized) lasso (corresponding to $f=\|D\cdot\|_{1}$) and give a geometric view for some special classes of the function $f$.
  • Discuss the convergence property of $\{t^{k},s^{k}\}_{k\in\mathbb{N}}$ in the non-proximized ADMM (21), This is closely related to the interatctions between $A$, $B$ and $f$, $g$. Take a similar treatment to $\{u^{k}\}_{k\in\mathbb{N}}$ in (23).
Some theoretical aspects are also of great interest, for instance,
  • Explore the properties of $\mathcal{T}$, if $\mathcal{Q}$ fails to have closed range, as indicated in [7].
  • Study the behaviours of (2) if the inclusion problem (1) is inconsistent, as proposed in [3].
  • Find a simple interpretation of Assumption 4-(ii) could also be interesting for practitioners.
8 Acknowledgements
We are gratefully indebted to Mr. Brecht Evens (KU Leuven, Belgium) for bringing [22, Theorem 3.4] to our attention and inspiring comments on our manuscript. Our thanks also go to Mr. Shuchang Zhang (NUDT, China) for detailed discussions on [35, Theorem 2.1].