### v3

# On degenerate metric resolvent: well-definedness and new convergence analysis

Abstract
We study the basic properties of resolvent equipped with degenerate metric based on restricted maximal monotonicity, and further propose several conditions for the well-definedness and weak convergence of the fixed point iterations. The results help to understand the behaviours of many operator splitting algorithms in the kernel space of degenerate metric.
keywords:
Resolvent, degenerate metric, restricted maximal monotonicity, proximal point algorithm (PPA), operator splitting algorithms.
{AMS}
47H05, 49M29, 49M27, 90C25
1 Introduction
1.1 Motivation
In nonlinear analysis and optimization, a generic and central assignment is to obtain numerical solutions to the inclusion problem
$\textrm{find }x\in\mathcal{H},\quad\textrm{s.t. }0\in\mathcal{A}x$ (1)
where $\mathcal{H}$ is a real Hilbert space and $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is a set-valued operator. Built on the classical proximal point algorithm (PPA) in 1970s [24, 27], many splitting methods for solving (1) have been developed in the literature during the past five decades. Very recently, a degenerate preconditioned version of PPA was studied in [10], which provided a powerful framework for analyzing various existing splitting methods and also for devising new and flexible schemes. The degenerate preconditioned PPA can be described by the following fixed-point iterate
$x^{k+1}=\mathcal{T}x^{k},\quad\textrm{where\ }\mathcal{T}:=(\mathcal{Q}+\mathcal{A})^{-1}\mathcal{Q},$ (2)
where $\mathcal{Q}:\mathcal{H}\mapsto\mathcal{H}$ is a linear, bounded, self-ajoint and positive semi-definite operator. To ensure the iterate to be well-defined, the operate $\mathcal{T}$ has to be single-valued and of full domain. However, this key condition was only stated as a basic assumption in [10, Definition 2.1] and hence motivates us to find sufficient guarantees for the well-definedness of $\mathcal{T}$.
Among many operator splitting algorithms prevails the degeneracy phenomenon, which makes the behaviours of (2) more subtle. For instance, the celebrated Douglas–Rachford splitting (DRS), for solving $\min_{u\in\mathcal{U}}(u)+g(u)$, reads as
$\left\lfloor\begin{array}[]{lll}u^{k+1}&=&\arg\min_{u}g(u)+\frac{1}{2\tau}\big{\|}u-z^{k}\big{\|}^{2},\\ w^{k+1}&=&\arg\min_{w}f(w)+\frac{1}{2\tau}\big{\|}w-(2u^{k+1}-z^{k})\big{\|}^{2},\\ z^{k+1}&=&z^{k}+(w^{k+1}-u^{k+1}),\end{array}\right.$ (3)
which exactly fits the form of (2) according to
$\mathcal{H}=\mathcal{U}^{3},\ x^{k}=\begin{bmatrix}u^{k}\\ w^{k}\\ z^{k}\end{bmatrix},\ \mathcal{A}=\begin{bmatrix}\tau\partial g&I&-I\\ -I&\tau\partial f&I\\ I&-I&0\end{bmatrix},\ \mathcal{Q}=\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&0&I\end{bmatrix}.$ (4)
Here, we adopt a different (but essentially equivalent) scheme (3) from [10, Eq.(1.6)] for two reasons: (i) (3) explicitly preserves both solution trajectory $\{u^{k}\}_{k\in\mathbb{N}}$ and shadow sequence $\{z^{k}\}_{k\in\mathbb{N}}$; (ii) the degenerate metric $\mathcal{Q}$ in (4) is easier to analyze than that in [10, Eq.(1.5)]. Then, one may ask: Can we conclude the well-definedness of the iterations and the convergence of $\{u^{k}\}_{k\in\mathbb{N}}$ from (4) only (without ad hoc analysis as [30] did)? More broadly, this further motivates us to study a general setting of (2) and find the sufficient conditions on $\mathcal{A}$ and $\mathcal{Q}$, to guarantee the well-definedness of $\mathcal{T}$ and the convergence of $\{x^{k}\}_{k\in\mathbb{N}}$.
1.2 Related works
When $\mathcal{Q}$ is replaced by a general non-linear operator, the operator $\mathcal{T}$ in (2) was coined the term warped resolvent in [13, Definition 1.1] and was ensured to be well-defined by requiring that $\mathsf{ran}\mathcal{Q}\subseteq\mathsf{ran}(\mathcal{A}+\mathcal{Q})$ and disjoint injectivity of $\mathcal{A}+\mathcal{Q}$ (see [12, Proposition 2.1] for instance). Besides, a group of sufficient conditions for well-definedness of $\mathcal{T}$ are listed in [13, Proposition 3.9] such as uniform/strict/strong monotonicity. However, these conditions are much strong and hard to verify. Since in this study we only focus on the operator $\mathcal{T}$ with $\mathcal{Q}$ being linear, we are able to find more mild conditions on $\mathcal{A}$ and $\mathcal{Q}$ so that the operator $\mathcal{T}$ to be well-defined.
The convergence issue of degenerate version of (2) has been first tackled in [16, 22] for specific algorithms (corresponding to particular $\mathcal{A}$ and $\mathcal{Q}$) in finite dimensional setting. However, the continuity of the proximity operator—as a main assumption proposed therein—is generally not satisfied in weak topology of infinite dimensional space. For infinite dimensional case, [11] showed the weak convergence for a class of preconditioned ADMM (associated with degenerate preconditioner) based on demiclosedness principle (under additional conditions listed in [11, Lemma 3.4]), [10] further performed a systematic study for general $\mathcal{A}$ and $\mathcal{Q}$. We in this paper for the more general scheme (2) provide more new convergence results with weakened conditions than existing works.
1.3 Contributions
This work aims to provide a fundamental analysis and understanding of the degenerate metric resolvent. The paper is organized as follows. Sect. 2 presents the basic properties of $\mathcal{T}$ under the restricted monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$, which are used as preliminary results in Sect. 35 for analyzing full domain, single-valuedness and convergence, respectively. Finally, Sect. 6 exemplifies the results with several splitting algorithms. The contributions are in order.
(1) Full domain: We in Theorem 3.1 give a sufficient and necessary condition for the full domain of $\mathcal{T}$ and further in Corollary 3.7 provide a sufficient condition that is easy to verify in practice. These results are obtained by generalized Moreau’s decomposition identity (cf. Lemma A.1) and restricted Minty’s theorem (cf. Theorem 3.4), based on a notion of restricted maximal monotonicity.
(2) Single-valuedness: Sect. 4 presents sufficient (and necessary) conditions for the single-valuedness of $\mathcal{T}$. Our results weaken the conditions of existing works, from injectivity to restricted injectivity.
(3) Convergence: Theorem 5.3 shows the weak convergence in the range space of $\mathcal{Q}$ under restricted maximal monotonicity only. This result extends [22, Theorem 3.4] to infinite dimensional setting with milder condition than [22]. Notably, the operator $\mathcal{A}$ is not assumed to be maximal monotone for this result. We further show in Theorem 5.6 the whole convergence of $\{x^{k}\}_{k\in\mathbb{N}}$ under additional conditions. The comparisons with existing works are also discussed in details.
1.4 Notations, assumptions and preliminaries
[…]
Throughout the paper, we make the following assumption on $\mathcal{Q}$.
Assumption 1
(cf. [33, Assumption 2.1])
(i) [Basic] $\mathcal{Q}:\mathcal{H}\mapsto\mathcal{H}$ is linear, bounded, self-adjoint and positive semi-definite, and $\mathsf{dom}\mathcal{Q}=\mathcal{H}$;
(ii) [Degeneracy] $\mathcal{Q}$ is degenerate, i.e., $\mathsf{ker}\mathcal{Q}\backslash\{0\}\neq\varnothing$;
(iii) [Closedness] $\mathcal{Q}$ is closed, i.e., $\overline{\mathsf{ran}\mathcal{Q}}=\mathsf{ran}\mathcal{Q}$.
Remark 1
A simple criterion for the closedness is: $\exists\alpha>0$, such that $\|x\|_{\mathcal{Q}}^{2}\geq\alpha\|x\|^{2},\ \forall x\in\mathsf{ran}\mathcal{Q}$ [4, Fact 2.26].
A few basic properties of $\mathcal{Q}$ are listed here.
Fact 1
Under Assumption 1, the following hold.
(i) [4, Fact 2.25] An orthogonal decomposition of $\mathcal{H}$ is: $\mathcal{H}=\mathsf{ran}\mathcal{Q}+\mathsf{ker}\mathcal{Q}$.
(ii) [4, Proposition 3.30] Any point $x\in\mathcal{H}$ can be decomposed as $x=x_{\mathsf{r}}+x_{\mathsf{k}}$, where $x_{\mathsf{r}}:=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x$ and $x_{\mathsf{k}}:=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}x$ are the $\mathsf{ran}\mathcal{Q}$- and $\mathsf{ker}\mathcal{Q}$-parts of $x$.
(iii) [9, Theorem 4] $\mathcal{Q}$ admits an (actually unique) self-adjoint square-root, denoted as $\sqrt{\mathcal{Q}}$, such that $\mathcal{Q}=\sqrt{\mathcal{Q}}\sqrt{\mathcal{Q}}$.
2 Basic properties under restricted monotonicity
In this part, we first introduce a restricted version of monotonicity, from which follow several well-known properties such as firm nonexpansiveness and demiclosdedness in terms of $\mathcal{Q}$-norm. These preliminary results will be essential for Sect. 35 later.
2.1 Restricted monotonicity
Observing (2), the monotonicity of $\mathcal{A}$ should be associated with the degeneracy of $\mathcal{Q}$, which gives rise to the following important definition.
Definition 2.1 (restricted monotonicity).
Under the condition of $\mathsf{ran}\mathcal{A}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing$, $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ is monotone, if
$\langle x-y|u-v\rangle\geq 0,\ \forall(x,u),(y,v)\in\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q}).$
The restricted monotonicity is equivalent to the $\mathcal{M}$-monotonicity [10, Proposition 2.5] and also closely related to the local monotonicity [4, Definition 25.7], which requires $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times C)$ to be monotone with the set $C$ being open. The essential difference between the restricted and local versions of monotonicity lies in that $\mathsf{int}\ \mathsf{ran}\mathcal{Q}$ could be empty while $\mathsf{int}\ C$ is always nonempty.
It is well-known that $\mathcal{A}$ is monotone if and only if $\mathcal{A}^{-1}$ is monotone [4, Proposition 20.22]. This fact can be extended directly as below and will provide us with a new perspective for studying the well-definedness of $\mathcal{T}$.
Proposition 2.2.
$\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ is monotone, if and only if $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone.
For convenience of later references, we put Definition 2.1 as
Assumption 2
(i) [Basic viability] $\mathsf{ran}\mathcal{A}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing$;
(ii) [Restricted monotonicity] $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ is monotone.
2.2 Basic properties
First, it is easy to recognize
Fact 1.
Given $\mathcal{T}$ as (2) under Assumption 1, the following hold:
(i) $\mathcal{Q}=\mathcal{Q}\circ\mathcal{P}_{\mathsf{ran}\mathcal{Q}}=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{Q}$;
(ii) $\mathcal{T}\circ\mathcal{P}_{\mathsf{ran}\mathcal{Q}}=\mathcal{T}$, i.e., $\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}=\mathcal{T}$.
Now, we show several basic properties on $\mathcal{T}$ based on Assumption 2 and Fact 1. The first is the single-valuedness of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$.
Lemma 2.3.
Under Assumptions 1 and 2, $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ is single-valued.
Proof 2.4.
Let $x\in\mathsf{dom}\mathcal{T}$ and assume that $y_{1}\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$ and $y_{2}\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$. Then, we have
$\mathcal{Q}(x-y_{1})\in\mathcal{A}y_{1},\quad\mathcal{Q}(x-y_{2})\in\mathcal{A}y_{2}.$ (5)
Thus, by Assumptions 1 and 2, we derive that
00 0 $\geq$ $-\|y_{1}-y_{2}\|_{\mathcal{Q}}^{2}=\big{\langle}y_{1}-y_{2}|\mathcal{Q}(y_{2}-y_{1})\big{\rangle}\quad$
$=$ $\big{\langle}y_{1}-y_{2}|\mathcal{Q}(x-y_{1})-\mathcal{Q}(x-y_{2})\big{\rangle}\geq 0,$
which yields $\|y_{1}-y_{2}\|_{\mathcal{Q}}=0$, i.e., $\mathcal{Q}y_{1}=\mathcal{Q}y_{2}$, or equivalently, $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}y_{1}=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}y_{2}$.
[7, Proposition 4.2] and [11, Lemma 3.2] showed that $\mathcal{T}$ is $\mathcal{Q}$-firmly nonexpansive under maximally monotone $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ and non-degenerate $\mathcal{Q}$. This was also discussed in [32, Sect. 2]. As emphasized in [10, Lemma 2.6], this key property is preserved under restricted monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ for degenerate case.
Fact 2.
[10, Lemma 2.6] Under Assumptions 1 and 2, $\mathcal{T}$ given as (2) is $\mathcal{Q}$-firmly nonexpansive, i.e.,
$\big{\|}\mathcal{T}x-\mathcal{T}y\big{\|}_{\mathcal{Q}}^{2}\leq\big{\langle}\mathcal{T}x-\mathcal{T}y\big{|}\mathcal{Q}(x-y)\big{\rangle},\quad\forall x,y\in\mathsf{dom}\mathcal{T},$
or equivalently,
$\big{\|}\mathcal{T}x-\mathcal{T}y\big{\|}_{\mathcal{Q}}^{2}+\big{\|}(\mathcal{I}-\mathcal{T})x-(\mathcal{I}-\mathcal{T})y\big{\|}_{\mathcal{Q}}^{2}\leq\big{\|}x-y\big{\|}_{\mathcal{Q}}^{2},\quad\forall x,y\in\mathsf{dom}\mathcal{T}.$ (6)
Remark 2.5.
There is still a slight difference between our result and [10, Lemma 2.6]. The latter required $\mathcal{T}$ to be well-defined, which, however, is not assumed in our result. Nevertheless, we clarify that all the expressions in above inequalities are well-defined, even if $\mathcal{T}$ is multi-valued. Indeed, $\|\mathcal{T}x\|_{\mathcal{Q}}=\|\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{T}x\|_{\mathcal{Q}}$ and $\langle\mathcal{T}x|\mathcal{Q}y\rangle=\langle\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{T}x|\mathcal{Q}y\rangle$, $\forall(x,y)\in\mathsf{dom}\mathcal{T}\times\mathcal{H}$. They are well-defined, since $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ is single-valued (by Lemma 2.3). To put it more explicitly, combining with Fact 1, one can equivalently rewrite (6) as $\forall x_{\mathsf{r}},y_{\mathsf{r}}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{dom}\mathcal{T})$:
$\big{\|}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}})-(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(y_{\mathsf{r}})\big{\|}_{\mathcal{Q}}^{2}+\big{\|}(\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}})-(\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(y_{\mathsf{r}})\big{\|}_{\mathcal{Q}}^{2}\leq\big{\|}x_{\mathsf{r}}-y_{\mathsf{r}}\big{\|}_{\mathcal{Q}}^{2}.$ (7)
The concept of demiclosedness [4, Definition 4.26] and Browder’s demiclosedness principle [4, Theorem 4.27] are crucial for convergence analysis, see [11, Lemma 3.4, Theorem 3.1] for example. Here it is natural (or evenly, trivial) to extend the results to any $\mathcal{Q}$-firmly nonexpansive operator $\mathcal{S}$.
Definition 2.6 ($\mathcal{Q}$-demiclosedness).
Under Assumption 1, let $D$ be a nonempty weakly sequentially closed subset of $\mathsf{ran}\mathcal{Q}$, let $\mathcal{S}:D\mapsto\mathsf{ran}\mathcal{Q}$, and let $u\in\mathsf{ran}\mathcal{Q}$. Then $\mathcal{S}$ is $\mathcal{Q}$-demiclosed at $u$ if, for every sequence $\big{\{}x^{k}\big{\}}_{k\in\mathbb{N}}$ in $D$ and every $x\in D$ such that $x^{k}\rightharpoonup x$ and $\mathcal{S}x^{k}\rightarrow u$, we have $\mathcal{S}x=u$. In addition, $\mathcal{S}$ is $\mathcal{Q}$-demiclosed if it is demiclosed at every point in $D$.
Lemma 2.7 ($\mathcal{Q}$-demiclosedness principle).
Under Assumption 1, let $D$ be a nonempty weakly sequentially closed subset of $\mathsf{ran}\mathcal{Q}$ and let $\mathcal{S}:D\mapsto\mathsf{ran}\mathcal{Q}$ be a $\mathcal{Q}$-nonexpansive operator, i.e.,
$\big{\|}\mathcal{S}x_{1}-\mathcal{S}x_{2}\big{\|}_{\mathcal{Q}}\leq\big{\|}x_{1}-x_{2}\big{\|}_{\mathcal{Q}},\quad\forall x_{1},x_{2}\in D\subseteq\mathsf{ran}\mathcal{Q}\cap\mathsf{dom}\mathcal{S}.$
Then $\mathcal{I}-\mathcal{S}$ is $\mathcal{Q}$-demiclosed.
[…]
Lemma 2.8 ($\mathcal{Q}$-demiclosedness of $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}$).
Given $\mathcal{T}$ as (2) satisfying Assumptions 1 and 2, let $D$ be a nonempty closed subset of $\mathsf{ran}\mathcal{Q}\cap\mathsf{dom}\mathcal{T}$, then $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}:D\mapsto\mathsf{ran}\mathcal{Q}$ is $\mathcal{Q}$-demiclosed.
Proof 2.9.
In view of Lemma 2.7, it suffices to show that $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}:D\mapsto\mathsf{ran}\mathcal{Q}$ is $\mathcal{Q}$-nonexpansive. Indeed, this immediately follows from Fact 2.
2.3 Generalized Moreau’s decomposition identity
Lemma 2.10 develops some equivalent forms of $\mathcal{T}$, which generalize the classic Moreau’s decomposition identity [4, Propositioin 23.20], and will be used for proving Theorem 3.1. The proof is postponed to Appendix A.
Lemma 2.10.
Under Assumptions 1 and 2, the following equalities hold:
$\mathcal{Q}-\mathcal{Q}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}$
[…]
where $\sqrt{\mathcal{Q}}$ is defined as Fact 1-(iii).
3 Full domain
3.1 Main result
Theorem 3.1 provides a sufficient and necessary condition on $\mathcal{A}$ and $\mathcal{Q}$ for the full domain of $\mathcal{T}$.
Theorem 3.1.
Under Assumption 1, $\mathcal{T}$ given as (2) has full domain, if and only if $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ (or equivalently saying, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$) $:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ is maximally monotone.
Let us first discuss the maximal monotonicity of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$, from which then readily follows the proof of Theorem 3.1. This concept is essentially a restricted version of maximal monotonicity from $\mathcal{H}$ to $2^{\mathcal{H}}$, by confining both domain and range of the operator to $\mathsf{ran}\mathcal{Q}$. Indeed, this restricted maximal monotonicity is closely related to Definition 2.1, due to the following lemma.
Lemma 3.2.
The following 3 statements are equivalent.
(i) $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ is monotone;
(ii) $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone;
(iii) $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone.
Proof 3.3.
(i)$\Longleftrightarrow$(ii): Proposition 2.2.
(ii)$\Longrightarrow$(iii): If $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone, then we have:
$\langle u-v|x-y\rangle\geq 0,\ \forall(u,x)\in\mathsf{gra}(\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}),\ \forall(v,y)\in\mathsf{gra}(\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}).$
Note that $\langle u-v|x-y\rangle=\langle u-v|x_{\mathsf{r}}-y_{\mathsf{r}}\rangle$, since $u,v\in\mathsf{ran}\mathcal{Q}$. This is equivalent to saying that
$\langle u-v|x_{\mathsf{r}}-y_{\mathsf{r}}\rangle\geq 0,\ \forall(u,x_{\mathsf{r}})\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}),\ \forall(v,y_{\mathsf{r}})\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}).$
This shows that $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone.
(iii)$\Longrightarrow$(ii): If $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone, then we have:
$\langle u-v|\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}y\rangle\geq 0,\ \forall(u,x)\in\mathsf{gra}(\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}),\ \forall(v,y)\in\mathsf{gra}(\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}).$
Note that $\langle u-v|\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}y\rangle=\langle\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(u-v)|x-y\rangle=\langle u-v|x-y\rangle\geq 0$, since $u,v\in\mathsf{ran}\mathcal{Q}$. This shows that $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone.
It is natural to develop a restricted version of classic Minty’s theorem [4, Theorem 21.1] for the restricted maximality of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$.
Theorem 3.4 (Restricted Minty’s theorem).
Let $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ be monotone. Then $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, if and only if $\mathsf{ran}(\mathcal{I}+\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}})=\mathsf{ran}\mathcal{Q}$.
The proof is exactly the same as [4, Theorem 21.1], if the subspace $\mathsf{ran}\mathcal{Q}$ is regarded as the ambient space. Actually, our specific $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ in Theorem 3.4 can be replaced by any monotone operator $\mathcal{B}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$, of course, including $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$.
We are now ready to prove Theorem 3.1.
Proof 3.5.
Combining Lemma 2.10-(4) and Theorem 3.4, we claim that $\mathcal{T}$ has full domain, if and only if $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone. What remains to show is the equivalence of maximal monotonicity between $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ and $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$.
First, if $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, then for any given $(u,x)\in\mathsf{ran}\mathcal{Q}\times\mathsf{ran}\mathcal{Q}$, we have:
$(u,x)\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}})\Longleftrightarrow\forall(v,y)\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}),\ \langle u-v|x-y\rangle\geq 0.$
Since $u,x,v,y\in\mathsf{ran}\mathcal{Q}$, let $u=\sqrt{\mathcal{Q}}u^{\prime}$ and $v=\sqrt{\mathcal{Q}}v^{\prime}$ for $u^{\prime},v^{\prime}\in\mathsf{ran}\mathcal{Q}$. Then, $x\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}(u)=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}u=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}u^{\prime}$, and similarly, $y\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}v^{\prime}$. Then, we obtain
$(u^{\prime},x)\in\mathsf{gra}\Big{(}\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}\Longleftrightarrow\forall(v^{\prime},y)\in\mathsf{gra}\Big{(}\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)},\ \langle\sqrt{\mathcal{Q}}(u^{\prime}-v^{\prime})|x-y\rangle\geq 0.$
Let $x^{\prime}=\sqrt{\mathcal{Q}}x$ and $y^{\prime}=\sqrt{\mathcal{Q}}y$, then it becomes
$(u^{\prime},x^{\prime})\in\mathsf{gra}\Big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}\Longleftrightarrow\forall(v^{\prime},y^{\prime})\in\mathsf{gra}\Big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)},\ \langle u^{\prime}-v^{\prime}|x^{\prime}-y^{\prime}\rangle\geq 0.$
This shows that $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone.
Conversely, if $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, then for any given $(u,x)\in\mathsf{ran}\mathcal{Q}\times\mathsf{ran}\mathcal{Q}$, we have:
$(u,x)\in\mathsf{gra}(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}})\Longleftrightarrow\forall(v,y)\in\mathsf{gra}(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}),\ \langle u-v|\sqrt{\mathcal{Q}}x-\sqrt{\mathcal{Q}}y\rangle\geq 0.$
Let $x^{\prime}=\sqrt{\mathcal{Q}}x$, $y^{\prime}=\sqrt{\mathcal{Q}}y$, $u^{\prime}=\sqrt{\mathcal{Q}}u$, $v^{\prime}=\sqrt{\mathcal{Q}}v$, we have:
$(u^{\prime},x^{\prime})\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}})\Longleftrightarrow\forall(v^{\prime},y^{\prime})\in\mathsf{gra}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}),\ \langle u^{\prime}-v^{\prime}|x^{\prime}-y^{\prime}\rangle\geq 0.$
This shows the maximal monotonicity of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$.
Remark 3.6.
According to Lemma 3.2, Assumption 2 has already been implied by the maximal monotonicity of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$. Thus, there is no need to put Assumption 2 in Theorem 3.1.
3.2 Sufficient condition
The next question regarding Theorem 3.1 naturally arises: what is the condition on $\mathcal{A}$ and $\mathcal{Q}$, for which $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ is maximally monotone? The following important corollary gives a sufficient condition.
Corollary 3.7.
Under Assumption 1, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, if the following conditions are fulfilled:
(i) $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is maximally monotone;
(ii) $0\in\mathsf{sri}(\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A})$ [^1] The definition of strong relative interior $(\mathsf{sri})$ is given by [4, Definition 6.9]..
In particular, if $\mathcal{Q}$ is non-degenerate, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, if and only if $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is maximally monotone.
Proof 3.8.
In view of [4, Corollary 25.6 or Proposition 25.41-(iv)].
For convenience of reference, we put Theorem 3.1 and Corollary 3.7 as the following two sets of assumptions, respectively.
Assumption 3
$\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ (or $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$): $\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ is maximally monotone.
Assumption 4
(i) $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is maximally monotone;
(ii) $0\in\mathsf{sri}(\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A})$.
[…]
$\mathcal{A}=\begin{bmatrix}\partial f&0\\ -B&\partial g\end{bmatrix}:\mathcal{X}\times\mathcal{Y}\mapsto 2^{\mathcal{X}}\times 2^{\mathcal{Y}},\quad\mathcal{Q}=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}:\mathcal{X}\times\mathcal{Y}\mapsto\mathcal{X}\times\{0\},$
[…]
In addition, without Assumption 4-(ii), we claim that $\mathcal{T}$almost’ has full domain.
Proposition 3.9.
Under Assumptions 1 and 4-(i), $\overline{\mathsf{ran}}(\mathcal{A}+\mathcal{Q})\supseteq\mathsf{ran}\mathcal{Q}$.
Proof 3.10.
First, $\mathcal{Q}$ is maximally monotone by [4, Corollary 20.28], and $3^{*}$ monotone by Fact 1-(iv) and [4, Example 25.17]. Then, $\mathcal{A}+\mathcal{Q}$ is maximally monotone by [4, Corollary 25.5].
Then, noting $\mathsf{dom}\mathcal{A}\subseteq\mathsf{dom}\mathcal{Q}=\mathcal{H}$, [4, Theorem 25.24] yields
$\overline{\mathsf{ran}}(\mathcal{A}+\mathcal{Q})=\overline{\mathsf{ran}\mathcal{A}+\mathsf{ran}\mathcal{Q}}\supseteq\overline{0+\mathsf{ran}\mathcal{Q}}=\overline{\mathsf{ran}}\mathcal{Q}=\mathsf{ran}\mathcal{Q},$
where the last equality is due to the closedness of $\mathsf{ran}\mathcal{Q}$.
The above result always holds, without Assumption 4-(ii). One can check Example 3.12, where $\mathsf{ran}(\mathcal{A}+\mathcal{Q})\nsupseteq\mathsf{ran}\mathcal{Q}$ but $\overline{\mathsf{ran}}(\mathcal{A}+\mathcal{Q})=\mathbb{R}\times\mathbb{R}_{+}\supset\mathbb{R}\times\{0\}=\mathsf{ran}\mathcal{Q}$. Unfortunately, it is generally difficult to investigate the relation between $\overline{\mathsf{ran}}(\mathcal{A}+\mathcal{Q})$ and $\mathsf{ran}(\mathcal{A}+\mathcal{Q})$. We believe there are some connections between Assumption 4-(ii) and $\overline{\mathsf{ran}}(\mathcal{A}+\mathcal{Q})$. This is left to future work.
3.3 Examples
We list several examples to show
(1) $\mathcal{A}+\mathcal{Q}$ is often non-surjective;
(2) $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ may not be maximally monotone, even if $\mathcal{A}$ is maximally monotone;
(3) Assumption 4 is tight, but not necessary for the full domain of $\mathcal{T}$.
Example 3.11.
Consider a toy example
$\mathcal{H}=\mathbb{R}^{2},\quad\mathcal{A}=\partial f,\text{\ where\ }f(x,y)=|x|,\quad\mathcal{Q}=\begin{bmatrix}0&0\\ 0&1\end{bmatrix}.$
It is easy to see that $\mathsf{ran}\mathcal{A}=[-1,1]\times\{0\}$, $\mathsf{ran}\mathcal{Q}=\{0\}\times\mathbb{R}$, $\mathsf{ran}(\mathcal{A}+\mathcal{Q})=[-1,1]\times\mathbb{R}\neq\mathcal{H}=\mathbb{R}^{2}$, but it already suffices to guarantee that $\mathsf{ran}(\mathcal{A}+\mathcal{Q})\supset\mathsf{ran}\mathcal{Q}$. One further has $\mathsf{ran}\mathcal{A}-\mathsf{ran}\mathcal{Q}=[-1,1]\times\mathbb{R}$, which contains 0 as a relative interior.
Is $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ maximally monotone? First, noting that $\mathsf{dom}\mathcal{A}^{-1}=\mathsf{ran}\mathcal{A}=[-1,1]\times\{0\}$, $\mathcal{A}^{-1}$ is given as:
$\mathcal{A}^{-1}:(x,y)\mapsto\left\{\begin{array}[]{ll}\mathbb{R}_{+}\times\mathbb{R},&\textrm{if\ }x=1,\ y=0;\\ \mathbb{R}_{-}\times\mathbb{R},&\textrm{if\ }x=-1,\ y=0;\\ \{0\}\times\mathbb{R},&\textrm{if\ }x\in(-1,1),\ y=0.\\ \end{array}\right.$
where +=$\mathbb{R}_{+}=\{x\in\mathbb{R}:x\geq 0\}$, =$\mathbb{R}_{-}=\{x\in\mathbb{R}:x\leq 0\}$.
We now check whether $\mathsf{ran}(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})=\{0\}\times\mathcal{R}=\mathsf{ran}\mathcal{Q}$. Indeed, we develop:
$\mathsf{ran}\Big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}$ $\supseteq$ $\Big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}\begin{bmatrix}0\\ 0\end{bmatrix}$
$=$ $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\begin{bmatrix}0\\ 0\end{bmatrix}=\sqrt{\mathcal{Q}}(\{0\}\times\mathbb{R})$
$=$ $\{0\}\times\mathbb{R}=\mathsf{ran}\mathcal{Q},$
which shows $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone by restricted Minty’s theorem.
Example 3.12.
In [10, Remark 2.2], the authors proposed an interesting counter-example that deserves particular treatment. The problem is
$\mathcal{H}=\mathbb{R}^{2},\mathcal{A}=\partial f,\text{\ where\ }f(x,y)=\max\{e^{y}-x,0\}\text{\ and\ }\mathcal{Q}=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}.$
Based on basic subdifferential calculus, one obtains $\mathcal{A}$ given as
$\mathcal{A}:(x,y)\mapsto\left\{\begin{array}[]{ll}(-1,e^{y}),&\textrm{if\ }y>\ln x,x>0\textrm{\ or\ }x\leq 0,\forall y\in\mathbb{R};\\ (0,0),&\textrm{if\ }y<\ln x,x>0;\\ {}[-1,0]\times[0,e^{y}],&\textrm{if\ }y=\ln x,x>0.\end{array}\right.$
Thus, $\mathcal{A}+\mathcal{Q}$ is given as
$\mathcal{A}+\mathcal{Q}:(x,y)\mapsto\left\{\begin{array}[]{ll}(x-1,e^{y}),&\textrm{if\ }y>\ln x,x>0\textrm{\ or\ }x\leq 0,\forall y\in\mathbb{R};\\ (x,0),&\textrm{if\ }y<\ln x,x>0;\\ (u,v)\in[x-1,x]\times[0,e^{y}],&\textrm{if\ }y=\ln x,x>0.\end{array}\right.$
We then deduce that $\mathsf{ran}\mathcal{A}=[-1,0]\times\mathbb{R}_{+}$, $\mathsf{ran}(\mathcal{A}+\mathcal{Q})=(-\infty,-1]\times(0,+\infty)\cup(-1,+\infty]\times[0,+\infty)=\mathbb{R}\times[0,+\infty)\backslash(-\infty,-1]\times\{0\}$. Clearly, $\mathsf{ran}(\mathcal{A}+\mathcal{Q})\nsupseteq\mathsf{ran}\mathcal{Q}=\mathbb{R}\times\{0\}$. Note that $(\mathcal{A}+\mathcal{Q})^{-1}(x,0)=\varnothing$ for $x\leq-1$, though $(x,0)\in\mathsf{ran}\mathcal{Q}$.
Since $\mathcal{T}$ is not of full domian, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ must not be maximally monotone by Theorem 3.1. Let us check it. First, we have $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}:(x,0)\mapsto\mathbb{R}_{++}\times\{0\}$, $\forall x\in[-1,0]$, where ++=$\mathbb{R}_{++}=\{x\in\mathbb{R}:x>0\}$. Then, $\mathsf{ran}(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})=(-1,+\infty)\times\{0\}\neq\mathbb{R}\times\{0\}$, and $(-\infty,-1]\nsubseteq\mathsf{ran}(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})$. Thus, by Theorem 3.4, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is not maximally monotone.
Finally, it is easy to check that Assumption 4-(ii) is not satisfied. Indeed, $\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A}=\mathbb{R}\times(-\infty,0]$, and thus, $0\notin\mathsf{sri}(\mathsf{ran}\mathcal{A}-\mathsf{ran}\mathcal{Q})$. This shows that Assumption 4 is tight.
Example 3.13.
We consider Example 3.12 but with the metric $\mathcal{Q}=\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$. Then $\mathcal{A}+\mathcal{Q}$ is given as
$\mathcal{A}+\mathcal{Q}:(x,y)\mapsto\left\{\begin{array}[]{ll}(1,e^{y}+y),&\textrm{if\ }y>\ln x,x>0\textrm{\ or\ }x\leq 0,\forall y\in\mathbb{R};\\ (0,y),&\textrm{if\ }y<\ln x,x>0;\\ {}[-1,0]\times[y,e^{y}+y],&\textrm{if\ }y=\ln x,x>0.\end{array}\right.$
We then deduce that $\mathsf{ran}(\mathcal{A}+\mathcal{Q})=[-1,0]\times\mathbb{R}$. Clearly, $\mathsf{ran}(\mathcal{A}+\mathcal{Q})\supseteq\mathsf{ran}\mathcal{Q}=\{0\}\times\mathbb{R}$. By Theorem 3.1, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ must be maximally monotone. Indeed, if the input $y=0$, we then have $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}:0\mapsto\{0\}\times\mathbb{R}$. Then, $\mathsf{ran}(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})\supseteq(\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})(0)=(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}})(0)=\{0\}\times\mathbb{R}=\mathsf{ran}\mathcal{Q}$, and thus by restricted Minty’s theorem, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone.
However, $\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A}=[0,1]\times\mathbb{R}$, $0\notin\mathsf{sri}(\mathsf{ran}\mathcal{A}-\mathsf{ran}\mathcal{Q})$. This shows that the condition of $0\in\mathsf{sri}(\mathsf{ran}\mathcal{A}-\mathsf{ran}\mathcal{Q})$ is sufficient but not necessary condition for the full domain of $\mathcal{T}$.
4 Single-valuedness
Let us now study the condition for the single-valuedness of $\mathcal{T}$.
4.1 Restricted injectivity
As mentioned in Sect. 1.2, a standard condition for the single-valuedness of $\mathcal{T}$ is disjoint injectivity of $\mathcal{A}+\mathcal{Q}$ [5, Theorem 2.1-(ix)], [13, Proposition 3.8]. Moreover, a sufficient and necessary condition, proposed in [12, Proposition 2.1-(2)], is that $\mathcal{A}+\mathcal{Q}$ is injective on $\mathsf{ran}\mathcal{Q}$. Here, Theorem 4.2 claims that this condition can be further weakened to to the disjoint injectivity on $\mathsf{ran}\mathcal{Q}$ w.r.t. $\mathsf{ker}\mathcal{Q}$ only, under Assumptions 1 and 2.
Definition 4.1.
We say that $\mathcal{B}$ is disjointly injective on $\mathsf{ran}\mathcal{Q}$ w.r.t. $\mathsf{ker}\mathcal{Q}$, if
$\mathcal{B}y_{1}\cap\mathcal{B}y_{2}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing\Longrightarrow y_{1,\mathsf{k}}=y_{2,\mathsf{k}},\quad\forall y_{1},y_{2}\in\mathsf{dom}\mathcal{B}.$ (8)
Theorem 4.2.
Under Assumptions 1 and 2, for any given input $x_{\mathsf{r}}\in\mathsf{ran}\mathcal{Q}\cap\mathsf{dom}\mathcal{T}$, $y\in\mathcal{T}x_{\mathsf{r}}$ is single-valued, if and only if one of the following equivalent conditions holds:
(i) $\mathcal{A}+\mathcal{Q}$ is disjointly injective on $\mathsf{ran}\mathcal{Q}$ w.r.t. $\mathsf{ker}\mathcal{Q}$;
(ii) For any input $x_{\mathsf{r}}\in\mathsf{ran}\mathcal{Q}$, given the corresponding $y_{\mathsf{r}}=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}})$, the constrained inclusion problem of $\mathcal{A}(y_{\mathsf{r}}+y_{\mathsf{k}})\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$, s.t. $\mathcal{Q}y_{\mathsf{k}}=0$, has a unique solution of $y_{\mathsf{k}}$;
(iii) Suppose there exists a linear, bounded and surjective operator $\mathcal{E}:\mathcal{H}\mapsto D^{\prime}$, for some real Hilbert space $D^{\prime}$, such that $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}=\mathcal{E}^{\top}\mathcal{E}$. Then, for any input $x_{\mathsf{r}}\in\mathsf{ran}\mathcal{Q}$ given the corresponding $y_{\mathsf{r}}=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}})$, the inclusion problem $\mathcal{A}(y_{\mathsf{r}}+\mathcal{E}^{\top}b)\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$ has a unique solution of $b\in D^{\prime}$;
(iv) $\mathsf{ker}\mathcal{A}\cap(\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})-y)=\{0\}$.
Proof 4.3.
[…]
Single-$\Longrightarrow$(i): Let us prove (8) by contradiction. If (8) does not hold, it implies that $\exists y_{1},y_{2}\mathsf{dom}\mathcal{A}$, such that $(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing$, and $y_{1,\mathsf{k}}\neq y_{2,\mathsf{k}}$. Then, take $x$, such that $\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}$, i.e., $\{y_{1},y_{2}\}\subseteq\mathcal{T}x$. It has been shown in Lemma 2.3 that $y_{1,\mathsf{r}}=y_{2,\mathsf{r}}$. Combining with $y_{1,\mathsf{k}}\neq y_{2,\mathsf{k}}$, we obtain $y_{1}\neq y_{2}$. This means that $\mathcal{T}$ is not single-valued, which is contradiction.
Single-$\Longleftrightarrow$(ii): We deduce
$y\in\mathcal{T}x\Longleftrightarrow\mathcal{A}(y_{\mathsf{r}}+y_{\mathsf{k}})=\mathcal{A}y\owns\mathcal{Q}(x-y)=\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}}),$ (9)
which results in
$y_{\mathsf{r}}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}}),\quad y_{\mathsf{k}}\in\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}}).$ (10)
Lemma 2.3 shows that $y_{\mathsf{r}}$ is uniquely determined by $x_{\mathsf{r}}$. Thus, the single-valuedness of $\mathcal{T}$ totally depends on the kernel part—$y_{\mathsf{k}}$, as given above. This completes the proof, noting that $x_{\mathsf{r}}$ and $y_{\mathsf{r}}$ are irrelevant to $y_{\mathsf{k}}$.
(ii)$\Longleftrightarrow$(iii): clear.
(i)$\Longleftrightarrow$(ii): Since $(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing$, pick a point $x$, such that
$\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})y_{1}\cap(\mathcal{A}+\mathcal{Q})y_{2}.$ (11)
Lemma 2.3 and (8) show that $y_{1,\mathsf{r}}=y_{2,\mathsf{r}}:=y_{\mathsf{r}}$ and $y_{1,\mathsf{k}}=y_{2,\mathsf{k}}:=y_{\mathsf{k}}$. On the other hand, (11) is equivalent to (9) for $y_{1}$ and $y_{2}$, respectively.
[…]
Corollary 4.4.
Under Assumptions 1 and 2, $\mathcal{T}$ is single-valued, if $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ is single-valued.
Proof 4.5.
(9) can be rewritten as $y_{\mathsf{r}}+y_{\mathsf{k}}\in\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$. Then, apply the projection $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}$ to both sides yields $y_{\mathsf{k}}\in\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$, which completes the proof.
Remark 4.6.
(i) The condition of Corollary 4.4 is sufficient, but not necessary. Actually, the single-valuedness of $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ is much stronger than Theorem 4.2-(iv): Even if $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ is multi-valued, $y_{\mathsf{k}}$ could still be uniquely determined by (9). This depends on the interaction between $\mathcal{A}$ and $\mathcal{Q}$.
(ii) There is no need to enforce $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$ to be single-valued, since the uniqueness of $y_{\mathsf{r}}$ has been guaranteed by Lemma 2.3. This is irrelevant to single-valuedness of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}$, which is, indeed, often multi-valued in practice.
(iii) A big merit of condition Theorem 4.2-(iii), compared to (ii), is that the unconstrained version is easier to use and solve.
(iv) Theorem 4.2-(iv) has similar spirit with the most recent result of [17, Theorem 3.1], for regularized optimization problems. Their potential connections deserve further explorations in future work.
We present the following result as summary for the well-definedness of $\mathcal{T}$.
Theorem 4.7.
Under Assumptions 1 and 2, $\mathcal{T}$ given as (2) is well-defined,
(i) if and only if Assumption 3 and any condition in Theorem 4.2 are fulfilled;
(ii) if Assumption 4 and the condition of Corollary 4.4 are satisfied.
4.2 More remarks and examples
The following guiding examples show how the single-valuedness of $\mathcal{T}$ depends on the choice of $\mathcal{Q}$. Given a function $f:\mathbb{R}^{2}\mapsto\mathbb{R}:(x_{1},x_{2})\mapsto|x_{1}|$, consider a minimization problem w.r.t. $x=(x_{1},x_{2})\in\mathbb{R}^{2}$:
$\min_{x\in\mathbb{R}^{2}}f(x)+\frac{1}{2}(x_{1}-b)^{2}\Longleftrightarrow 0\in\partial|\cdot|(x_{1}^{\star})+(x_{1}^{\star}-b),$ (15)
This is further written as the PPA form (2):
$\begin{bmatrix}0\\ 0\end{bmatrix}\in\begin{bmatrix}\partial|\cdot|&0\\ 0&0\end{bmatrix}\begin{bmatrix}x_{1}^{\star}\\ x_{2}^{\star}\end{bmatrix}+\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\begin{bmatrix}x_{1}^{\star}-b\\ x_{2}^{\star}-0\end{bmatrix}$
with the input variable $y=(b,0)$.
Obviously, the optimal solution set is $X^{\star}=\{x_{1}^{\star}\}\times\mathbb{R}$, where $x_{1}^{\star}$ is given by a well-known soft-thresholding function $S(\cdot)$ of $b$:
$x_{1}^{\star}=S(b):=\left\{\begin{array}[]{ll}b-1,&\textrm{if\ }b>1;\\ b+1,&\textrm{if\ }b<-1;\\ 0,&\textrm{if\ }|b|\leq 1.\end{array}\right.$
One can easily check that neither Theorem 4.2 nor Corollary 4.4 is satisfied. For instance,
$\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(y_{\mathsf{r}}-x^{\star}_{\mathsf{r}})=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(y_{\mathsf{r}}-x^{\star}_{\mathsf{r}})=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\begin{bmatrix}b-S(b)\\ 0\end{bmatrix}=\{0\}\times\mathbb{R},$
which is multi-valued.
However, given the same $f$, (15) is changed as:
$\min_{x\in\mathbb{R}^{2}}f(x)+\frac{1}{2}(x_{2}-b)^{2}\Longleftrightarrow 0\in\partial\partial|\cdot|(x_{1}^{\star})+x_{2}^{\star}-b.$
This corresponds to
$\begin{bmatrix}0\\ 0\end{bmatrix}\in\begin{bmatrix}\partial|\cdot|&0\\ 0&0\end{bmatrix}\begin{bmatrix}x_{1}^{\star}\\ x_{2}^{\star}\end{bmatrix}+\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\begin{bmatrix}x_{1}^{\star}-0\\ x_{2}^{\star}-b\end{bmatrix}$
with input $y=(0,b)$. The unique solution is obviously $x^{\star}=(0,b)$. On the other hand, it is easy to check that $\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}(y_{\mathsf{r}}-x^{\star}_{\mathsf{r}})=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}\mathcal{A}^{-1}\begin{bmatrix}0\\ 0\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix}$. This shows, by Corollary 4.4, the uniqueness of solution, $\forall b\in\mathbb{R}$.
5 Convergence analysis
5.1 Convergence in the range space
All the results presented above are concerned with the single step of (2) only. Now, we turn to the convergence analysis for the generated iterate sequence with the following assumption of existence of solutions.
Assumption 5
[Existence of solution to (1)] $\mathsf{zer}\mathcal{A}=\mathcal{A}^{-1}(0)\neq\varnothing$.
The following is an intermediate result from the proof of [22, Theorem 3.4], which is also crucial for our convergence analysis.
Lemma 5.1.
Under Assumptions 1, 2 and 5, given $\mathcal{T}$ as (2), then
$\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A}).$
Proof 5.2.
(1) $x\in\mathsf{zer}\mathcal{A}\Longrightarrow 0\in\mathcal{A}x\Longrightarrow\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})x\Longrightarrow x\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x\Longrightarrow\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x)=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x)\Longrightarrow\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x\in\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$. On the other hand, $x\in\mathsf{zer}\mathcal{A}\Longrightarrow\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$. Thus, we obtain $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})\subseteq\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$.
(2) Let $x\in\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})\Longrightarrow x=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})x\Longrightarrow\mathcal{Q}x=\mathcal{Q}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})x=\mathcal{Q}\mathcal{T}x$. On the other hand, $\mathcal{T}=(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\Longrightarrow 0\in\mathcal{A}(\mathcal{T}x)+\mathcal{Q}\mathcal{T}x-\mathcal{Q}x$, $\forall x\in\mathsf{dom}\mathcal{T}$. Combining both above, we then have $x\in\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})\Longrightarrow x\in\mathcal{A}(\mathcal{T}x)\Longrightarrow\mathcal{T}x\subseteq\mathsf{zer}\mathcal{A}\Longrightarrow(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x)\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$. Since $x=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x)$, we have $x\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$. Thus, $\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})\subseteq\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$.
Let us now discuss the convergence of (2). More concretely, we improve the weak convergence of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x^{k}\rightharpoonup\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x^{\star}$ established previously in [22, Theorem 3.4] by removing the finite dimensional restrict. For notational simplicity, we denote $x_{\mathsf{r}}^{k}:=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x^{k}$, $\forall k\in\mathbb{N}$.
Theorem 5.3.
[Weak convergence in $\mathrm{ran}\mathcal{Q}$] Under Assumptions 1, 3 and 5, let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by (2). Then $x_{\mathsf{r}}^{k}\rightharpoonup\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x^{\star}$ for some $x^{\star}\in\mathsf{zer}\mathcal{A}$, as $k\rightarrow\infty$.
Proof 5.4.
The proof is divided into 3 steps:
Step-1: For every $x_{\mathsf{r}}^{\star}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$, $\lim_{k\rightarrow\infty}$ $\|x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}}$ exists.
Step-2: $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ has at least one weak sequential cluster point lying in $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$;
Step-3: The weak sequential cluster point of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ is unique.
Step-1: First, we define $x_{\mathsf{r}}^{\star}\in\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$. We also have $x_{\mathsf{r}}^{\star}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$ by Lemma 5.1, which implies that $\exists x^{\star}\in\mathsf{zer}\mathcal{A}$, such that $x_{\mathsf{r}}^{\star}=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x^{\star}$. According to Fact 2, substituting $x=x_{\mathsf{r}}^{k}$ and $y=x_{\mathsf{r}}^{\star}$ into (7) yields
$\big{\|}x_{\mathsf{r}}^{k+1}-x_{\mathsf{r}}^{\star}\big{\|}_{\mathcal{Q}}^{2}\leq\big{\|}x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\big{\|}_{\mathcal{Q}}^{2}-\big{\|}x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{k+1}\|_{\mathcal{Q}}^{2}.$ (16)
(16) guarantees that $\{\|x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}}\}_{k\in\mathbb{N}}$ is non-increasing, and bounded from below (always being non-negative), and thus, convergent, i.e., $\lim_{k\rightarrow\infty}$ $\|x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}}$ exists.
Step-2: (16) implies the boundedness of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$. Indeed, from (16) and the closedness of $\mathsf{ran}\mathcal{Q}$, there exists $\alpha>0$, such that
$\alpha\big{\|}x_{\mathsf{r}}^{k}\big{\|}-\alpha\big{\|}x_{\mathsf{r}}^{\star}\big{\|}\leq\alpha\big{\|}x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\big{\|}\leq\big{\|}x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\big{\|}_{\mathcal{Q}}\leq\big{\|}x_{\mathsf{r}}^{0}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}},\quad\forall k\in\mathbb{N}.$
Thus, $\big{\|}x_{\mathsf{r}}^{k}\big{\|}\leq\big{\|}x_{\mathsf{r}}^{\star}\big{\|}+\frac{1}{\alpha}\big{\|}x_{\mathsf{r}}^{0}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}}:=\rho$, $\forall k\in\mathbb{N}$, and $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}\subseteq B(0,\rho)\cap\mathsf{ran}\mathcal{Q}$. $B(0,\rho)$ is weakly sequentially compact (by [4, Lemma 2.45]), and also is $B(0,\rho)\cap\mathsf{ran}\mathcal{Q}\subseteq B(0,\rho)$ (by [4, Lemma 1.34]). Then, $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ must possess at lease one weak sequential cluster point $v^{*}\in B(0,\rho)\cap\mathsf{ran}\mathcal{Q}$, i.e., there exists a subsequence $\{x_{\mathsf{r}}^{k_{i}}\}_{k\in\mathbb{N}}$ that weakly converges to $v^{*}$, as $k_{i}\rightarrow\infty$. We here need to show that $v^{*}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$, and more generally, every weak sequential cluster point of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ belongs to $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$. To this end, summing up (16) from $k=0$ to $K-1$, and taking $K\rightarrow\infty$, we have
$\sum_{k=0}^{\infty}\big{\|}x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{k+1}\big{\|}_{\mathcal{Q}}^{2}\leq\big{\|}x_{\mathsf{r}}^{0}-x_{\mathsf{r}}^{\star}\big{\|}_{\mathcal{Q}}^{2}<+\infty.$
This implies that $x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{k+1}=(\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}}^{k})\rightarrow 0$, and also $(\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{\mathsf{r}}^{k_{i}})\rightarrow 0$ as $k_{i}\rightarrow\infty$. Combining with $x_{\mathsf{r}}^{k_{i}}\rightharpoonup v^{*}$ and the $\mathcal{Q}$-demiclosedness of $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ (shown in Lemma 2.8), we obtain $(\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(v^{*})=0\Longrightarrow v^{*}\in\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$. Since $\{x_{\mathsf{r}}^{k_{i}}\}_{k_{i}\in\mathbb{N}}$ is an arbitrary weakly convergence subsequence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$, we conclude that every weak sequential cluster point of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$, and furthermore, in $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$ as well, again by Lemma 5.1.
Step-3: We need to show that $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ cannot have two distinct weak sequential cluster point in $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$. Indeed, let $v_{1}^{*},v_{2}^{*}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$ be two cluster points of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$. Since $\lim_{k\rightarrow\infty}\|x_{\mathsf{r}}^{k}-x_{\mathsf{r}}^{\star}\|_{\mathcal{Q}}$ exists as proved in Step-1, set $l_{1}=\lim_{k\rightarrow\infty}\|x_{\mathsf{r}}^{k}-v_{1}^{*}\|_{\mathcal{Q}}$, and $l_{2}=\lim_{k\rightarrow\infty}\|x_{\mathsf{r}}^{k}-v_{2}^{*}\|_{\mathcal{Q}}$. Take a subsequence $\{x_{\mathsf{r}}^{k_{i}}\}_{k_{i}\in\mathbb{N}}$ weakly converging to $v_{1}^{*}$, as $k_{i}\rightarrow\infty$. From the identity:
$\big{\|}x_{\mathsf{r}}^{k_{i}}-v_{1}^{*}\big{\|}_{\mathcal{Q}}^{2}-\big{\|}x_{\mathsf{r}}^{k_{i}}-v_{2}^{*}\big{\|}_{\mathcal{Q}}^{2}=\big{\|}v_{1}^{*}-v_{2}^{*}\big{\|}_{\mathcal{Q}}^{2}+2\big{\langle}v_{1}^{*}-v_{2}^{*}\big{|}v_{2}^{*}-x_{\mathsf{r}}^{k_{i}}\big{\rangle}_{\mathcal{Q}},$
taking $k_{i}\rightarrow\infty$ on both sides, the last term becomes:
$\lim_{k_{i}\rightarrow\infty}\big{\langle}v_{1}^{*}-v_{2}^{*}\big{|}v_{2}^{*}-x_{\mathsf{r}}^{k_{i}}\big{\rangle}_{\mathcal{Q}}=\lim_{k_{i}\rightarrow\infty}\big{\langle}v_{1}^{*}-v_{2}^{*}\big{|}v_{2}^{*}-x_{\mathsf{r}}^{k_{i}}\big{\rangle}_{\mathcal{Q}}=-\|v_{1}^{*}-v_{2}^{*}\|_{\mathcal{Q}}^{2},$
then, we deduce that $l_{1}^{2}-l_{2}^{2}=-\big{\|}v_{1}^{*}-v_{2}^{*}\big{\|}_{\mathcal{Q}}^{2}$. Similarly, take a subsequence $\{x^{k_{j}}\}_{k_{j}\in\mathbb{N}}$ weakly converging to $v_{2}^{*}$, as $k_{j}\rightarrow\infty$, which yields that $l_{1}^{2}-l_{2}^{2}=\big{\|}v_{1}^{*}-v_{2}^{*}\big{\|}_{\mathcal{Q}}^{2}$. Consequently, $\big{\|}v_{1}^{*}-v_{2}^{*}\big{\|}_{\mathcal{Q}}=0\Longrightarrow v_{1}^{*}=v_{2}^{*}$, since $v_{1}^{*},v_{2}^{*}\in\mathsf{ran}\mathcal{Q}$. This shows the uniqueness of the weak sequential cluster point.
Finally, combining the above 3 step and [4, Lemma 2.46], we conclude the weak convergence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$, and denote the weak limit as $x_{\mathsf{r}}^{\star}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A})$.
5.2 Discussions and related works
There are lots of technical details in the proof of Theorem 5.3 that need to be clarified, discussed and further compared to other related works.
First, Lemma 5.1 makes the connection between the degenerate PPA (2) and the original inclusion problem (1). One can always show the weak convergence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ in $\mathsf{Fix}(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})$ without Lemma 5.1. This is Lemma 5.1 that tells us what the degenerate PPA (2) solves for is actually a point of $\mathsf{zer}\mathcal{A}$ projected onto the range space of degenerate metric.
Assumption 4—a stronger condition than Assumption 3—is not needed in Theorem 5.3, and thus, $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is also not necessarily maximally monotone here. The single-valued conditions listed in Theorem 4.2 are also not needed. In other words, $\mathcal{T}$ is allowed to be multi-valued, since we are only concerned with $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$, which is already well-defined by Assumptions 1 and 3. This is fundamentally different from [22, Theorem 3.4], where $\mathcal{T}$ has to be well-defined.
[22, Theorem 3.4] discussed the convergence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ in the finite-dimensional setting, where the proof heavily relied on the assumption of continuity of $\mathcal{T}$. We here argue that the continuity of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ would suffice for [22, Theorem 3.4], which is obviously true in finite-dimensional case by Fact 2. However, $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}$ is not weakly continuous generally in infinite-dimensional case. Instead, we exploited the weak-to-strong continuity of $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ in Lemma 2.8.
[…]
Our proof uses the $\mathcal{Q}$-demiclosedness of $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}$. An alternative way is to use the maximal monotonicity of $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{|}_{\mathsf{ran}\mathcal{Q}}$ (i.e., Assumption 3) to finish the proof as well, see Appendix B.
Finally, we stress that $\{x^{k}\}_{k\in\mathbb{N}}$ may be unbounded due to the uncontrolled component of $x_{\mathsf{k}}^{k}:=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}x^{k}$. Thus, $\{x^{k}\}_{k\in\mathbb{N}}$ may not even possess any weak sequential cluster point, and there is no conclusion about the convergence of $\{x^{k}\}_{k\in\mathbb{N}}$. The only exception occurs in the non-degenerate case, where the weak convergence of $\{x_{\mathsf{r}}^{k}\}_{k\in\mathbb{N}}$ is obviously equivalent to that of $\{x^{k}\}_{k\in\mathbb{N}}$. Assumption 3 in this case boils down to the usual maximal monotonicity of $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$, as stated in Corollary 3.7.
5.3 Convergence in the whole space
To proceed further, if $\mathcal{T}$ is well-defined, it is then natural to ask when the iterations (2) converge in the whole space. At least, the boundedness of $\{x^{k}\}_{k\in\mathbb{N}}$ should be a basic requirement. To this end, [10, Theorem 2.9] assumed the Lipschitz continuity of $(\mathcal{A}+\mathcal{Q})^{-1}$. Here, we propose the equivalent conditions as
Assumption 6
For any two inputs $x_{i,\mathsf{r}}\in\mathsf{ran}\mathcal{Q}$ for $i=1$ and 2, and the corresponding $y_{i,\mathsf{r}}=(\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T})(x_{i,\mathsf{r}})$,
(i) the solution to the constrained inclusion of $\mathcal{A}(y_{i,\mathsf{k}}+y_{i,\mathsf{r}})\owns\mathcal{Q}(x_{i,\mathsf{r}}-y_{i,\mathsf{r}})$, s.t. $\mathcal{Q}y_{i,\mathsf{k}}=0$, is unique, and satisfies $\|y_{1,\mathsf{k}}-y_{2,\mathsf{k}}\|\leq\xi\|x_{1,\mathsf{r}}-x_{2,\mathsf{r}}\|$ for some constant $\xi$;
(ii) the solution to the inclusion $\mathcal{A}(\mathcal{E}^{\top}b_{i}+y_{i,\mathsf{r}})\owns\mathcal{Q}(x_{i,\mathsf{r}}-y_{i,\mathsf{r}})$ is unique, and satisfies $\|b_{1}-b_{2}\|\leq\xi\|x_{1,\mathsf{r}}-x_{2,\mathsf{r}}\|$ for some constant $\xi$, where $\mathcal{E}$ is defined in Theorem 4.2-(iii).
The equivalence result is presented below.
Proposition 5.5.
Under Assumptions 1, 3 and the same setting of Theorem 4.2, then the following 3 conditions are equivalent:
(i) Assumption 6-(i);
(ii) Assumption 6-(ii);
(iii) Lipschitz continuity of $(\mathcal{A}+\mathcal{Q})^{-1}$ [10, Theorem 2.9].
The proof is postponed in Appendix C. Indeed, all the conditions in Proposition 5.5 are equivalent to $\|\mathcal{T}x_{1}-\mathcal{T}x_{2}\|\leq\xi\|x_{1}-x_{2}\|_{\mathcal{Q}}$, which can be found in [32, Theorem 3.3] and [11, Lemma 3.2].
Then, the convergence in the whole space immediately follows.
Theorem 5.6.
Let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by (2), which satisfies Assumptions 1, 3, 5 and any condition of Theorem 4.2 and Proposition 5.5. Furthermore, if any of the following conditions hold:
(i) every weak sequential cluster point of $\{x^{k}\}_{k\in\mathbb{N}}$ lies in $\mathsf{Fix}\mathcal{T}$;
(ii) Assumption 4;
(iii) $\mathcal{I}-\mathcal{T}$ is demiclosed,
then $\{x^{k}\}_{k\in\mathbb{N}}$ weakly converges to some $x^{\star}\in\mathsf{zer}\mathcal{A}$, as $k\rightarrow\infty$.
Proof 5.7.
Proposition 5.5 guarantees the boundedness of $\{x^{k}\}_{k\in\mathbb{N}}$. Then there exist two subsequences $\{x^{k_{i}}\}_{k\in\mathbb{N}}$ and $\{x^{k_{j}}\}_{k\in\mathbb{N}}$, such that $x^{k_{i}}\rightharpoonup v_{1}^{*}$ and $x^{k_{j}}\rightharpoonup v_{2}^{*}$. By Step-3 of the proof of Theorem 5.3, we have $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}v_{1}^{*}=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}v_{2}^{*}$ and further $\mathcal{T}v_{1}^{*}=\mathcal{T}v_{2}^{*}$, i.e., all the weak sequential cluster points of $\{x^{k}\}_{k\in\mathbb{N}}$ share the same projection onto $\mathsf{ran}\mathcal{Q}$.
Condition (i): it implies that $v_{1}^{*}=\mathcal{T}v_{1}^{*}=\mathcal{T}v_{2}^{*}=v_{2}^{*}$. This shows that the weak sequential cluster point of $\{x^{k}\}_{k\in\mathbb{N}}$ is unique, and lies in $\mathsf{Fix}\mathcal{T}=\mathsf{zer}\mathcal{A}$. The weak convergence of $\{x^{k}\}_{k\in\mathbb{N}}$ is obtained by [4, Lemma 2.46].
Condition (ii): Assumption 4-(i) implies that $\mathsf{gra}\mathcal{A}$ is closed in $\mathcal{H}^{\text{weak}}\times\mathcal{H}^{\text{strong}}$. Take a subsequence $\{x^{k_{i}}\}_{k\in\mathbb{N}}$, such that $x^{k_{i}}\rightharpoonup v^{*}$. Then $\mathcal{T}x^{k_{i}}\rightharpoonup v^{*}$, since $x^{k_{i}}-\mathcal{T}x^{k_{i}}\rightarrow 0$. Considering $\mathcal{A}\mathcal{T}x^{k_{i}}\owns\mathcal{Q}(x^{k_{i}}-\mathcal{T}x^{k_{i}})$, combining with $\mathcal{T}x^{k_{i}}\rightharpoonup v^{*}$ and $x^{k_{i}}-\mathcal{T}x^{k_{i}}\rightarrow 0$, it then follows from the closedness of $\mathsf{gra}\mathcal{A}$ that $v^{\star}\in\mathsf{zer}\mathcal{A}=\mathsf{Fix}\mathcal{T}$, which leads to Condition (i).
Condition (iii): Considering the subsequence $\{x^{k_{i}}\}_{k\in\mathbb{N}}$ in (ii), the demiclosedness of $\mathcal{I}-\mathcal{T}$ yields that $(\mathcal{I}-\mathcal{T})v^{*}=0$, i.e., $v^{*}\in\mathsf{Fix}\mathcal{T}$, which also leads to Condition (i).
Conditions (i) and (ii) correspond to [10, Theorem 2.9, Corollary 2.10], respectively. Condition (iii), as mentioned in [10, Remark 2.11], is hard to control. The key problem here is that $\mathcal{T}$ is not necessarily nonexpansive, and thus the general demiclosedness of $\mathcal{I}-\mathcal{T}$ cannot be obtained by Browder’s principle [4, Theorem 4.27]. However, it can be tackled case-by-case, by exploiting particular structures of $\mathcal{A}$ and $\mathcal{Q}$, and additional assumptions. For instance, [11, Lemma 3.4], in a context of a specific class of preconditioned ADMM, proposed several additional assumptions (e.g., compactness, weak closedness) to make $\mathcal{I}-\mathcal{T}$ demiclosed.
[16, Theorem 3.3] seems the first attempt to establish the whole convergence under degenerate metric in a finite-dimensional setting, based on the continuity of proximity operator. This technique, together with [22, Theorem 3.4], is not applicable for infinite-dimensional setting, since $\mathcal{T}$ is generally not weakly continuous.
6 Applications to operator splitting algorithms
6.1 Douglas–Rachford splitting (DRS)
The scheme of DRS has been given as (3), with the degenerate PPA interpretation of (4). The well-definedness of the DRS is obvious from (3). It can also be verified from (4) only via Assumption 4. First, it is clear that $\mathcal{A}:\mathcal{U}^{3}\mapsto 2^{\mathcal{U}}\times 2^{\mathcal{U}}\times\mathcal{U}$ is maximally monotone. Second, observe that $\mathsf{ran}\mathcal{A}=\mathcal{U}^{3}=\mathcal{H}$ and $0\in\mathsf{ran}\mathcal{Q}$, and thus, $0\in\mathsf{sri}(\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A})$, which fulfills Assumption 4-(ii). Then it follows that the scheme (3) admits any $(u,w,z)\in\mathcal{U}^{3}$ in any iterative step $k$.
A standard result in the seminal work [23] claims $z^{k}\rightharpoonup z^{\star}$ as $k\rightarrow\infty$, if $\tau>0$. This can also easily be obtained by our Theorem 5.3. The weak convergence of the solution trajectory $\{u^{k}\}_{k\in\mathbb{N}}$ had been a longstanding problem until the advent of [30]. To obtain this result, [30] used some sophisticated tools, e.g., quasi-Fejér convergence, strong/weak topologies. In [33, Theorem 4.2], the author used many variants of PPA forms of DRS (especially the complicated prediction+correction framework initialized by [21, 20, 19]). Now, the next Proposition 6.1 shows that the convergence of $\{u^{k}\}_{k\in\mathbb{N}}$ and $\{w^{k}\}_{k\in\mathbb{N}}$ can be directly obtained from (4) alone and via our Theorem 5.6. This is an improvement, compared to [33, Remark 3.1-(iii)], which once claimed that the convergences of $\{u^{k}\}_{k\in\mathbb{N}}$ and $\{w^{k}\}_{k\in\mathbb{N}}$ remain unclear from the PPA form of (4) only. This demonstrates the great potential of using the PPA interpretation. In addition, similar to [30], Proposition 6.1 also removes the assumption of maximal monotonicity of $\partial f+\partial g$, but without using the instrumental result of [3, Corollary 3] as [30] did.
Proposition 6.1.
Let $\{u^{k},w^{k},z^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by (3). If $\tau>0$, then $(u^{k},w^{k},z^{k})\rightharpoonup(u^{\star},w^{\star},z^{\star})$ as $k\rightarrow\infty$, where $u^{\star}$ is an optimal solution to $\min_{u}f(u)+g(u)$.
Proof 6.2.
First, $\tau>0$ guarantees the positive semi-definiteness of $\mathcal{Q}$. Since it is difficult to compute $\mathcal{A}^{-1}$, let us consider the inclusion of $\mathcal{A}(\mathcal{E}^{\top}b+y_{\mathsf{r}})\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$, and check if it has a unique solution of $b$ and if $b$ satisfies Assumption 6.
In the setting of (4), $\mathcal{E}=\begin{bmatrix}I&0&0\\ 0&I&0\end{bmatrix}$. Let $b=\begin{bmatrix}b_{1}\\ b_{2}\end{bmatrix}$, $x_{\mathsf{r}}=(0,0,a)$, and the corresponding $y_{\mathsf{r}}=(0,0,c)$ [^3] There is no need to find the specific expression of $c$ in terms of $a$. Of great importance here is the fact that $c$ is a nonexpansive mapping of $a$ by Fact 2., then the concrete form of $\mathcal{A}(\mathcal{E}^{\top}b+y_{\mathsf{r}})\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$ becomes:
$\begin{bmatrix}\tau\partial g&I&-I\\ -I&\tau\partial f&I\\ I&-I&0\end{bmatrix}\begin{bmatrix}b_{1}\\ b_{2}\\ c\end{bmatrix}\owns\begin{bmatrix}0\\ 0\\ a-c\end{bmatrix}.$
Putting $b_{2}=b_{1}-a+c$ (indicated by the last line) into the first line, we obtain $\tau\partial g(b_{1})+b_{1}-a+c-c=\tau\partial g(b_{1})+b_{1}-a\owns 0$, leading to $b_{1}=\operatorname{prox}_{\tau g}(a)$, and thus, $b_{2}=\operatorname{prox}_{\tau g}(a)-a+c=c-\operatorname{prox}_{g^{*}/\tau}(a)$. $b$ is obviously single-valued, which meets Theorem 4.2. Then, from the well-known nonexpansiveness of proximity operator (see [15, Lemma 2.4] for example), we have
$\|b-b^{\prime}\|$$\leq$ $\|b_{1}-b^{\prime}_{1}\|+\|b_{2}-b^{\prime}_{2}\|$
$=$ $\big{\|}\operatorname{prox}_{\tau g}(a)-\operatorname{prox}_{\tau g}(a^{\prime})\big{\|}+\big{\|}c-\operatorname{prox}_{g^{*}/\tau}(a)-c^{\prime}+\operatorname{prox}_{g^{*}/\tau}(a^{\prime})\big{\|}$
$\leq$ $\big{\|}a-a^{\prime}\big{\|}+\big{\|}a-a^{\prime}\big{\|}+\big{\|}c-c^{\prime}\big{\|}$
$\leq$ $3\big{\|}a-a^{\prime}\big{\|},$
where the last inequality is due to the $\|c-c^{\prime}\|\leq\|a-a^{\prime}\|$ by Fact 2. Finally, the weak convergence of $\{u^{k},w^{k},z^{k}\}_{k\in\mathbb{N}}$ follows by Theorem 5.6.
It is worth noting that the weak convergence of $\{u^{k}\}_{k\in\mathbb{N}}$ cannot be directly obtained from $u^{k}=\operatorname{prox}_{\tau g}(z^{k})$ and the weak convergence of $\{z^{k}\}_{k\in\mathbb{N}}$, since the proximity operator is not weak-to-weak continuous. The Lipschitz continuity of $u^{k}$ (w.r.t. and bounded by $z^{k}$) has to be incorporated into the degenerate PPA framework.
6.2 Alternating direction method of multipliers (ADMM)
Considering an equality constrained problem:
$\min_{s,t}f(s)+g(t),\quad\text{s.t.\ }As+Bt=0,$
the ADMM is given as [33, Eq.(22)]:
$\left\lfloor\begin{array}[]{lll}s^{k+1}&\in&\operatorname{Arg}\min_{s}f(s)+\frac{\tau}{2}\big{\|}As+Bt^{k}-\frac{1}{\tau}u^{k}\big{\|}^{2},\\ t^{k+1}&\in&\operatorname{Arg}\min_{t}g(t)+\frac{\tau}{2}\big{\|}As^{k+1}+Bt-\frac{1}{\tau}u^{k}\big{\|}^{2},\\ u^{k+1}&=&u^{k}-\tau(As^{k+1}+Bt^{k+1}).\\ \end{array}\right.$ (17)
Due to the equivalence between ADMM and DRS, we now transform the ADMM into a DRS form. Let $z^{k}=u^{k}+\tau B^{\top}t^{k}$, then, (17) becomes:
$\left\lfloor\begin{array}[]{lll}t^{k}&\in&\operatorname{Arg}\min_{t}g(t)+\frac{\tau}{2}\big{\|}Bt-\frac{1}{\tau}z^{k}\big{\|}^{2},\\ s^{k+1}&\in&\operatorname{Arg}\min_{s}f(s)+\frac{\tau}{2}\big{\|}As+2Bt^{k}-\frac{1}{\tau}z^{k}\big{\|}^{2},\\ z^{k+1}&=&z^{k}-\tau(As^{k+1}+Bt^{k}).\\ \end{array}\right.$ (18)
Similar to [33, Remark 3.1-(ii)], (18) can be rewritten in terms of infinal postcomposition [4, Definition 12.34]:
$\left\lfloor\begin{array}[]{lll}q^{k}&=&\arg\min_{q}G(q)+\frac{\tau}{2}\big{\|}q-\frac{1}{\tau}z^{k}\big{\|}^{2},\\ r^{k+1}&=&\arg\min_{r}F(r)+\frac{\tau}{2}\big{\|}r+2q^{k}-\frac{1}{\tau}z^{k}\big{\|}^{2},\\ z^{k+1}&=&z^{k}-\tau(r^{k+1}+q^{k}),\\ \end{array}\right.$ (19)
where $F=A\triangleright f$ and $G=B\triangleright g$ are infimal postcompositions of $f$ and $g$ by $A$ and $B$, respectively. Then, it is easy to see that (19) exactly fits into the form of DRS (3). Finally, by Proposition 6.1, it is easy to show the weak convergence of $\big{\{}q^{k},r^{k},z^{k}\big{\}}_{n\in\mathbb{N}}$.
The limiting case of the well-known Chambolle–Pock (CP) algorithm [14, Algorithm 1] can also be analyzed in a similar manner, based an augmentation technique [25], which equivalently expresses the CP algorithm as the DRS form (also see [33, Sect. 5.3] for details).
6.3 Augmented Lagrangian method (ALM)
Let us now consider a constrained optimization problem:
$\min_{u\in\mathcal{U}}f(u),\qquad\text{s.t.\ }Bu=b,$
where $B:\mathcal{U}\mapsto\mathcal{Y}$. The ALM reads as (see [31, Eq.(7.2)] for example):
$\left\lfloor\begin{array}[]{lll}u^{k+1}&\in&\operatorname{Arg}\min_{u}f(u)+\frac{\tau}{2}\big{\|}Bu-b-\frac{1}{\tau}p^{k}\big{\|}^{2},\\ p^{k+1}&=&p^{k}-\tau(Bu^{k+1}-b).\\ \end{array}\right.$ (20)
Using the notion of infimal postcomposition mentioned above, we obtain the equivalent form to (20):
$\left\lfloor\begin{array}[]{lll}q^{k+1}&=&\arg\min_{q}F(q)+\frac{\tau}{2}\big{\|}q-b-\frac{1}{\tau}p^{k}\big{\|}^{2},\\ p^{k+1}&=&p^{k}-\tau(q^{k+1}-b),\\ \end{array}\right.$ (21)
where $F^{*}=B\triangleright f$. The corresponding PPA interpretation is
$\begin{bmatrix}0\\ 0\end{bmatrix}\in\underbrace{\begin{bmatrix}\partial F&-I\\ I&\partial l\end{bmatrix}}_{\mathcal{A}}\begin{bmatrix}q^{k+1}\\ p^{k+1}\end{bmatrix}+\underbrace{\begin{bmatrix}0&0\\ 0&\frac{1}{\tau}I\end{bmatrix}}_{\mathcal{Q}}\begin{bmatrix}q^{k+1}-q^{k}\\ p^{k+1}-p^{k}\end{bmatrix},$ (22)
where the linear function $l$ is defined as $l=-\langle\cdot|b\rangle$.
Theorem 6.3.
Let $\{q^{k},p^{k}\}_{k\in\mathbb{N}}$ be a sequence generated by (21). If $\tau>0$, then $(q^{k},p^{k})\rightharpoonup(b,p^{\star})$ as $k\rightarrow\infty$.
Proof 6.4.
First, observing (22) obtains that $\mathcal{A}$ is maximally monotone, $\mathsf{ran}\mathcal{A}=\mathcal{Y}^{2}=\mathcal{H}$, and thus, $0\in\mathsf{sri}(\mathsf{ran}\mathcal{Q}-\mathsf{ran}\mathcal{A})$, which fulfills Assumption 4. Then it follows that the scheme (21) admits any input $x=(q,p)\in\mathcal{Y}^{2}$.
Then, we check Theorem 4.2 and 6. Here, $\mathcal{E}=\begin{bmatrix}I&0\end{bmatrix}$. Given an input $x_{\mathsf{r}}=(0,a)$ and the corresponding output $y_{\mathsf{r}}=(0,c)$, the inclusion of $\mathcal{A}(\mathcal{E}^{\top}v+y_{\mathsf{r}})\owns\mathcal{Q}(x_{\mathsf{r}}-y_{\mathsf{r}})$ becomes:
$\begin{bmatrix}\partial F&-I\\ I&\partial l\end{bmatrix}\begin{bmatrix}v\\ c\end{bmatrix}\owns\begin{bmatrix}0\\ \frac{1}{\tau}(a-c)\end{bmatrix}.$
The second line yields that $v=\frac{1}{\tau}(a-c)+b$, which is obviously single-valued (satisfying Theorem 4.2). And furthermore, $\|v-v^{\prime}\|=\frac{1}{\tau}\|(a-a^{\prime})-(c-c^{\prime})\|\leq\frac{2}{\tau}\|a-a^{\prime}\|$, i.e., $\frac{2}{\tau}$-Lipschitz continuous w.r.t. $a$. Finally, the weak convergence of $\{q^{k},p^{k}\}_{k\in\mathbb{N}}$ follows by Theorem 5.6.
The basic Bregman iteration [26] can be analyzed in a similar manner, due to its essential equivalence to the ALM (20).
7 Concluding remarks
This work provided a fundamental understanding of degenerate metric resolvent and its associated convergence issue. Our results can be further exemplified with more applications:
  • Analyze the existence and uniqueness of solution to the general regularized least-squares problem: $\min_{x\in\mathbb{R}^{n}}f(x)+\frac{1}{2}\|Ax-b\|^{2}$, particularly generalized lasso (corresponding to $f=\|D\cdot\|_{1}$), and make close connections to [17].
  • Discuss the convergence property of $\{t^{k},s^{k}\}_{k\in\mathbb{N}}$ in the non-proximized ADMM (18), This depends on the interatctions between $A$, $B$ and $f$, $g$. Take a similar treatment to $\{u^{k}\}_{k\in\mathbb{N}}$ in (20).
Some theoretical aspects are also of great interest, for instance,
  • Explore the properties of $\mathcal{T}$, if $\mathcal{Q}$ fails to have closed range, as indicated in [6].
  • Study the behaviours of (2) if the inclusion problem (1) is inconsistent, as proposed in [2].
  • Find a simple interpretation of Assumption 4-(ii) could also be interesting for practitioners.
8 Acknowledgements
We are gratefully indebted to Mr. Brecht Evens (KU Leuven, Belgium) for inspiring comments and bringing [22, Theorem 3.4] to our attention in the early stages of this work. We would also like to thank Mr. Shuchang Zhang (NUDT, China) for detailed discussions on [33, Theorem 2.1].
Appendix A Generalized Moreau’s idendity and Proof of Lemma 2.10
Lemma A.1 extends the classical Moreau’s identity to the degenerate case. This is a key ingredient for proving Lemma 2.10. We stress that this extension holds under Assumption 1 only, even without the restricted monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ only. This greatly weakens the previous result of [33, Lemma 2.3].
Lemma A.1 (Generalized Moreau’s decomposition identity).
Under Assumption 1, we have $\mathcal{I}\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}+(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}$, for any input $x\in\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}=\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$;
Proof A.2.
For any input $x\in\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}$, pick $y$, such that $y\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$. Then, we develop $\mathcal{Q}x\in\mathcal{A}y+\mathcal{Q}y\Longrightarrow\mathcal{Q}(x-y)\in\mathcal{A}y\Longrightarrow y\in\mathcal{A}^{-1}\mathcal{Q}(x-y)\Longrightarrow x-y\in x-\mathcal{A}^{-1}\mathcal{Q}(x-y)\Longrightarrow x-y\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x\Longrightarrow x\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x+(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$, from which follows that $\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}\subseteq\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$.
Conversely, For any input $x\in\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$, pick $y$, such that $y\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x$. Then, we develop $y+\mathcal{A}^{-1}\mathcal{Q}y\owns x\Longrightarrow x-y\in\mathcal{A}^{-1}\mathcal{Q}y\Longrightarrow\mathcal{Q}y\in\mathcal{A}(x-y)\Longrightarrow\mathcal{Q}(y-x+x)\in\mathcal{A}(x-y)\Longrightarrow\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})(x-y)\Longrightarrow x-y\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x\Longrightarrow x\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x+(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$, from which follows that $\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}\supseteq\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$.
Both directions show the desired result, with $\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}=\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$.
Let us now show that all the expressions appear in Lemma 2.10 are single-valued under Assumptions 1 and 2.
Fact 3.
Under Assumptions 1 and 2, the following hold.
(i) $\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}$ is single-valued;
(ii) $(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}$ is single-valued;
(iii) $\sqrt{\mathcal{Q}}\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}$ is single-valued.
Proof A.3.
(i) Given $x\in\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$, let $\{y_{1},y_{2}\}\subseteq(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x$, then, $y_{1}+\mathcal{A}^{-1}\mathcal{Q}y_{1}\owns x$ and $y_{2}+\mathcal{A}^{-1}\mathcal{Q}y_{2}\owns x$, leading to $y_{2}-y_{1}\in\mathcal{A}^{-1}\mathcal{Q}y_{1}-\mathcal{A}^{-1}\mathcal{Q}y_{2}$. Then,
$0\geq-\|y_{1}-y_{2}\|_{\mathcal{Q}}^{2}=\big{\langle}\mathcal{Q}y_{1}-\mathcal{Q}y_{2}\big{|}y_{2}-y_{1}\big{\rangle}\geq 0,$
where the last inequality is due to $y_{2}-y_{1}\in\mathcal{A}^{-1}\mathcal{Q}y_{1}-\mathcal{A}^{-1}\mathcal{Q}y_{2}$ and the local monotonicity of $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$. Finally, we have $\mathcal{Q}y_{1}=\mathcal{Q}y_{2}$, which shows the single-valuedness of $\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}$.
(ii) Take $x\in\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}\big{)}$. Assume, by contradiction, that $\{y_{1},y_{2}\}\subseteq(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}x$, then, $y_{1}+\mathcal{Q}\mathcal{A}^{-1}y_{1}\owns\mathcal{Q}x$ and $y_{2}+\mathcal{Q}\mathcal{A}^{-1}y_{2}\owns\mathcal{Q}x$. This implies that $\{y_{1},y_{2}\}\subseteq\mathsf{ran}\mathcal{Q}$. Let $y_{1}=\mathcal{Q}v_{1}$ and $y_{2}=\mathcal{Q}v_{2}$, we have $\mathcal{Q}v_{1}+\mathcal{Q}\mathcal{A}^{-1}\mathcal{Q}v_{1}\owns\mathcal{Q}x$ and $\mathcal{Q}v_{2}+\mathcal{Q}\mathcal{A}^{-1}\mathcal{Q}v_{2}\owns\mathcal{Q}x$. Thus, $v_{1}+\mathcal{A}^{-1}\mathcal{Q}v_{1}\owns x+s_{1}$ and $v_{2}+\mathcal{A}^{-1}\mathcal{Q}v_{2}\owns x+s_{2}$, $\forall s_{1},s_{2}\in\mathsf{ker}\mathcal{Q}$. Thus, due to the monotonicity of $\mathcal{A}^{-1}$, we have:
00 0 $\leq$ $\langle\mathcal{Q}v_{1}-\mathcal{Q}v_{2}|t_{1}-t_{2}\rangle\quad\forall t_{i}\in\mathcal{A}^{-1}\mathcal{Q}v_{i}\text{\ for\ }i=1,2$
$=$ $\langle\mathcal{Q}v_{1}-\mathcal{Q}v_{2}|(x+s_{1}-v_{1})-(x+s_{2}-v_{2})\rangle\quad\text{choose\ }t_{i}=x+s_{i}-v_{i}\text{\ for\ }i=1,2$
$=$ $\langle v_{1}-v_{2}|\mathcal{Q}(s_{1}-v_{1}-s_{2}+v_{2})\rangle$
$=$ $-\|v_{1}-v_{2}\|_{\mathcal{Q}}^{2},\quad\text{since $s_{1},s_{2}\in\mathsf{ker}\mathcal{Q}$}$
which leads to $\mathcal{Q}v_{1}=\mathcal{Q}v_{2}$, i.e., $y_{1}=y_{2}$.
(iii) It suffices to show that $\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}$ is single-valued. Take $x\in\mathsf{dom}\big{(}\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}\big{)}$. Assume, by contradiction, that $\{y_{1},y_{2}\}\subseteq\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}x$. Then, $y_{1}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}y_{1}\owns\sqrt{\mathcal{Q}}x$ and $y_{2}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}y_{2}\owns\sqrt{\mathcal{Q}}x$. This implies that $\{y_{1},y_{2}\}\subseteq\mathsf{ran}\mathcal{Q}$. Let $y_{1}=\sqrt{\mathcal{Q}}v_{1}$ and $y_{2}=\sqrt{\mathcal{Q}}v_{2}$, we have $\sqrt{\mathcal{Q}}v_{1}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}v_{1}\owns\sqrt{\mathcal{Q}}x$ and $\sqrt{\mathcal{Q}}v_{2}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}v_{2}\owns\sqrt{\mathcal{Q}}x$. Thus, $v_{1}+\mathcal{A}^{-1}\mathcal{Q}v_{1}\owns x+s_{1}$ and $v_{2}+\mathcal{A}^{-1}\mathcal{Q}v_{2}\owns x+s_{2}$, $\forall s_{1},s_{2}\in\mathsf{ker}\mathcal{Q}$. Thus, due to the monotonicity of $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$, we have:
00 0 $\leq$ $\langle\mathcal{Q}v_{1}-\mathcal{Q}v_{2}|t_{1}-t_{2}\rangle\quad\forall t_{i}\in\mathcal{A}^{-1}\mathcal{Q}v_{i}\text{\ for\ }i=1,2$
$=$ $\langle\mathcal{Q}v_{1}-\mathcal{Q}v_{2}|(x+s_{1}-v_{1})-(x+s_{2}-v_{2})\rangle\quad\text{choose\ }t_{i}=x+s_{i}-v_{i}\text{\ for\ }i=1,2$
$=$ $\langle v_{1}-v_{2}|\mathcal{Q}(s_{1}-v_{1}-s_{2}+v_{2})\rangle$
$=$ $-\|v_{1}-v_{2}\|_{\mathcal{Q}}^{2},\quad\text{since $s_{1},s_{2}\in\mathsf{ker}\mathcal{Q}$}$
which leads to $\mathcal{Q}v_{1}=\mathcal{Q}v_{2}$, i.e., $y_{1}=y_{2}$.
We are now ready to prove Lemma 2.10.
Proof A.4.
First, note that the single-valuedness of all the expressions here has been shown in Fact 3. We show the equalities following the direction of $(1)\Longrightarrow(2)\Longrightarrow(3)\Longrightarrow(4)\Longrightarrow(1)$.
(1): Apply $\mathcal{Q}$ to both sides of Lemma A.1, use Lemma 2.3 and Fact 3.
(2): We begin with $\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}$. Take $x\in\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}$. Let $v\in(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}x$ and $y=\mathcal{Q}v$. Then, we deduce that $v+\mathcal{A}^{-1}\mathcal{Q}v\owns x\Longrightarrow\mathcal{Q}v+\mathcal{Q}\mathcal{A}^{-1}\mathcal{Q}v\owns\mathcal{Q}x\Longrightarrow y+\mathcal{Q}\mathcal{A}^{-1}y\owns\mathcal{Q}x\Longrightarrow y=(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}x$. This also shows that $\mathsf{dom}\big{(}\mathcal{Q}(\mathcal{I}+\mathcal{A}^{-1}\mathcal{Q})^{-1}\big{)}\subseteq\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}\big{)}$.
(3): We begin with $(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}$. Take $x\in\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}\big{)}$. Let $y=(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}x$, then, $y+\mathcal{Q}\mathcal{A}^{-1}y\owns\mathcal{Q}x$. This implies that $y\in\mathsf{ran}\mathcal{Q}$. Let $y=\sqrt{\mathcal{Q}}v$ with $v\in\mathsf{ran}\mathcal{Q}$ [^4] Strictly speaking, we should write $y=\sqrt{\mathcal{Q}}(v+s)$ with $s\in\mathsf{ker}\mathcal{Q}$. This is essentially equivalent to $y=\sqrt{\mathcal{Q}}v$.. Then, the inclusion becomes $\sqrt{\mathcal{Q}}v+\mathcal{Q}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}v\owns\mathcal{Q}x\Longrightarrow v+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}v\owns\sqrt{\mathcal{Q}}x\Longrightarrow v=\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}x\Longrightarrow y=\sqrt{\mathcal{Q}}v=\sqrt{\mathcal{Q}}\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}x$. This also shows that $\mathsf{dom}\big{(}(\mathcal{I}+\mathcal{Q}\mathcal{A}^{-1})^{-1}\mathcal{Q}\big{)}\subseteq\mathsf{dom}\big{(}\sqrt{\mathcal{Q}}\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}\big{)}$.
(4): It suffices to show $\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}=\Big{(}\mathcal{I}+\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}^{-1}\sqrt{\mathcal{Q}}$. Take $x\in\mathsf{dom}\big{(}\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}\big{)}$. Let $y=\big{(}\mathcal{I}+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}x$, then, $y+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}y\owns\sqrt{\mathcal{Q}}x$. This implies that $y\in\mathsf{ran}\mathcal{Q}$, which already satisfies the domain restricted in $\mathsf{ran}\mathcal{Q}$, i.e., $y+\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}|_{\mathsf{ran}\mathcal{Q}}(y)\owns\sqrt{\mathcal{Q}}x$, leading to the right-hand side of (4).
(4$\Longrightarrow$(1): It suffices to show that $\Big{(}\mathcal{I}+\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}^{-1}\sqrt{\mathcal{Q}}=\sqrt{\mathcal{Q}}-\sqrt{\mathcal{Q}}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}$. Take $x\in\mathsf{dom}\big{(}\big{(}\mathcal{I}+(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}})|_{\mathsf{ran}\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}\big{)}$. Let $y=\big{(}\mathcal{I}+(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}})|_{\mathsf{ran}\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}x$, then, $y+(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}})|_{\mathsf{ran}\mathcal{Q}}(y)\owns\sqrt{\mathcal{Q}}x$. This implies that $y\in\mathsf{ran}\mathcal{Q}$, and thus, the above inclusion can be simplified as $y+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}y\owns\sqrt{\mathcal{Q}}x$, and let $y=\sqrt{\mathcal{Q}}v$ with $v\in\mathsf{ran}\mathcal{Q}$. Then, we have $\sqrt{\mathcal{Q}}v+\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\mathcal{Q}v\owns\sqrt{\mathcal{Q}}x\Longrightarrow v+\mathcal{A}^{-1}\mathcal{Q}v\owns x\Longrightarrow\mathcal{A}^{-1}\mathcal{Q}v\owns x-v\Longrightarrow\mathcal{Q}v\in\mathcal{A}(x-v)\Longrightarrow\mathcal{Q}(v-x+x)\in\mathcal{A}(x-v)\Longrightarrow\mathcal{Q}x\in(\mathcal{A}+\mathcal{Q})(x-v)\Longrightarrow x-v\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x\Longrightarrow v\in x-(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x\Longrightarrow y=\sqrt{\mathcal{Q}}v=\sqrt{\mathcal{Q}}(\mathcal{I}-(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q})x$. This also shows that $\mathsf{dom}\big{(}\big{(}\mathcal{I}+(\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}})|_{\mathsf{ran}\mathcal{Q}}\big{)}^{-1}\sqrt{\mathcal{Q}}\big{)}\subseteq\mathsf{dom}\big{(}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}\big{)}$.
Finally, we show all the equalities with exactly same domains.
Appendix B An alternative proof of Theorem 5.3
The proof is presented below.
Proof B.1.
Applying both sides of Lemma 2.10-(4) to $x_{\mathsf{r}}^{k}$, we have:
$\sqrt{\mathcal{Q}}x_{\mathsf{r}}^{k}-\sqrt{\mathcal{Q}}x_{\mathsf{r}}^{k+1}=\Big{(}\mathcal{I}+\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}\big{|}_{\mathsf{ran}\mathcal{Q}}\Big{)}^{-1}\sqrt{\mathcal{Q}}x_{\mathsf{r}}^{k},$ (23)
where the single-valuedness of the right-hand side has been shown in the proof of Lemma 2.10. Let $q_{\mathsf{r}}^{k}:=\sqrt{\mathcal{Q}}x_{\mathsf{r}}^{k}$, $\mathcal{B}:=\big{(}\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}\big{)}\big{|}_{\mathsf{ran}\mathcal{Q}}$, then, (23) is rewritten as (by maximal monotonicity of $\mathcal{B}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ and celebrated Moreau’s identity):
$q_{\mathsf{r}}^{k+1}=q_{\mathsf{r}}^{k}-(\mathcal{I}+\mathcal{B})^{-1}q_{\mathsf{r}}^{k}=\big{(}\mathcal{I}+\mathcal{B}^{-1}\big{)}^{-1}q_{\mathsf{r}}^{k}\Longrightarrow\mathcal{B}^{-1}q_{\mathsf{r}}^{k+1}\owns q_{\mathsf{r}}^{k}-q_{\mathsf{r}}^{k+1}.$
It is known from the proof of Theorem 5.3 that $q_{\mathsf{r}}^{k}-q_{\mathsf{r}}^{k+1}\rightarrow 0$ and $q_{\mathsf{r}}^{k}\rightharpoonup q^{\star}$, as $k\rightarrow\infty$. Since $\mathcal{B}^{-1}$ is also maximally monotone (by [4, Proposition 20.22]), then $\mathsf{gra}\mathcal{B}^{-1}$ is sequentially closed in $(\mathsf{ran}\mathcal{Q})^{\text{weak}}\times(\mathsf{ran}\mathcal{Q})^{\text{strong}}$, i.e., $(\mathsf{ran}\mathcal{Q},\|\cdot\|_{\mathcal{Q}})\times(\mathsf{ran}\mathcal{Q},\langle\cdot|\cdot\rangle_{\mathcal{Q}})$, by [4, Proposition 20.38]. Thus, we obtain
$0\in\mathcal{B}^{-1}q^{\star}\Longrightarrow q^{\star}\in\mathcal{B}(0)=\sqrt{\mathcal{Q}}\mathcal{A}^{-1}(0)\Longrightarrow x_{\mathsf{r}}^{\star}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}(0)=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{zer}\mathcal{A}).$
Appendix C Proof of Proposition 5.5
Before proving this, we present several simple facts.
Fact 4.
Under Assumptions 1 and 2, given $\mathcal{T}$ in (2), then $\forall x\in\mathsf{ran}\mathcal{Q}$, $\exists\lambda_{\inf},\lambda_{\sup}>0$, such that
(i) $\lambda_{\inf}\|x\|^{2}\leq\|x\|_{\mathcal{Q}}^{2}\leq\lambda_{\sup}\|x\|^{2}$;
(ii) $\lambda_{\inf}\|x\|\leq\|\mathcal{Q}x\|\leq\lambda_{\sup}\|x\|$;
(iii) $\|\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{T}x_{1}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{T}x_{2}\|\leq\sqrt{\frac{\lambda_{\sup}}{\lambda_{\inf}}}\|x_{1,\mathsf{r}}-x_{2,\mathsf{r}}\|$, $\forall x_{1},x_{2}\in\mathsf{dom}\mathcal{T}$.
Proof C.1.
(i)-(ii): Due to the boundedness and closedness of $\mathcal{Q}$ as Assumption 1.
(iii): by Fact 4-(i) and Fact 2.
Then, we prove Proposition 5.5.
Proof C.2.
(i)$\Longleftrightarrow$(ii): $\|y_{1,\mathsf{k}}-y_{2,\mathsf{k}}\|^{2}=\|\mathcal{E}^{\top}b_{1}-\mathcal{E}^{\top}b_{2}\|^{2}=\|b_{1}-b_{2}\|^{2}$, since $\mathcal{E}\mathcal{E}^{\top}=I$.
(i)$\Longrightarrow$(iii):
$\|(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x_{1}-(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x_{2}\|$
$=$ $\|y_{1}-y_{2}\|=\|y_{1,\mathsf{r}}+y_{1,\mathsf{k}}-y_{2,\mathsf{r}}-y_{2,\mathsf{k}}\|$
$\leq$ $\|y_{1,\mathsf{r}}-y_{2,\mathsf{r}}\|+\|y_{1,\mathsf{k}}-y_{2,\mathsf{k}}\|$
[…]
(iii)$\Longrightarrow$(i):
$\|y_{1,\mathsf{k}}-y_{2,\mathsf{k}}\|$$\leq$ $\|y_{1}-y_{2}\|+\|y_{1,\mathsf{r}}-y_{2,\mathsf{r}}\|$
[…]